### v2

# parskip=half

Abstract
Causal survival analysis combines survival analysis and causal inference to evaluate the effect of a treatment or intervention on a time-to-event outcome, such as survival time. It offers an alternative to relying solely on Cox models for assessing these effects. In this paper, we present a comprehensive review of estimators for the average treatment effect measured with the restricted mean survival time, including regression-based methods, weighting approaches, and hybrid techniques. We investigate their theoretical properties and compare their performance through extensive numerical experiments. Our analysis focuses on the finite-sample behavior of these estimators, the influence of nuisance parameter selection, and their robustness and stability under model misspecification. By bridging theoretical insights with practical evaluation, we aim to equip practitioners with both state-of-the-art implementations of these methods and practical guidelines for selecting appropriate estimators for treatment effect estimation. Among the approaches considered, G-formula two-learners, AIPCW-AIPTW, Buckley-James estimators, and causal survival forests emerge as particularly promising.
1 Introduction
1.1 Context and motivations
Causal survival analysis is a growing field that integrates causal inference (Rubin,, 1974; Hernán MA,, 2020) with survival analysis (Kalbfleisch and Prentice,, 2002) to evaluate the impact of treatments on time-to-event outcomes, while accounting for censoring situations where only partial information about an event’s occurrence is available.
Although it is a relatively new domain, the literature has quickly expanded—often in a fragmented way. Researchers face a wide array of estimators and must make several methodological decisions, often without clear guidance. Moreover, many of the proposed methods lack accessible implementations. This is particularly important given that, in many real-world scenarios where randomized trials are infeasible, well-conducted observational studies—with explicit and plausible causal assumptions—remain the only viable source of evidence about the treatment effect (Makary and Prasad,, 2025). This calls for a structured review that consolidates available approaches, clarifies their assumptions, and provides an evaluation of estimator performance to make informed methodological choices — particularly in finite sample settings. Such a survey also has the potential to help identify remaining methodological challenges that need to be addressed.
Related efforts have already been made to clarify methodological options in survival analysis. For instance, Denz et al., (2023) focus on confounders-adjusted survival curves under proportional hazards and explore model misspecification in that context. However, existing reviews often remain within the Cox model framework, and may not fully address the diversity of estimators and causal estimands available—particularly those applicable when the proportional hazards assumption does not hold or when censoring is not independent.
Despite its widespread use in biomedical research, we decided not to cover the estimation of the Hazard Ratio (HR) which presents several well-documented limitations. First, when using Cox model, it relies on the proportional hazards assumption, which is not necessarely satisfied in practice. Second, hazard ratios are generally not causally interpretable—even in randomized controlled trials—due to built-in selection bias that can distort the comparison between groups unless the treatment has no effect at all (Martinussen and Vansteelandt,, 2013; Martinussen et al.,, 2020). Third, the HR is a non-collapsible measure, meaning that the effect estimated in the entire population (the marginal effect) can differ from the effect estimated within subgroups (the conditional effect), even when there is no confounding (Huitfeldt et al.,, 2019; Greenland et al.,, 1999). Finally, the HR could be difficult to interpret clinically, particularly because it reflects an instantaneous risk rather than a cumulative or absolute survival benefit (Averbuch et al.,, 2025). In contrast, survival curves are more informative than HRs and can be easily generated (Hernán,, 2010).
For all these reasons, we focus in this paper on the estimation of the Restricted Mean Survival Time (RMST), a popular causal measure in survival analysis that offers an intuitive interpretation of the average survival time over a specified time horizon.
Numerous methods have been developed for causal survival analysis, including approaches based on pseudo-observations (Andersen et al.,, 2017; Wang,, 2018; van Hage et al.,, 2024) and targeted maximum likelihood estimation (TMLE) (Stitelman et al.,, 2012; Talbot et al.,, 2025). In this work, we focus on estimators based on outcome regression, inverse probability weighting, and hybrid strategies.
1.2 Definition of the estimand: the RMST
We set the analysis in the potential outcome framework, where a patient, described by a vector of covariates $X\in\mathbb{R}^{p}$, either receives a treatment ($A=1$) or is in the control group ($A=0$). The patient will then survive up to a certain time $T^{(0)}\in\mathbb{R}^{+}$ in the control group, or up to a time $T^{(1)}\in\mathbb{R}^{+}$ in the treatment group. In practice, we cannot simultaneously have access to $T^{(0)}$ and $T^{(1)}$, as one patient is either treated or control, but only to $T$ defined as follows:
Assumption 1.1.
(Stable Unit Treatment Value Assumption — SUTVA)
$T=AT^{(1)}+(1-A)T^{(0)}$
Due to potential censoring, the outcome $T$ is not completely observed. The most common form of censoring is right-censoring (also known as type II censoring), which occurs when the event of interest has not taken place by the end of the observation period, indicating that it may have occurred later if the observation had continued (Turkson et al.,, 2021). We focus in this study on this type of censoring only and we assume that we observe $\tilde{T}=T\wedge C=\min(T,C)$ for some censoring time $C\in\mathbb{R}^{+}$. When an observation is censored, the observed time is equal to the censoring time.
We also assume that we know whether an outcome is censored or not. In other words, we observe the event status variable $\Delta=\mathbb{I}\{T\leqslant C\}$, where $\mathbb{I}\{\cdot\}$ is the indicator function. $\Delta$ is $1$ if the true outcome is observed, and 0 if it is censored.
We assume observing a $n$-sample of variables $(X,A,\widetilde{T},\Delta)$ stemming from an $n$-sample of the partially unobservable variables $(X,A,T^{(0)},T^{(1)},C)$. A toy example of such data is given in Table 1.
Table 1: Example of a survival dataset. Each observation corresponds to a single patient. In practice, only $X$, AA, $\Delta$, and $\widetilde{T}$ are observed. The column Obs. Outcome represents the observed survival outcome.
Covariates Treatment Potential outcome Outcome Censoring Status Obs. Outcome
$X_{1}$ $X_{2}$ $X_{3}$ AA $\quad T^{(0)}\;\;$ $^{(1)} TT$ $C$ $\Delta$ $\widetilde{T}$
1 1.5 4 1 ? 200 200 ? 1 200
5 1 2 0 100 ? 100 ? 1 100
9 0.5 3 1 ? ? ? 200 0 200
Our aim is to estimate the Average Treatment Effect (ATE) defined as the difference between the Restricted Mean Survival Time (Royston and Parmar 2013).
Definition 1.1.
(Causal effect: Difference between Restricted Mean Survival Time)
$\theta_{\mathrm{RMST}}=\mathbb{E}\left[T^{(1)}\wedge\tau-T^{(0)}\wedge\tau\right],$
where $a\wedge b:=\min(a,b)$ for $a,b\in\mathbb{R}$, and $\tau>0$ is a pre-specified time horizon over which survival is evaluated.
We define the survival functions $S^{(a)}(t):=\mathbb{P}(T^{(a)}>t)$ for $a\in\{0,1\}$, i.e., the probability that an individual will survive beyond a given time $t$ if treated or if not treated. Likewise, we let $S(t):=\mathbb{P}(T>t)$, and $S_{C}(t):=\mathbb{P}(C>t)$. We also let $G(t):=\mathbb{P}(C\geqslant t)$ be the left-limit of the survival function $S_{C}$. Because $T^{(a)}\wedge\tau$ are non-negative random variables, one can easily express the restricted mean survival time using the survival functions:
$\mathbb{E}(T^{(a)}\wedge\tau)=\int_{0}^{\infty}\mathbb{P}(T^{(a)}\wedge\tau>t)dt=\int_{0}^{\tau}S^{(a)}(t)dt.$} (1)
Consequently, $\theta_{\mathrm{RMST}}$ can be interpreted as the mean difference between the survival function of treated and control until a fixed time horizon $\tau$. A RMST difference $\theta_{\mathrm{RMST}}=10$ days with $\tau=200$ means that, on average, the treatment increases the survival time by 10 days at 200 days. We give a visual interpretation of RMST in Figure 1.
Figure 1: Plot of the estimated survival curves on synthetic toy-data. The $\theta_{\mathrm{RMST}}$ at $\tau=50$ corresponds to the yellow shaded area between the two survival curves. The curves have been estimated using the Kaplan-Meier estimator, see Section 2.1.
Although the present work focuses on the estimation of the difference in RMST, we would like to stress that the causal effect can be assessed through other measures, such as for instance the difference between the survival functions
$\theta_{\mathrm{SP}}:=S^{(1)}(\tau)-S^{(0)}(\tau)$
for some time $\tau$, see for instance (Ozenne et al.,, 2020). As mentioned in Section 1.1, another widely used measure (though not necessarily causal) is the hazards ratio, defined as
$\theta_{\mathrm{HR}}:=\frac{\lambda^{(1)}(\tau)}{\lambda^{(0)}(\tau)},$
where the hazard function $\lambda^{(a)}$ is defined as
$\lambda^{(a)}(t):=\lim_{h\to 0^{+}}\frac{\mathbb{P}(T^{(a)}\in[t,t+h]|T^{(a)}\geqslant t)}{h}.$
in a continuous setting, or as $\lambda^{(a)}(t):=\mathbb{P}(T^{(a)}=t|T^{(a)}\geqslant t)$ when the survival times are discrete. The hazard functions and the survival functions are linked through the identities
$S^{(a)}(t)=\exp\left(-\Lambda^{(a)}(t)\right)\quad\text{where}\quad\Lambda^{(a)}(t):=\int_{0}^{t}\lambda^{(a)}(s)\mathop{}\!\mathrm{d}s,$} (2)
in the continuous case. The functions $\Lambda^{(a)}$ are called the cumulative hazard functions. In the discrete case, we have in turn
$S^{(a)}(t)=\prod_{t_{k}\leqslant t}\left(1-\lambda^{(a)}(t_{k})\right),$} (3)
where $\{t_{1},\dots,t_{K}\}$ are the discrete event times of $T^{(a)}$. These hazard functions are classically used to model the survival times and the censoring times, see Section 2.2.1.
1.3 Organization of the paper
In this paper, we detail the minimal theoretical framework that allows the analysis of established RMST estimators in the context of both Randomized Controlled Trials (Section 2) and observational data (Section 3). We give their statistical properties (consistency, asymptotic normality) along with proofs when possible. We then conduct in Section 6 a numerical study of these estimators through simulations under various experimental conditions, including independent and conditionally independent censoring and correct and incorrect model specifications. We conclude in Section 7 with practical recommendations on estimator selection, based on criteria such as asymptotic behavior, computational complexity, and efficiency.
1.4 Notations
We provide in Table LABEL:tbl-reminder_notations a summary of the notation used throughout the paper.
Table 2: Summary of the notations.
Symbol Description
$X$ Covariates
AA Treatment indicator $(A=1$ for treatment, $A=0$ for control))
$T$ Survival time
$T^{(a)},a\in\{0,1\}$ Potential survival time respectively with and without treatment
$S^{(a)},a\in\{0,1\}$ Survival function $S^{(a)}(t)=\mathbb{P}(T^{(a)}>t)$ of the potential survival times
$\lambda^{(a)},a\in\{0,1\}$ Hazard function $\lambda^{(a)}(t)=\lim_{h\to 0^{+}}\mathbb{P}(T^{(a)}\in[t,t+h)|T^{(a)}\geqslant t)/h$ of the potential survival times
$\Lambda^{(a)},a\in\{0,1\}$ Cumulative hazard function of the potential survival times
$C$ Censoring time
$S_{C}$ Survival function $S_{C}(t)=\mathbb{P}(C>t)$ of the censoring time
$G$ Left-limit of the survival function $G(t)=\mathbb{P}(C\geqslant t)$ of the censoring time
$\widetilde{T}$ Observed time ($T\wedge C$)
$\Delta$ Event status $\mathbb{I}\{T\leqslant C\}$
$\Delta^{\tau}$ Event status of the restricted time $\Delta^{\tau}=\max\{\Delta,\mathbb{I}\{\widetilde{T}\geqslant\tau\}\}$
$\{t_{1},t_{2},\dots,t_{K}\}$ Discrete times
$e(x)$ Propensity score $\mathbb{E}[A|X=x]$
$\mu(x,a),a\in\{0,1\}$ $\mathbb{E}[T\wedge\tau\mid X=x,A=a]$
$S(t|x,a),a\in\{0,1\}$ Conditional survival function, $\mathbb{P}(T>t|X=x,A=a)$.
$\lambda^{(a)}(t|x),a\in\{0,1\}$ Conditional hazard functions of the potential survival times
$G(t|x,a),a\in\{0,1\}$ left-limit of the conditional survival function of the censoring $\mathbb{P}(C\geqslant t|X=x,A=a)$
$Q_{S}(t|x,a),a\in\{0,1\}$ $\mathbb{E}[T\wedge\tau\mid X=x,A=a,T\wedge\tau>t]$
2 Causal survival analysis in Randomized Controlled Trials
Randomized Controlled Trials (RCTs) are the gold standard for establishing the effect of a treatment on an outcome, because treatment allocation is controlled through randomization, which ensures (asymptotically) the balance of covariates between treated and controls, and thus avoids problems of confounding between treatment groups. The core assumption in a classical RCT is the random assignment of the treatment (D. B. Rubin 1974).
Assumption 2.1.
(Random Treatment Assignment)
$A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X$
We also assume that there is a positive probability of receiving the treatment, which we rephrase under the following assumption.
Assumption 2.2.
(Trial Positivity)
$0<\mathbb{P}(A=1)<1$
Under Assumptions 1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) and 2.2 (.Causal survival analysis in Randomized Controlled Trials), classical causal identifiability equations can be written to express $\theta_{\mathrm{RMST}}$ without potential outcomes.
$\theta_{\mathrm{RMST}}$ $=\mathbb{E}[T^{(1)}\wedge\tau-T^{(0)}\wedge\tau]$ (4)
$=\mathbb{E}[T^{(1)}\wedge\tau|A=1]-\mathbb{E}[T^{(0)}\wedge\tau|A=0]$ (Random treatment assignment)
$=\mathbb{E}[T\wedge\tau|A=1]-\mathbb{E}[T\wedge\tau|A=0].$ (SUTVA)
However, Equation 4 still depends on $T$, which remains only partially observed due to censoring. To ensure that censoring does not compromise the identifiability of treatment effects, we must impose certain assumptions on the censoring process, standards in survival analysis, namely, independent censoring and conditionally independent censoring. These assumptions lead to different estimation approaches. We focus on two strategies: those that aim to estimate $\mathbb{E}[T\wedge\tau|A=a]$ directly (e.g., through censoring-unbiased transformations, see Section 2.2), and those that first estimate the survival curves to derive RMST via Equation 1 (such as the Kaplan-Meier estimator and all its variants, see the next Section).
2.1 Independent censoring: the Kaplan-Meier estimator
In a first approach, one might assume that the censoring times are independent on the rest of the variables.
Assumption 2.3.
(Independent Censoring)
$C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X,A.$
Under Assumption 2.3, subjects censored at time $t$ are representative of all subjects who remain at risk at time $t$.
Figure 2 represents the causal graph when the study is randomized and outcomes are observed under independent censoring. The treatment assignment $A$ is randomized and hence independent of baseline covariates $X$, which may influence the potential survival outcomes $T^{(a)}$. The observed survival time $T$ depends on both $A$ and $X$, while the censoring time $C$ is assumed to be conditionally independent of the potential outcomes given $X$, reflecting an independent censoring mechanism. This structure encodes key assumptions used throughout this section, including consistency, random treatment assignment, and independent censoring.
Figure 2: Causal graph in RCT survival data with independent censoring.
We also assume that there is no almost-sure upper bound on the censoring time before $\tau$, which we rephrase under the following assumption.
Assumption 2.4.
(Positivity of the Censoring Process)
There exists $\varepsilon>0$ such that
$G(t)\geqslant\varepsilon\quad\text{for all}\quad t\in[0,\tau).$
If indeed it was the case that $\mathbb{P}(C<t)=1$ for some $t<\tau$, then we would not be able to infer anything on the survival function on the interval $t,\tau]$ as all observation times $\widetilde{T}_{i}$ would be in $[0,t]$ almost surely. In practice, adjusting the threshold time $\tau$ can help satisfy the positivity assumption. For instance, in a clinical study, if a subgroup of patients has zero probability of remaining uncensored at a given time, $\tau$ can be modified to ensure that participants have a feasible chance of remaining uncensored up to the revised threshold.
The two Assumptions 2.3 (.Independent censoring: the Kaplan-Meier estimator) and 2.4 (.Independent censoring: the Kaplan-Meier estimator) together allow the distributions of $T^{(a)}$ to be identifiable, in the sense that there exists an identity that expresses $S^{(a)}$ as a function of the joint distribution of $(\widetilde{T},\Delta,A=a)$, see for instance (Ebrahimi et al.,, 2003) for such a formula in a non-causal framework. This enables several estimation strategies, the most well-known being the Kaplan-Meier product-limit estimator.
To motivate the definition of the latter and make explicit the identifiability identity, we set the analysis in the discrete case. We let $\{t_{k}\}_{k\geqslant 1}$ be a set of positive and increasing times and assume that $T\in\{t_{k}\}_{k\geqslant 1}$ almost surely. Then for any $t\in[0,\tau]$, it holds, letting $t_{0}=0$ by convention, thanks to Equation 3,
$S(t|A=a)$ $=\mathbb{P}(T>t|A=a)=\prod_{t_{k}\leqslant t}\left(1-\mathbb{P}(T=t_{k}|T>t_{k-1},A=a)\right)$
$=\prod_{t_{k}\leqslant t}\left(1-\frac{\mathbb{P}(T=t_{k},A=a)}{\mathbb{P}(T\geqslant t_{k},A=a)}\right).$
Using Assumptions 2.3 (.Independent censoring: the Kaplan-Meier estimator) and 2.4 (.Independent censoring: the Kaplan-Meier estimator), we find that
$\frac{\mathbb{P}(T=t_{k},A=a)}{\mathbb{P}(T\geqslant t_{k},A=a)}=\frac{\mathbb{P}(T=t_{k},C\geqslant t_{k},A=a)}{\mathbb{P}(T\geqslant t_{k},C\geqslant t_{k},A=a)}=\frac{\mathbb{P}(\widetilde{T}=t_{k},\Delta=1,A=a)}{\mathbb{P}(\widetilde{T}\geqslant t_{k},A=a)},$} (5)
yielding the final identity
$S(t|A=a)=\prod_{t_{k}\leqslant t}\left(1-\frac{\mathbb{P}(\widetilde{T}=t_{k},\Delta=1,A=a)}{\mathbb{P}(\widetilde{T}\geqslant t_{k},A=a)}\right).$} (6)
Notice that the right hand side only depends on the distribution of the observed tuple $(A,\widetilde{T},\Delta)$. This last equation suggests in turn to introduce the quantities
$D_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A=a)\quad\text{and}\quad N_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A=a),$} (7)
which correspond respectively to the number of deaths $D_{k}(a)$ and of individuals at risk $N_{k}(a)$ at time $t_{k}$ in the treated group (a=1) or in the control group (a=0).
Definition 2.1.
(Kaplan-Meier estimator, (Kaplan and Meier,, 1958)) With $D_{k}(a)$ and $N_{k}(a)$ defined in Equation 7, we let
$\widehat{S}_{\mathrm{KM}}(t|A=a):=\prod_{t_{k}\leqslant t}\left(1-\frac{D_{k}(a)}{N_{k}(a)}\right).$} (8)
The associated RMST estimator is then simply defined as
$\widehat{\theta}_{\mathrm{KM}}=\int_{0}^{\tau}\widehat{S}_{KM}(t|A=1)-\widehat{S}_{KM}(t|A=0)dt.$} (9)
The Kaplan-Meier estimator is the Maximum Likelihood Estimator (MLE) of the survival functions, see for instance Kaplan and Meier, (1958). Furthermore, because $D_{k}(a)$ and $N_{k}(a)$ are sums of i.i.d. random variables, the Kaplan-Meier estimator inherits some convenient statistical properties. In Annex 9 (Section 9.1) we show that it exhibits exponentially fast bias decay (Proposition 9.1), asymptotic normality with explicitly characterized variances (Proposition 9.2). We finally mention that the KM estimator as defined in Definition 2.1 still makes sense in a non-discrete setting, and one only needs to replace the fixed grid $\{t_{k}\}$ by the values at which we observed an event $(\widetilde{T}_{i}=t_{k},\Delta_{i}=1)$. We refer to Breslow, (1974) for a study of this estimator in the continuous case and to Aalen et al., (2008), Sec 3.2 for a general study of the KM estimator through the prism of point processes.
2.2 Conditionally independent censoring
An alternative hypothesis in survival analysis that relaxes the assumption of independent censoring is conditionally independent censoring, also referred sometimes as informative censoring. It allows to model more realistic censoring processes, in particular in situations where there are reasons to believe that $C$ may be dependent on $A$ and $X$ (for instance, if a patient is more likely to leave the study when treated because of side-effects of the treatment).
Assumption 2.5.
(Conditionally Independent Censoring)
$C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X,A$
Under Assumption 2.5, subjects within a same stratum defined by $X=x$ and $A=a$ have equal probability of censoring at time $t$, for all $t$. In case of conditionally independent censoring, we also need to assume that all subjects have a positive probability to remain uncensored at their time-to-event.
Assumption 2.6.
(Positivity / Overlap for Censoring)
There exists $\varepsilon>0$ such that for all $t\in[0,\tau]$, it almost surely holds
$G(t|A,X)\geqslant\varepsilon.$
Figure 3 represents the causal graph when the study is randomized with conditionally independent censoring.
Figure 3: Causal graph in RCT survival data with dependent censoring.
Under dependent censoring, the Kaplan-Meier estimator as defined in Definition 2.1 can fail to estimate survival, in particular because Equation 5 no longer holds. Alternatives include plug-in G-formula estimators (Section 2.2.1) and unbiased transformations (Section 2.2.2).
2.2.1 The G-formula and the Cox Model
Because the censoring is now independent on the potential outcome conditionally to the covariates, it would seem natural to model the response of the survival time conditionally to these covariates too:
$\mu(x,a):=\mathbb{E}[T\wedge\tau|X=x,A=a].$
Building on Equation 4, one can express the RMST as a function of $\mu$:
$\theta_{\mathrm{RMST}}=\mathbb{E}\left[\mathbb{E}[T\wedge\tau|X,A=1]\right]-\mathbb{E}\left[T\wedge\tau|X,A=0]\right]=\mathbb{E}[\mu(X,1)-\mu(X,0)].$
An estimator $\widehat{\mu}$ of $\mu$ would then straightforwardly yields an estimator of the RMST difference through the so-called G-formula plug-in estimator:
$\widehat{\theta}_{\mathrm{G-formula}}=\frac{1}{n}\sum_{i=1}^{n}\widehat{\mu}\left(X_{i},1\right)-\widehat{\mu}\left(X_{i},0\right).$} (10)
We would like to stress that a G-formula approach works also in a observational context as the one introduced in Section 3.2. However, because the estimation strategies presented in the next sections rely on estimating nuisance parameters, and that this latter task is often done in the same way as we estimate the conditional response $\mu$, we decided to not delay the introduction of the G-formula any further, and we present below a few estimation methods for $\mu$. These methods are subdivided in two categories: $T$-learners (i.e. two learners), where $\mu(\cdot,1)$ is estimated separately from $\mu(\cdot,0)$, and $S$-learner (i.e. single learner), where $\widehat{\mu}$ is obtained by fitting a single model based on covariates $(X,A)$.
Cox Model. There are many ways to model $\mu$ in a survival context, the most notorious of which being the Cox proportional hazards model (Cox,, 1972). It relies on a semi-parametric model of the conditional hazard functions and specifies the function under treatment $a$, $\lambda^{(a)}(t|X)$, as
$\lambda^{(a)}(t|X)=\lambda_{0}^{(a)}(t)\exp(X^{\top}\beta^{(a)}),$
where $\lambda^{(a)}_{0}$ is a baseline hazard function under treatment $a$ and $\beta^{(a)}$ has the same dimension as the vector of covariate $X$ (We use the superscript $(a)$ here to emphasize that a separate Cox model is fit for each treatment level, since the treatment assignment $A$ is not included as a covariate in $X$). The conditional survival function then takes the simple form (in the continuous case)
$S^{(a)}(t|X)=S^{(a)}_{0}(t)^{\exp(X^{\top}\beta^{(a)})},$
where $S^{(a)}_{0}(t)$ is the survival function associated with $\lambda_{0}^{(a)}$. The vector $\beta^{(a)}$ is classically estimated by maximizing the so-called partial likelihood function as introduced in the original paper of Cox, (1972). The cumulative baseline hazard function can be estimated through the Breslow’s estimator (Breslow,, 1974):
$\widehat{\Lambda}_{0}^{(a)}(t)=\sum_{\Delta_{i}=1,\widetilde{T}_{i}\leqslant t}\frac{1}{\displaystyle\sum_{\widetilde{T}_{j}\geqslant\widetilde{T}_{i}}\exp(X_{j}^{\top}\widehat{\beta}^{(a)})}$
where $\widehat{\beta}^{(a)}$ is a partial likelihood maximizer. One can show that $(\widehat{\beta}^{(a)},\widehat{\Lambda}_{0}^{(a)})$ is the MLE of the true likelihood, when $\widehat{\Lambda}_{0}^{(a)}$ is optimized over all step functions of the form
$\Lambda_{0}(t):=\sum_{\Delta_{i}=1}h_{i},\quad h_{i}\in\mathbb{R}^{+}.$
This fact was already hinted in the original paper by Cox and made rigorous in many subsequent papers, see for instance Fan et al., (2010). Furthermore, if the true distribution follows a Cox model, then both $\widehat{\beta}^{(a)}$ and $\widehat{\Lambda}_{0}^{(a)}$ are strongly consistent and asymptotically normal estimator of the true parameters $\beta^{(a)}$ and $\Lambda^{(a)}$, see Kalbfleisch and Prentice, (2002), Sec 5.7. When using a $T$-learner approach, one simply finds $(\widehat{\beta}^{(a)},\widehat{\Lambda}_{0}^{(a)})$ for $a\in\{0,1\}$ by considering the control group and the treated group separately. When using a $S$-learner approach, the treatment status $A$ becomes a covariate a the model becomes
$\lambda(t|X,A)=\lambda_{0}(t)\exp(X^{\top}\beta+\alpha A).$} (11)
for some $\alpha\in\mathbb{R}$. One main advantage of Cox model is that it makes it very easy to interpret the effect of a covariate on the survival time. If indeed $\alpha>0$, then the treatment has a negative effect on the survival times. Likewise, if $\beta_{i}>0$, then the $i$-th coordinate of $X$ has a negative effect as well. We finally mention that the hazard ratio takes a particularly simple form under the later model since
$\theta_{\mathrm{HR}}=e^{\alpha}.$
In particular, it does not depends on the time horizon $\tau$, and is thus sometimes referred to as proportional hazard. Figure 4 illustrates the estimation of the difference in Restricted Mean Survival Time using G-formula with Cox models.
Finally, while we have focused here on Cox-based specifications due to its widespread use in biomedical research. However, alternative approaches exist. For example, Accelerated Failure Time (AFT) models (Saikia and Barman,, 2017) provide a semi-parametric alternative that directly models the survival time and does not rely on the proportional hazards (PH) assumption.
Figure 4: Illustration of the G-formula for estimating $\theta_{\mathrm{RMST}}$ in an RCT when only one covariate $X_{1}$ influences the outcome.
Weibull Model. A popular parametric model for survival is the Weibull Model, which amounts to assume that
$\lambda^{(a)}(t|X)=\lambda_{0}^{(a)}(t)\exp(X^{\top}\beta)$
where $\lambda_{0}^{(a)}(t)$ is the instantaneous hazard function of a Weibull distribution, that is to say a function proportional to $t^{\gamma}$ for some shape parameter $\gamma>0$. We refer to Zhang, (2016) for a study of this model.
Survival Forests. On the non-parametric front, we mention the existence of survival forests (Ishwaran et al.,, 2008).
2.2.2 Censoring-unbiased transformations
Censoring-unbiased transformations involve applying a transformation to $T$. Specifically, we compute a new time $T^{*}$ of the form
$T^{*}:=T^{*}(\widetilde{T},X,A,\Delta)=\begin{cases}\phi_{0}(\widetilde{T}\wedge\tau,X,A)\quad&\text{if}\quad\Delta^{\tau}=0,\\ \phi_{1}(\widetilde{T}\wedge\tau,X,A)\quad&\text{if}\quad\Delta^{\tau}=1.\end{cases}$} (12)
for two wisely chosen transformations $\phi_{0}$ and $\phi_{1}$, and where
$\Delta^{\tau}:=\mathbb{I}\{T\wedge\tau\leqslant C\}=\Delta+(1-\Delta)\mathbb{I}(\widetilde{T}\geqslant\tau)$} (13)
is the indicator of the event where the individual is either uncensored or censored after time $\tau$. The idea behind the introduction of $\Delta^{\tau}$ is that because we are only interested in computing the expectation of the survival time thresholded by $\tau$, any censored observation coming after time $\tau$ can in fact be considered as uncensored ($\Delta^{\tau}=1$).
A censoring-unbiased transformation $T^{*}$ shall satisfy: for $a\in\{0,1\}$, it holds
$\mathbb{E}[T^{*}|A=a,X]=\mathbb{E}[T^{(a)}\wedge\tau|X]\quad\text{almost surely.}$} (14)
A notable advantage of this approach is that it enables the use of the full transformed dataset $(X_{i},A_{i},T^{*}_{i})$ as if no censoring occurred. Because it holds
$\mathbb{E}[\mathbb{E}[T^{*}|A=a,X]]=\mathbb{E}\left[\frac{\mathbb{I}\{A=a\}}{\mathbb{P}(A=a)}T^{*}\right],$} (15)
One can derive an estimator of the RMST difference from any censoring-unbiased transformation $T^{*}$ as:
$\widehat{\theta}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{i}$} (16)
where $\pi=\mathbb{P}(A=1)\in(0,1)$ by Assumption 2.2 and where $T^{*}_{i}=T^{*}(\widetilde{T}_{i},X_{i},A_{i},\Delta_{i})$.
The following proposition establishes the theoretical guarantees for the class of estimators constructed from bounded censoring-unbiased transformations:
Proposition 2.1.
Under Assumptions 2.1 (.Causal survival analysis in Randomized Controlled Trials) and 2.2 (.Causal survival analysis in Randomized Controlled Trials), the estimator $\widehat{\theta}$ derived as in Equation 16 from square integrable censoring-unbiased transformations satisfying Equation 14 is an unbiased, strongly consistent, and asymptotically normal estimator of the RMST difference.
Square integrability will be ensured any time the transformation is bounded, which will always be the case of the ones considered in this work.
It is natural in a RCT setting to assume that probability of being treated $\pi$ is known. If not, it is usual to replace $\pi$ by its empirical counterpart $\widehat{\pi}=n_{1}/n$ where $n_{a}=\sum_{i}\mathds{1}\{A=a\}$. The resulting estimator takes the form
$\widehat{\theta}=\frac{1}{n_{1}}\sum_{A_{i}=1}T^{*}_{i}-\frac{1}{n_{0}}\sum_{A_{i}=0}T^{*}_{i}.$} (17)
While this estimator introduces a slight bias due to the estimation of $\pi$ (see for instance Colnet et al., (2022), Lemma 2), it remains strongly consistent and asymptotically normal under the same assumptions:
Proposition 2.2.
Under Assumptions 2.1 (.Causal survival analysis in Randomized Controlled Trials) and 2.2 (.Causal survival analysis in Randomized Controlled Trials), the estimator $\widehat{\theta}$ derived as in Equation 17 from a square integrable censoring-unbiased transformations satisfying Equation 14 is a strongly consistent, and asymptotically normal estimator of the RMST difference.
The two most popular transformations are Inverse-Probability-of-Censoring Weighting (Koul et al.,, 1981) and Buckley-James (Buckley and James,, 1979), both illustrated in Figure 5 and detailed below. In the former, only non-censored observations are considered and they are weighted while in the latter, censored observations are imputed with an estimated survival time.
Figure 5: Illustration of Inverse-Probability-of-Censoring and Buckley-James transformations.
The Inverse-Probability-of-Censoring Weighted transformation
The Inverse-Probability-of-Censoring Weighted (IPCW) transformation, introduced by Koul et al., (1981) in the context of censored linear regression, involves discarding censored observations and applying weights to uncensored data. More precisely, we let
$T^{*}_{\mathrm{IPCW}}=\frac{\Delta^{\tau}}{G(\widetilde{T}\wedge\tau|X,A)}\widetilde{T}\wedge\tau,$} (18)
where we recall that $G(t|X,A):=\mathbb{P}(C\geqslant t|X,A)$ is the left limit of the conditional survival function of the censoring. This estimator assigns higher weights to uncensored subjects within a covariate group that is highly prone to censoring, thereby correcting for conditionally independent censoring and reducing selection bias (Howe et al.,, 2016).
We establish in Appendix 9, Section 9.2, (Proposition 9.3) that the IPCW transformation is indeed a censoring-unbiased transformation in the sense of Equation 14.
The IPCW depends on the unknown conditional survival function of the censoring $G(\cdot|X,A)$, which thus needs to be estimated. Estimating conditional censoring or the conditional survival function can be approached similarly, as both involve estimating a time—whether for survival or censoring. Consequently, we can use semi-parametric methods, such as the Cox model, or non-parametric approaches like survival forests, and we refer to Section 2.2.1 for a development on these approaches. Once an estimator $\widehat{G}(\cdot|A,X)$ of the later is provided, one can construct an estimator of the difference in RMST based on Equation 16 or Equation 17
$\widehat{\theta}_{\mathrm{IPCW}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{\mathrm{IPCW},i},$} (19)
or
$\widehat{\theta}_{\mathrm{IPCW}}=\frac{1}{n_{1}}\sum_{A_{i}=1}T^{*}_{\mathrm{IPCW},i}-\frac{1}{n_{0}}\sum_{A_{i}=0}T^{*}_{\mathrm{IPCW},i}.$} (20)
where we recall that $n_{a}:=\#\{i\in[n]~|~A_{i}=a\}$. By Proposition 2.1, Proposition 2.2 and Proposition 9.3, we easily deduce that $\widehat{\theta}_{\mathrm{IPCW}}$ is asymptotically consistent as soon as $\widehat{G}$ is.
Corollary 2.1.
Under Assumptions1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.2 (.Causal survival analysis in Randomized Controlled Trials), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{G}$ is uniformly weakly (resp. strongly) consistent then so is $\widehat{\theta}_{\mathrm{IPCW}}$, either as in defined in Equation 19 or in Equation 20.
This result simply comes from the fact that $\widehat{\theta}_{\mathrm{IPCW}}$ depends continuously on $\widehat{G}$ and that $G$ is lower-bounded (Assumption 2.6 .Conditionally independent censoring). Surprisingly, we found limited use of this estimator in the literature, beside its first introduction in Koul et al., (1981). This could potentially be explained by the fact that, empirically, we observed that this estimator is highly variable. Consequently, we do not explore its properties further and will not include it in the numerical experiments. A related and more popular estimator is the IPCW-Kaplan-Meier, defined as follows.
Definition 2.2.
(IPCW-Kaplan-Meier) We let again $\widehat{G}(\cdot|X,A)$ be an estimator of the (left limit of) the conditional censoring survival function and we introduce
$D_{k}^{\mathrm{IPCW}}(a)$ $:=\sum_{i=1}^{n}\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
$\quad\text{and}\quad N^{\mathrm{IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
be the weight-corrected numbers of deaths and of individuals at risk at time $t_{k}$. The IPCW version of the KM estimator is defined as:
$\widehat{S}_{\mathrm{IPCW}}(t|A=a)$ $=\prod_{t_{k}\leqslant t}\left(1-\frac{D_{k}^{\mathrm{IPCW}}(a)}{N_{k}^{\mathrm{IPCW}}(a)}\right).$
Note that the quantity $\pi$ is not present in the definition of $D_{k}^{\mathrm{IPCW}}(a)$ and $N_{k}^{\mathrm{IPCW}}(a)$ because it would simply disappear in the ratio $D_{k}^{\mathrm{IPCW}}(a)/N_{k}^{\mathrm{IPCW}}(a)$. The subsequent RMST estimator then takes the form
$\widehat{\theta}_{\mathrm{IPCW-KM}}=\int_{0}^{\tau}\widehat{S}_{\mathrm{IPCW}}(t|A=1)-\widehat{S}_{\mathrm{IPCW}}(t|A=0)dt.$} (21)
Like before for the classical KM estimator, this new reweighted KM estimator enjoys good statistical properties.
We show in Proposition 9.4 (Appendix 9) that the oracle version of this estimator is strongly consistent and asymptotically normal. These properties also hold for the estimator using a consistently estimated $\widehat{G}$.
Corollary 2.2.
Under Assumptions1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{G}$ is uniformly weakly (resp. strongly) consistent then so is $\widehat{S}_{\mathrm{IPCW}}(t|A=a)$.
Because the evaluation of $N_{k}^{\textrm{IPCW}}(a)$ now depends on information gathered after time $t_{k}$ (through the computation of the weights), the previous properties on the absence of bias (Proposition 9.1) and on the derivation of the asymptotic variance (Proposition 9.2) do not hold in this setting. Whether its bias is exponentially small and whether the asymptotic variance can be derived in a closed form are questions left open. We are also not aware of any estimation schemes for the asymptotic variance in this context. In the case where we do not have access to the oracle survival function $G$, we can again still achieve consistency if the estimator $\widehat{G}(\cdot|A,X)$ that we provide is consistent.
The Buckley-James transformation
One weakness of the IPCW transformation is that it discards all censored data. The Buckley-James (BJ) transformation, introduced by Buckley and James, (1979), takes a different path by leaving all uncensored values untouched, while replacing the censored ones by an extrapolated value. Formally, it is defined as follows:
$T^{*}_{\mathrm{BJ}}$ $=\Delta^{\tau}(\widetilde{T}\wedge\tau)+(1-\Delta^{\tau})Q_{S}(\widetilde{T}\wedge\tau|X,A),$ (22)
where, for $t\leqslant\tau$,
$Q_{S}(t|X,A):=\mathbb{E}[T\wedge\tau|X,A,T\wedge\tau>t]=\frac{1}{S(t|X,A)}\int_{t}^{\tau}S(u|X,A)\mathop{}\!\mathrm{d}u$
where $S(t|X,A=a):=\mathbb{P}(T^{(a)}>t|X)$ is the conditional survival function. We refer again to Figure 5 for a diagram of this transformation.
We show in Appendix 9 (Proposition 9.5) that this transformation satisfies the censoring-unbiasedness condition defined in Equation 14.
Again, the BJ transformation depends on a nuisance parameter (here $Q_{S}(\cdot|X,A)$) that needs to be estimated. We again refer to Section 2.2.1 for a brief overview of possible estimation strategies for $Q_{S}$. Once provided with an estimator $\widehat{Q}_{S}(\cdot|A,X)$, a very natural estimator of the RMST based on the BJ transformation and on Equation 16 or Equation 17 would be
$\widehat{\theta}_{\mathrm{BJ}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{\mathrm{BJ},i},$} (23)
or
$\widehat{\theta}_{\mathrm{BJ}}=\frac{1}{n_{1}}\sum_{A_{i}=1}T^{*}_{\mathrm{BJ},i}-\frac{1}{n_{0}}\sum_{A_{i}=0}T^{*}_{\mathrm{BJ},i}.$} (24)
Like for the IPCW transformation, the BJ transformation yields a consistent estimate of the RMST as soon as the model is well-specified.
Corollary 2.3.
Under Assumptions1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{Q}_{S}$ is uniformly weakly (resp. strongly) consistent then so is $\widehat{\theta}_{\mathrm{BJ}}$ defined as in Equation 23 or Equation 24.
The proof is again a mere application of Propositions 2.1, 2.2 and 9.5, and relies on the continuity of $S\mapsto Q_{S}$.
Additionally, the Buckley-James transformation enjoys a notable optimality property: it minimizes the mean squared error among all square-integrable censoring-unbiased transformations. We restate formally as Theorem 9.1 in Appendix 9. This result, originally stated without proof in Fan and Gijbels, (1994), is proved there for completeness.
2.2.3 Augmented corrections
The main disadvantage of the two previous transformations is that they heavily rely on the specification of good estimator $\widehat{G}$ (for IPCW) or $\widehat{S}$ (for BJ). In order to circumvent this limitation, Rubin and van der Laan, (2007) proposed the following transformations, whose expression is based on theory of semi-parametric estimation developed in Van der Laan and Robins, (2003),
$T^{*}_{\mathrm{DR}}=\frac{\Delta^{\tau}\widetilde{T}\wedge\tau}{G(\widetilde{T}\wedge\tau|X,A)}+\frac{(1-\Delta^{\tau})Q_{S}(\widetilde{T}\wedge\tau|X,A)}{G(\widetilde{T}\wedge\tau|X,A)}-\int_{0}^{\widetilde{T}\wedge\tau}\frac{Q_{S}(t|X,A)}{G(t|X,A)^{2}}\mathop{}\!\mathrm{d}\mathbb{P}_{C}(t|X,A),$} (25)
where $\mathop{}\!\mathrm{d}\mathbb{P}_{C}(t|X,A)$ is the distribution of $C$ conditional on the covariates $X$ and $A$. We stress that this distribution is entirely determined by the $G(\cdot|X,A)$, so that this transformation only depends on the knowledge of both conditional survival functions $G$ and $S$, will be thus sometimes denoted $T^{*}_{\mathrm{DR}}(G,S)$ to underline this dependency. Importantly, this transformation not only satisfies the censoring-unbiasedness condition of Equation 14, but also enjoys a double robustness property: it remains valid as long as either $G$ or $S$ is correctly specified. The formal statement of this result is given in Proposition 9.6 in Appendix 9. The statement and proof of this results is found in Rubin and van der Laan, (2007) in the case where $C$ and $T$ are continuous.
3 Causal survival analysis in observational studies
Unlike RCT, observational data — such as from registries, electronic health records, or national healthcare systems — are collected without controlled randomized treatment allocation. In such cases, treated and control groups are likely unbalanced due to the non-randomized design, which results in the treatment effect being confounded by variables influencing both the survival outcome $T$ and the treatment allocation $A$. To enable identifiability of the causal effect, additional standard assumptions are needed.
Assumption 3.1.
(Conditional Exchangeability / Unconfoundedness)
It holds
$A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X$
Under Assumption 3.1, the treatment assignment is randomly assigned conditionally on the covariates $X$. This assumption ensures that there are no unmeasured confounders as the latter would make it impossible to distinguish correlation from causality.
Assumption 3.2.
(Positivity / Overlap for Treatment)
Letting $e(X):=\mathbb{P}(A=1\mid X)$ be the propensity score, there holds
$0<e(X)<1\quad\text{almost surely.}$
The Assumption 3.2 requires adequate overlap in covariate distributions between treatment groups, meaning every observation must have a non-zero probability of being treated. Because Assumption 2.1 (.Causal survival analysis in Randomized Controlled Trials) no longer holds anymore, neither does the previous identifiability Equation 4. In this new context, we can write
$\theta_{\mathrm{RMST}}$ $=\mathbb{E}[T^{(1)}\wedge\tau-T^{(0)}\wedge\tau]$ (26)
$=\mathbb{E}\left[\mathbb{E}[T^{(1)}\wedge\tau|X]-\mathbb{E}[T^{(0)}\wedge\tau|X]\right]$
$=\mathbb{E}\left[\mathbb{E}[T^{(1)}\wedge\tau|X,A=1]-\mathbb{E}[T^{(0)}\wedge\tau|X,A=0]\right]$ (unconfoundness)
$=\mathbb{E}\left[\mathbb{E}[T\wedge\tau|X,A=1]-\mathbb{E}[T\wedge\tau|X,A=0]\right].$ (SUTVA)
In another direction, one could wish to identify the treatment effect through the survival curve as in Equation 1:
$S^{(a)}(t)=\mathbb{P}(T^{(a)}>t)=\mathbb{E}\left[\mathbb{P}(T>t|X,A=a)\right].$} (27)
Again, both identities still depend on the unknown quantity $T$ and suggest two different estimation strategies. These strategies differ according to the censoring assumptions and are detailed below.
3.1 Independent censoring
Figure 6 illustrates a causal graph in observational survival data with independent censoring (Assumptions 3.1 and 2.3).
Figure 6: Causal graph in observational survival data with independent censoring.
Under Assumption 2.3 (.Independent censoring: the Kaplan-Meier estimator), we saw in Section 2.1 that the Kaplan-Meier estimator could conveniently handle censoring. Building on Equation 27, we can write
$S^{(1)}(t)=\mathbb{E}\left[\frac{\mathbb{E}[\mathbb{I}\{A=1,T>t\}|X]}{\mathbb{E}[\mathbb{I}\{A=1\}|X]}\right]=\mathbb{E}\left[\frac{A\mathbb{I}\{T>t\}}{e(X)}\right],$
which suggests to adapt the classical KM estimator by reweighting it by the propensity score. The use of propensity score in causal inference has been initially introduced by Rosenbaum and Rubin, (1983) and further developed in Hirano et al., (2003). It was extended to survival analysis by Xie and Liu, (2005) through the adjusted Kaplan-Meier estimator as defined below.
Definition 3.1.
[…]
$D_{k}^{\mathrm{IPTW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{a}{\widehat{e}(X_{i})}+\frac{1-a}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A_{i}=a)$
$\quad\text{and}\quad N^{\mathrm{IPTW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{a}{\widehat{e}(X_{i})}+\frac{1-a}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
be the numbers of deaths and of individuals at risk at time $t_{k}$, reweighted by the propensity score. The Inverse-Probability-of-Treatment Weighting (IPTW) version of the KM estimator is defined as:
$\widehat{S}_{\mathrm{IPTW}}(t|A=a)$ $=\prod_{t_{k}\leqslant t}\left(1-\frac{D_{k}^{\mathrm{IPTW}}(a)}{N_{k}^{\mathrm{IPTW}}(a)}\right).$ (28)
We let $S^{*}_{\mathrm{IPTW}}(t|A=a)$ be the oracle KM-estimator defined as above with $\widehat{e}(\cdot)=e(\cdot)$. Although the reweighting slightly changes the analysis, the good properties of the classical KM carry on to this setting.
In Appendix 9, Proposition 9.7 establishes that $S^{*}_{\mathrm{IPTW}}(t\mid A=a)$ is a strongly consistent and asymptotically normal estimator of the survival function $S^{(a)}(t)$. Corollary 3.1 further shows that the same properties hold for the practical estimator provided the estimated propensity score converges uniformly. This follow directly from $S^{*}_{\mathrm{IPTW}}$ is a continuous function of $e(\cdot)$, and because $e$ and 1e1-e are lower-bounded as per Assumptions 3.2 (.Causal survival analysis in observational studies).
Corollary 3.1.
[…]
The resulting RMST estimator simply takes the form:
$\widehat{\theta}_{\mathrm{IPTW-KM}}=\int_{0}^{\tau}\widehat{S}_{\mathrm{IPTW}}(t|A=1)-\widehat{S}_{\mathrm{IPTW}}(t|A=0)dt.$} (29)
which will be consistent under the same Assumptions as the previous Corollary. Note that we are not aware of any formal results concerning the bias and the asymptotic variance of the oracle estimator $S^{*}_{\mathrm{IPTW}}(t|A=a)$, and we refer to Xie and Liu, (2005) for heuristics concerning these questions.
3.2 Conditional independent censoring
Under Assumptions 3.1 (.Causal survival analysis in observational studies) and 2.5 (.Conditionally independent censoring), the causal effect is affected both by confounding variables and by conditional censoring. The associated causal graph is depicted in Figure 7. In this setting, one can still use the $G$-formula exactly as in Section 2.2.1.
A natural alternative approach is to weight the IPCW and BJ transformations from Section 2.2.2 by the propensity score to disentangle both confounding effects and censoring at the same time.
Figure 7: Causal graph in observational survival data with dependent censoring.
We mention that the G-formula approach developed in Section 2.2.1 does work in that context. In particular, Chen and Tsiatis, (2001) prove consistency and asymptotic normality results for Cox estimators in an observational study, and they give an explicit formulation of the asymptotic variance as a function of the parameters of the Cox model. In the non-parametric world, Foster et al., (2011) and Künzel et al., (2019) empirically study this estimator using survival forests, with the former employing it as a T-learner (referred to as Virtual Twins) and the latter as an S-learner.
3.2.1 IPTW-IPCW transformations
One can check that the IPCW transformation as introduced in Equation 18 is also a censoring-unbiased transformation in that context (formally stated in Proposition 9.8). Deriving an estimator of the difference in RMST is however slightly different in that context. In particular, Equation 15 rewrites
$\mathbb{E}[\mathbb{E}[T^{*}|X,A=1]]=\mathbb{E}\left[\frac{A}{e(X)}T^{*}\right],$
Which in turn suggests to define
$\widehat{\theta}_{\mathrm{IPTW-IPCW}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)T^{*}_{\mathrm{IPCW},i}.$} (30)
This transformation now depends on two nuisance parameters, namely the conditional survival function of the censoring (through $T^{*}_{\mathrm{IPCW}}$) and the propensity score. Proposition 3.1 ensures consistency of this estimator when both nuisance quantities are consistently estimated.
Proposition 3.1.
[…]
We can also use the same strategy as for the IPCW transformation and incorporate the new weights into a Kaplan-Meier estimator.
Definition 3.2.
[…]
$D_{k}^{\mathrm{IPTW-IPCW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
$\quad\text{and}\quad N^{\mathrm{IPTW-IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
be the weight-corrected numbers of deaths and of individual at risk at time $t_{k}$. The IPTW-IPCW version of the KM estimator is defined as:
$\widehat{S}_{\mathrm{IPTW-IPCW}}(t|A=a)$ $=\prod_{t_{k}\leqslant t}\left(1-\frac{D_{k}^{\mathrm{IPTW-IPCW}}(a)}{N_{k}^{\mathrm{IPTW-IPCW}}(a)}\right).$
The RMST difference estimated with IPTW-IPCW-Kaplan-Meier survival curves is then simply as
$\widehat{\theta}_{\mathrm{IPTW-IPCW-KM}}=\int_{0}^{\tau}\widehat{S}_{\mathrm{IPTW-IPCW}}(t|A=1)-\widehat{S}_{\mathrm{IPTW-IPCW}}(t|A=0)dt.$} (31)
In Appendix 9, Proposition 9.9 shows that the oracle version of this estimator is consistent and asymptotically normal. The result extends to the practical estimator: as shown in the corollary below (Corollary 3.2), uniform consistency of the nuisance estimators ensures the consistency of the estimated survival curves $\widehat{\theta}_{\mathrm{IPTW-IPCW-KM}}$.
Corollary 3.2.
Under Assumptions 1.1 (.Definition of the estimand: the RMST), 3.1 (.Causal survival analysis in observational studies), 3.2 (.Causal survival analysis in observational studies), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), and for all $t\in[0,\tau]$, if the nuisance estimators $\widehat{G}(\cdot|A,X)$ and $\widehat{e}$ are weakly (resp. strongly) uniformly consistent, then $\widehat{S}_{\mathrm{IPTW-IPCW}}(t|A=a)$ is a weakly (resp. strongly) consistent estimator of $S^{(a)}(t)$.
We are not aware of general formulas for the asymptotic variances in this context. We mention nonetheless that Schaubel and Wei, (2011) have been able to derive asymptotic laws in this framework in the particular case of Cox-models.
3.2.2 IPTW-BJ transformations
Like IPCW transformation, BJ transformation is again a censoring-unbiased transformation in an observational context. This result is formalized in Proposition 9.10 in Appendix 9.
The corresponding estimator of the difference in RMST is
$\hat{\theta}_{\mathrm{IPTW-BJ}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)T^{*}_{\mathrm{BJ},i}.$} (32)
This transformation depends on the conditional survival function $S$ (through $T^{*}_{\mathrm{BJ}}$) and the propensity score. Proposition 3.2 stated that consistency of the nuisance parameter estimators implies again consistency of the RMST estimator.
Proposition 3.2.
[…]
3.2.3 Double augmented corrections
Building on the classical doubly-robust AIPTW estimator from causal inference (Robins et al.,, 1994), we could incorporate the doubly-robust transformations for censoring of Section 2.2.3 to obtain a doubly robust transformation for censoring and treatment assignment
$\Delta^{*}_{\mathrm{QR}}=\Delta^{*}_{\mathrm{QR}}(G,S,\mu,e):=\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)(T^{*}_{\mathrm{DR}}(G,S)-\mu(X,A))+\mu(X,1)-\mu(X,0),$
where we recall that $T^{*}_{\mathrm{DR}}$ is defined in Section 2.2.3. This transformation depends on four nuisance parameters: $G$ and $S$ through $T^{*}_{\mathrm{DR}}$, and now the propensity score $e$ and the conditional response $\mu$. This transformation does not really fall into the scope of censoring-unbiased transform, but it is easy to show that $\Delta^{*}_{\mathrm{QR}}$ is doubly robust for censoring and treatment assignment. That is, among the four nuisance functions involved, only one from each pair $(S,G)$ and $(e,\mu)$ needs to be correctly specified. This property (similar to Ozenne et al., (2020), Thm 1) is formally stated in Proposition 9.11 in Appendix 9.
[…]
$=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}-\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)(T^{*}_{\mathrm{DR}}(\widehat{G},\widehat{S})_{i}-\widehat{\mu}(X_{i},A_{i}))+\widehat{\mu}(X_{i},1)-\widehat{\mu}(X_{i},0).$
This estimator enjoys good asymptotic properties under parametric models, as detailed in Ozenne et al., (2020)
4 Variance estimation
Some survival estimators presented in this work have known asymptotic variance expressions, such as the Kaplan-Meier estimator (via Greenwood’s formula) or weighted Kaplan-Meier estimators. Under assumptions of asymptotic normality and independence between the two survival curves, the variance of the RMST difference can then be straightforwardly obtained through the delta method.
However, in practical situations, the covariance between estimated survival curves is typically nonzero, complicating variance estimation. While explicit variance formulas have been derived in certain settings—such as for the G-formula combined with semiparametric estimators (see Chen and Tsiatis, (2001))—these expressions often remain complex and difficult to implement. Therefore, we recommend using the nonparametric bootstrap (Efron and Tibshirani,, 1994) as a robust and practical solution for variance estimation and confidence interval construction in general settings. The effectiveness of bootstrap methods in the context of censored data has been investigated in several studies, including Efron, (1981) and Chen and Lo, (1996), which provide both theoretical justification and empirical validation for its use with Kaplan–Meier estimator.
5 Implementation
In this section, we first summarize the various estimators and their properties. We then provide custom implementations for all estimators, even those available in existing packages. These manual implementations serve two purposes: first, to make the methods accessible to the community when no existing implementation is available; and second, to facilitate a deeper understanding of the methods by detailing each step, even when a package solution exists. Finally, we present the packages available for directly computing $\theta_{\mathrm{RMST}}$.
5.1 Summary of the estimators
Table LABEL:tbl-nuisance provides an overview of the estimators introduced in this paper, along with the corresponding nuisance parameters needed for their estimation and an overview of their statistical properties in particular regarding their sensitivity to misspecification of the nuisance parameters.
Table 3: Estimators of the RMST difference and nuisance parameters needed to compute each estimator. Empty boxes indicate that the nuisance parameter is not needed in the estimator thus misspecification has no effect. Estimators in italic are those that are already implemented in available packages.
Estimator RCT Obs Ind Cens Dep Cens Outcome model Censoring model Treatment model Robustness
[…]
5.2 Implementation of the estimators
Across different implementations, we use the following shared functions provided in the utilitary.R file.
  • estimate_propensity_score: function to estimate propensity scores $e(X)$ using either parametric (i.e. logistic regression with the argument type_of_model = "glm") or non-parametric methods (i.e. probability forest with the argument type_of_model ="probability forest" based on the probability_forest function from the grf (Tibshirani et al.,, 2017) package). This latter can include cross-fitting (n.folds 1).
  • estimate_survival_function: function to estimate the conditional survival model, which supports either Cox models (argument type_of_model = "cox") or survival forests (argument type_of_model = "survival forest") which uses the function survival_forest from the grf (Tibshirani et al.,, 2017) package. This latter can include cross-fitting (n.folds 1). The estimation can be done with a single learner (argument learner = "S-learner") or two learners (argument learner = "T-learner").
  • estimate_hazard_function: function to estimate the instantaneous hazard function by deriving the cumulative hazard function at each time point. This cumulative hazard function is estimated from the negative logarithm of the survival function.
  • Q_t_hat: function to estimate the remaining survival function at all time points and for all individuals which uses the former estimate_survival_function.
  • Q_Y: function to select the value of the remaining survival function from Q_t_hat at the specific time-to-event.
  • integral_rectangles: function to estimate the integral of a decreasing step function using the rectangle method.
  • expected_survival: function to estimate the integral with $x,y$ coordinate (estimated survival function) using the trapezoidal method.
  • integrate: function to estimate the integral at specific time points Y.grid of a given integrand function which takes initially its values on times using the trapezoidal method.
Unadjusted Kaplan-Meier
Although Kaplan-Meier is implemented in the survival package (Therneau,, 2001), we provide a custom implementation, Kaplan_meier_handmade, for completeness. The difference in Restricted Mean Survival Time, estimated using Kaplan-Meier as in Equation 9 can then be calculated with the RMST_1 function.
As an alternative, one can also use the survfit function in the survival package (Therneau,, 2001) for Kaplan-Meier and specify the rmean argument equal to $\tau$ in the corresponding summary function:
IPCW Kaplan-Meier
We first provide an adjusted.KM function which is then used in the IPCW_Kaplan_meier function to estimate the RMST difference $\hat{\theta}_{\mathrm{IPCW}}$ as in Equation 21. The survival censoring function $G(t|X)$ is computed with the estimate_survival_function utility function from the utilitary.R file.
One could also use the survfit function in the survival package (Therneau,, 2001) by adding IPCW weights for treated and control group and specify the rmean argument equal to $\tau$ in the corresponding summary function:
This alternative approach for IPCW Kaplan-Meier would also be valid for IPTW and IPTW-IPCW Kaplan-Meier.
Buckley-James-based estimator
The function BJ estimates $\theta_{\mathrm{RMST}}$ by implementing the Buckley-James estimator as in Equation 23. It uses two functions available in the utilitary.R file, namely Q_t_hat and Q_Y.
IPTW Kaplan-Meier
The function IPTW_Kaplan_meier implements the IPTW-KM estimator in Equation 29. It uses the estimate_propensity_score function from the utilitary.R.
G-formula
We implement two versions of the G-formula: g_formula_T_learner and g_formula_S_learner. In g_formula_T_learner, separate models estimate survival curves for treated and control groups, whereas g_formula_S_learner uses a single model incorporating both covariates and treatment status to estimate survival time. The latter approach is also available in the RISCA package (Foucher et al.,, 2019) but is limited to Cox models.
IPTW-IPCW Kaplan-Meier
The IPTW_IPCW_Kaplan_meier function implements the IPTW-IPCW Kaplan Meier estimator from Equation 31. It uses the utilitary functions from the utilitary.R file estimate_propensity_score and estimate_survival_function to estimate the nuisance parameters, and the function adjusted.KM which computes an adjusted Kaplan Meier estimator using the appropriate weight.
IPTW-BJ estimator
The IPTW_BJ implements the IPTW-BJ estimator in Equation 32. It uses the utilitary functions, from the utilitary.R file, estimate_propensity_score, Q_t_hat and Q_Y to estimate the nuisance parameters.
AIPTW-AIPCW
The AIPTW_AIPCW function implements the AIPTW_AIPCW estimator in Equation 33 using the utilitary function from the utilitary.R file estimate_propensity_score, Q_t_hat, Q_Y, and estimate_survival_function to estimate the nuisance parameters.
5.3 Available packages
Currently, there are limited number of maintained implementations available for estimating RMST in the presence of right censoring. Among the available tools, noteworthy packages include survRM2 (Hajime et al.,, 2015), grf (Tibshirani et al.,, 2017) and RISCA (Foucher et al.,, 2019). These implementations are used and wrapped in our provided utilitary.R code files.
SurvRM2
The RMST difference with Unadjusted Kaplan-Meier $\hat{\theta}_{KM}$ (Equation 9) can be obtained using the function rmst2 which requires as input the observed time-to-event, the status, the arm which corresponds to the treatment and $\tau$.
RISCA
The RISCA package (Foucher et al.,, 2019) provides several methods for estimating $\theta_{\mathrm{RMST}}$. The RMST difference with Unadjusted Kaplan-Meier $\hat{\theta}_{KM}$ (Equation 9) can be derived using the survfit function from the survival package (Therneau,, 2001) which estimates Kaplan-Meier survival curves for treated and control groups, and then the rmst function calculates the RMST by integrating these curves, applying the rectangle method (type=“s”), which is well-suited for step functions.
The IPTW Kaplan-Meier (Equation 28) can be applied using the ipw.survival and rmst functions. The ipw.survival function requires user-specified weights (i.e. propensity scores). To streamline this process, we define the RISCA_iptw function, which combines these steps and utilizes the estimate_propensity_score from the utilitary.R file.
A single-learner version of the G-formula, as introduced in Section 2.2.1, can be implemented using the gc.survival function. This function requires as input the conditional survival function which should be estimated beforehand with a Cox model via the coxph function from the survival package (Therneau,, 2001). Specifically, the single-learner approach applies a single Cox model incorporating both covariates and treatment, rather than separate models for each treatment arm. We provide a function RISCA_gf that consolidates these steps.
grf
The grf package (Tibshirani et al.,, 2017) enables estimation of the difference between RMST using the Causal Survival Forest approach (Cui et al.,, 2023), which extends the non-parametric causal forest framework to survival data. The RMST can be estimated with the causal_survival_forest function, requiring covariates $X$, observed event times, event status, treatment assignment, and the time horizon $\tau$ as inputs. The average_treatment_effect function then evaluates the treatment effect based on predictions from the fitted forest.
6 Simulations
We compare the behaviors and performances of the estimators through simulations conducted across various experimental contexts. These contexts include scenarios based on RCTs and observational data, with both independent and dependent censoring. We also use semi-parametric and non-parametric models to generate censoring and survival times, as well as logistic or nonlinear models to simulate treatment assignment. For each setting, we generated 100 independent datasets, computed the estimator on each, and summarized the resulting distribution using boxplots.
6.1 RCT
Data Generating Process
We generate RCTs with independent censoring (Scenario 1) and conditionally independent censoring (Scenario 2). We sample $n$ iid datapoints $(X_{i},A_{i},C,T_{i}(0),T_{i}(1))_{i\in[n]}$ where $T_{i}(0),T_{i}(1)$ and $C$ follows Cox models. More specifically, we set
  • $X\sim\mathcal{N}\left(\mu,\Sigma\right)$ where $\mu=(1,1,-1,1)$ and $\Sigma=\mathrm{Id}_{4}$.
  • The hazard function of $T^{(0)}$ is
    $\lambda^{(0)}(t|X)=0.01\exp\left\{\beta_{0}^{\top}X\right\}\quad\text{where}\quad\beta_{0}=(0.5,0.5,-0.5,0.5).$
  • The survival times in the treatment group are given by $T^{(1)}=T^{(0)}+10$.
  • The hazard function of the censoring time $C$ is simply taken as $\lambda_{C}(t|X)=0.03$ in Scenario 1, and in Scenario 2 as
    $\lambda_{C}(t|X)=0.03\cdot\exp\left\{\beta_{C}^{\top}X\right\}\quad\text{where}\quad\beta_{C}=(0.7,0.7,-0.25,-0.1).$
  • The treatment allocation is independent of $X$: $e(X)=0.5$.
  • The threshold time $\tau$ is set to 25.
The descriptive statistics of the two datasets are displayed in Annex (Section 10.1). The graph of the RMST difference as a function of $\tau$ for the two scenarios are displayed below; $\theta_{\mathrm{RMST}}$ is the same in both settings. For RCT scenarios 1 and 2, the ground truth value of the RMST difference at time $\tau=25$ was computed analytically and found to be approximately 7.1.
Estimation of the RMST
For each Scenario, we estimate the RMST difference using the methods summarized in Section 5.1. The methods used to estimate the nuisance components are indicated in brackets: either logistic regression or random forests for propensity scores and either Cox models or survival random forests for survival and censoring models. A naive estimator where censored observations are simply removed and the survival time is averaged for treated and controls is also provided for a naive baseline.
Figure 8 shows the distribution of the RMST difference for 100 simulations in Scenario 1 and different sample sizes: 500, 1000, 2000, 4000. The true value of $\theta_{\mathrm{RMST}}$ is indicated by red dashed line. For clarity, only a subset of estimators with particularly informative or interesting results is displayed; the full set of estimators is provided in Figure 15 (Appendix 11).
Figure 8: Estimation results of the ATE in the simulation of an RCT with independent censoring, using a restricted subset of estimators. The complete results are provided in the appendix (Section 11, Figure 15).
In this setting, and in accordance with the theory, the simplest estimator (unadjusted KM) performs just as well as the others, and presents an extremely small bias (as derived in Section 2.1).
The naive estimator is biased, as expected, and the bias in the manual G-formula S-learner arises because the treatment effect is additive $T^{(1)}=T^{(0)}+10$ and violates the assumption that $T$ would follow a Cox model in the variables $(X,A)$. However, $T|A=a$ is a Cox-model for $a\in\{0,1\}$, which explains the remarkable performance of G-formula (Cox/T-learners) and some of the other models based on a Cox estimation of $S$. Some forest-based versions also yield nearly unbiased results, although their performance can vary: for instance, the T-learner remains well-behaved, while the S-learner appears more sensitive and may require larger sample sizes to achieve convergence. G-formula (Cox/T-learners) achieves the lowest finite-sample variance at 8,000 observations.
For completeness, we have also evaluated the other estimators (e.g. IPTW KM, IPCW KM, BJ, IPTW-IPCW KM, IPTW-BJ, AIPTW-AIPCW, causal survival forest and package versions) in Figure 15 (Appendix 11).
Figure 9: Estimation results of the ATE for the simulation of a RCT with dependent censoring, using a restricted subset of estimators. The complete results are provided in the appendix (Section 11, Figure 16).
Figure 9 shows the results for the RCT simulation with conditionally independent censoring (Scenario 2), focusing on a subset of estimators. The naive estimator is biased, as expected. The IPCW KM (Cox) is slightly biased up to 4,000 observations and quite variable due to extreme values of censoring probabilities. In contrast, the Buckley-James estimator BJ (Cox) is unbiased even with as few as 500 observations. The BJ estimator also demonstrates smaller variance than IPCW methods. G-formula (Cox/T-learners) seem to perform well, even in small samples. As in Scenario 1, G-formula (Cox/ S-learner) remains biased. The forest-based estimators shown in Figure 9 appear biased. Notably, all estimators exhibit higher variability compared to Scenario 1 likely due to the higher censoring rate in this scenario and the estimation on censoring model.
For completeness, we have also evaluated the other estimators (e.g. IPTW KM, IPTW-IPCW KM, IPTW-BJ, AIPTW-AIPCW, causal survival forest and package versions) in Figure 16 (Appendix 11).
6.2 Observational data
Data Generating Process
We carry out two simulations of an observational study with both independent and conditional independent censoring as in scenario 1 and 2. The only difference lies in the simulation of the propensity score, which is no longer constant. For the simulation, an iid $n$-sample $(X_{i},A_{i},C,T_{i}(0),T_{i}(1))_{i\in[n]}$ is generated as in Section 6.1, except for the treatment allocation process that is given by:
$\operatorname{logit}(e(X))=\beta_{A}^{\top}X\quad\text{where}\quad\beta_{A}=(-1,-1,-2.5,-1),$
where we recall that $\operatorname{logit}(p)=\log(p/(1-p))$. The descriptive statistics for the two observational data with independent (Obs1) and conditionally independent censoring (Obs2) are displayed in Appendix (Section 10.2). Note that the survival distribution was not modified, so the target RMST difference remains identical to that of the RCT simulation, namely 7.1 (see Section 6.1).
Estimation of the RMST
Figure 10 below shows the distribution of the estimators of $\theta_{\mathrm{RMST}}$ for the observational study with independent censoring.
Figure 10: Estimation results of the ATE for the simulation of an observational study with independent censoring.
In the simulation of an observational study with independent censoring, confounding bias is introduced, setting it apart from RCT simulations. As expected, the naive estimator is biased while IPTW KM (Log.Reg.) is unbiased. The top-performing estimator in this scenario is the G-formula (Cox/T-learners), which remains unbiased even with as few as 500 observations and achieves the lowest variance. All estimators that use forests to estimate nuisance parameters are biased across sample sizes from 500 to 8000. This setting thus highlights that one should either have an a priori knowledge of the specification of the models or large sample size.
For completeness, we have also evaluated the other estimators (e.g. KM, IPCW KM, BJ, IPTW-IPCW KM, IPTW-BJ, AIPTW-AIPCW, causal survival forest and package versions) in Figure 17 (Appendix 11).
Figure 11 below shows the distribution of the $\theta_{\mathrm{RMST}}$ estimates for the observational study with conditionally independent censoring. The red dashed line represents the true $\theta_{\mathrm{RMST}}$ for $\tau=25$.
Figure 11: Estimation results of the ATE for the simulation of an observational study with dependent censoring.
In the simulation of an observational study with conditionally independent censoring, the top-performing estimators in this scenario are G-formula (Cox/T-learners) and AIPCW-AIPTW (Cox & Cox & Log.Reg.), which are unbiased even with 500 observations. The G-formula (Cox/T-learners) has the lowest variance. Surprisingly, the G-formula (Cox/S-learner) performs quite competitively, showing only a slight bias despite the violation of the proportional hazards assumption. All estimators that use forests to estimate nuisance parameters are biased across sample sizes from 500 to 8000. Although Causal Survival Forest and AIPTW-AIPCW (Forest) are expected to eventually converge, they remain extremely demanding in terms of sample size.
For completeness, we have also evaluated the other estimators (e.g. KM, IPCW KM, BJ, IPTW KM and package versions) in Figure 18 (Appendix 11).
6.3 Misspecification of nuisance components
Data Generating Process
We generate an observational study with covariate interactions and conditionally independent censoring. The objective is to assess the impact of misspecifying nuisance components; specifically, we will use models that omit interactions to estimate these components. This approach enables us to evaluate the robustness properties of various estimators. In addition, in this setting forest-based methods are expected to behave better.
We generate $n$ samples $(X_{i},A_{i},C,T_{i}(0),T_{i}(1))$ as follows:
  • $X\sim\mathcal{N}\left(\mu,\Sigma\right)$ and $\mu=(0.5,0.5,0.7,0.5)$, $\Sigma=\mathrm{Id}_{4}$.
  • The hazard function of $T^{(0)}$ is given by
[…]
    $\text{and}\quad Z$ $=(X_{1}^{2},X_{2}^{2},X_{3}^{2},X_{4}^{2},X_{1}X_{2},X_{1}X_{3},X_{1}X_{4},X_{2}X_{3},X_{2}X_{4},X_{3}X_{4}).$
  • The distribution of $T^{(1)}$ is the one of $T^{(0)}$ but shifted: $T^{(1)}=T^{(0)}+1$.
  • The hazard function of $C$ is given by
[…]
  • The propensity score is
[…]
    When the model is well-specified, the full vector $(X,Z)$ is given as an input of the nuisance parameter models. When it is not, only $X$ and the first half of $Z$ corresponding to $(X_{1}^{2},X_{2}^{2},X_{3}^{2},X_{4}^{2})$ is given as an input.
The true RMST difference at time 0.45 for this scenarios is estimated to be 0.26. The descriptive statistics are given in Appendix (Section 10.3).
Estimation of the RMST
First, we estimate $\theta_{\mathrm{RMST}}$ without any model misspecification to confirm the consistency of the estimators under correctly specified nuisance models. More specifically, it means that for parametric propensity score models, semi-parametric censoring and survival models, we use models including interactions and squared terms, assuming knowledge on the data generating process.
Next, we introduce misspecification individually for the treatment model, censoring model, and outcome model (Figure 13), i.e., we use models without interactions to estimate parametric and semi-parametric nuisance components while the data are generated with interactions.
We further examine combined misspecifications for pairs of models: treatment and censoring, treatment and outcome, and outcome and censoring. Finally, we assess the impact of misspecifying all nuisance models simultaneously (Figure 14).
Figure 12: Estimation results of the ATE for the simulation of an observational study with dependent censoring with interaction.
As before, for completeness, we have also evaluated the other estimators (e.g. KM, IPCW KM, BJ, IPTW KM and package versions) in Figure 19 (Appendix 11).
Figure 13: Estimation results of the ATE for an observational study with dependent censoring in case of a single misspecification.
When there is no misspecification in Figure 12, as expected, IPTW-BJ (Cox & Log.Reg), G-formula (Cox/T-learners) and AIPTW-AIPCW (Cox & Cox & Log.Reg) are unbiased. IPTW-IPCW KM (Cox & Log.Reg) exhibits a bias but seems to converge at larger sample size. Regarding forest-based methods, IPTW-BJ (Forest), AIPTW-AIPCW (Forest) and Causal Survival Forest estimate accurately the RMST difference. Surprisingly, G-formula (Forest/ T-learners), G-formula (Forest/ S-learner) and IPTW-IPCW KM (Forest) exhibit small bias.
Figure 13 shows that AIPTW-AIPCW (Cox & Cox & Log.Reg) remains unbiased when there is one nuisance misspecification. In contrary, the other estimators are biased when one of its nuisance parameter is misspecified.
The results for the other estimators (e.g., KM, IPCW KM, BJ, IPTW KM) are shown in Figure 20 (Appendix 11).
Figure 14: Estimation results of the ATE for an observational study with dependent censoring in case of two or more misspecifications.
Figure 14 shows that, as expected, when all nuisance models are misspecified, all estimators exhibit bias. AIPTW-AIPCW (Cox & Cox & Log.Reg.) seems to converge in case where either the outcome and censoring models, or the treatment and censoring models are misspecificed, which deviates from initial expectations (see Section 3.2.3. It was anticipated that AIPTW-AIPCW would converge solely when both the censoring and treatment models were misspecified.
The results for the other estimators (e.g., KM, IPCW KM, BJ, IPTW KM) are shown in Figure 21 (Appendix 11).
7 Conclusion
Summary of objectives and findings In this work, we provided a comprehensive overview of causal survival estimators targeting the restricted mean survival time (RMST), emphasizing both their theoretical foundations and practical performance. We reviewed a wide range of methods—including regression-based models, weighting approaches, or hybrid estimators-and assessed their behavior in both their parametric and non-parametric versions through extensive simulations. These scenarios covered diverse settings, including randomized trials and observational data, with varying censoring mechanisms and model specifications. Particular attention was given to robustness, sensitivity to misspecification, and estimator variability in finite samples.
Practical recommendations for RCTs In simple contexts such as randomized controlled trials (RCTs), whether subject to independent or dependent censoring, we advise against overcomplicating estimation strategies. Classical estimators like the Kaplan-Meier survival curve estimator (with correction for dependent correction if necessary) remain highly effective in these settings, providing reliable performance without requiring adjustment. Additionally, we recommend the use of the parametric G-formula, even in RCTs, despite its inclusion of covariates which may seem unnecessary when treatment assignment is randomized. Indeed, both theoretical results (Karrison and Kocherginsky,, 2018) and our simulations suggest that adjusting for prognostic covariates—even in randomized settings—can reduce the asymptotic variance of the RMST difference estimator. We also advocate for the use of the Buckley-James (BJ) transformation in this context, which demonstrated excellent performance in our simulations, particularly in reducing bias under independent or dependent censoring without introducing excessive variability.
Practical recommendations for Observational studies Beyond simple settings, non-robust weighted estimators such as IPTW-IPCW and IPTW-BJ were consistently marked by high variability, particularly when applied to observational data with dependent censoring. Parametric G-formula estimators yielded the best performance under correct model specification (as in Denz et al., (2023)), showing both low finite sample bias and variance. However, their sensitivity to model misspecification remains a limitation. In observational settings with dependent censoring, we therefore recommend more robust estimators such as AIPTW-AIPCW, which maintain good performance despite their complexity.
That said, parametric versions may still be preferable in small to moderate sample sizes, as they exhibited lower bias and variance in our simulations—especially for samples under 4,000 observations. In contrast, when sample sizes are sufficiently large and model structure is unknown or complex (e.g., involving interactions or nonlinearities), non-parametric estimators become viable alternatives. In such cases, we specifically recommend the AIPTW-AIPCW estimator, since the non-parametric G-formula lacks formal asymptotic guarantees and is prone to plug-in bias (Chernozhukov et al.,, 2018).
Limitations and Strengths It is important to highlight that our simulation design may disadvantage nonparametric methods, as survival and censoring times were generated from exponential distributions and treatment assignments via logistic regression, naturally favoring parametric models. Finally, to promote further research and ensure reproducibility, we provide full R code implementing all the estimators discussed in this work and reproducing the simulation results.
Perspectives A promising avenue for future research lies in exploring variable selection strategies within causal survival analysis. In standard causal inference, variable selection involves two main objectives. First, ensuring correct identification of the causal effect by appropriately adjusting for confounders—while avoiding the inclusion of instrumental variables, which may increase the variance without reducing bias (Lunceford and Davidian,, 2004). Second, improving the efficiency of the estimators by including precision variables—variables that are predictive of the outcome but not of the treatment assignment—known to reduce variance (Lunceford and Davidian,, 2004; Rotnitzky and Smucler,, 2019). In the context of causal survival analysis, a new variable appears: censoring. It is not clear a priori whether variables that are associated only with the censoring mechanism should be included in the models used for estimation (e.g., in IPCW-based methods). Including such variables might reduce bias in the estimation of the RMST difference but could also increase the variance of the final estimator. Thus, a promising line of research would be to investigate how different types of covariates—confounders, precision variables, instrumental variables, and censoring-related variables—affect both the bias and variance of RMST estimators under various modeling strategies.
Finally, RMST-based estimators provide a valuable alternative to hazard ratios, offering interpretable and robust effect estimates that are particularly well-suited for complex, real-world data. The tools and simulations provided with this work aim to facilitate their broader adoption in applied causal survival analysis.
8 Acknowledgments
This study was funded by Sanofi. Charlotte Voinot and Bernard Sebastien are Sanofi employees and may hold shares and/or stock options in the company. Imke Mayer and Julie Josse have nothing to disclose. The work of CB was supported by the Deutsche Foschungsgemeinschaft (German Research Foundation) on the French-German PRCI ANR ASCAI CA 1488/4-1 “Aktive und Batch-Segmentierung, Clustering und Seriation: Grundlagen der KI”.

### v3

@@
 Abstract
-Causal survival analysis combines survival analysis and causal inference to evaluate the effect of a treatment or intervention on a time-to-event outcome, such as survival time.
-It offers an alternative to relying solely on Cox models for assessing these effects.
-In this paper, we present a comprehensive review of estimators for the average treatment effect measured with the restricted mean survival time, including regression-based methods, weighting approaches, and hybrid techniques.
-We investigate their theoretical properties and compare their performance through extensive numerical experiments.
-Our analysis focuses on the finite-sample behavior of these estimators, the influence of nuisance parameter selection, and their robustness and stability under model misspecification.
-By bridging theoretical insights with practical evaluation, we aim to equip practitioners with both state-of-the-art implementations of these methods and practical guidelines for selecting appropriate estimators for treatment effect estimation.
-Among the approaches considered, G-formula two-learners, AIPCW-AIPTW, Buckley-James estimators, and causal survival forests emerge as particularly promising.
+The restricted mean survival time (RMST) difference offers an interpretable causal contrast to estimate the treatment effect for time-to-event outcomes, yet a wide range of available estimators leaves limited guidance for practice.
+We provide a unified review of RMST estimators for randomized trials and observational studies, establish identification and asymptotic properties, and supply new derivations where needed.
+Our extensive simulation study compares simple nonparametric methods (such as unweighted Kaplan–Meier estimators) alongside parametric and nonparametric implementations of the G-formula, weighting approaches, Buckley–James transformations, and augmented estimators under diverse censoring mechanisms and model specifications.
+Across scenarios, classical Kaplan–Meier estimators (weighted when required by the censoring process) and G-formula methods perform well in randomized settings, while in observational data G-formula estimators remain competitive; however, augmented estimators such as AIPTW–AIPCW generally offer robustness to model misspecification and a favorable bias–variance trade-off.
+Parametric estimators perform best under correct specification, whereas nonparametric methods avoid functional assumptions but require large sample sizes to achieve reliable performance.
+We offer practical recommendations for estimator choice and provide open-source R code to support reproducibility and application.
 1 Introduction
-1.1 Context and motivations
-Causal survival analysis is a growing field that integrates causal inference (Rubin,, 1974; Hernán MA,, 2020) with survival analysis (Kalbfleisch and Prentice,, 2002) to evaluate the impact of treatments on time-to-event outcomes, while accounting for censoring situations where only partial information about an event’s occurrence is available.
-Although it is a relatively new domain, the literature has quickly expanded—often in a fragmented way.
-Researchers face a wide array of estimators and must make several methodological decisions, often without clear guidance.
-Moreover, many of the proposed methods lack accessible implementations.
-This is particularly important given that, in many real-world scenarios where randomized trials are infeasible, well-conducted observational studies—with explicit and plausible causal assumptions—remain the only viable source of evidence about the treatment effect (Makary and Prasad,, 2025).
-This calls for a structured review that consolidates available approaches, clarifies their assumptions, and provides an evaluation of estimator performance to make informed methodological choices — particularly in finite sample settings.
-Such a survey also has the potential to help identify remaining methodological challenges that need to be addressed.
+Causal survival analysis is a rapidly evolving field that integrates causal inference (rubin_1974; Hernan2020) with survival analysis (kalbfleisch2002) to evaluate the impact of treatments on time-to-event outcomes, while accounting for censoring, i.e.
+situations where only partial information about an event’s occurrence is available.
+In this litterature, the Restricted Mean Survival Time (RMST) difference has emerged as a promising causal measure that offers an intuitive interpretation of the average survival time over a specified time horizon.
+Despite its recent emergence, the field has expanded quickly, though often in a fragmented manner.
+As a result, researchers are faced with a broad spectrum of estimators, from weighting and regression-based approaches to more complex hybrid strategies, and must navigate numerous methodological decisions.
+These include, for example, the choice of how to estimate nuisance components, whether through parametric models, or nonparametric techniques, frequently without clear guidance.
+The situation is further complicated by the limited availability of accessible software implementations.
+Addressing these challenges is especially crucial in real-world contexts where randomized trials are not feasible and carefully designed observational studies, grounded in explicit causal assumptions, represent the primary source of evidence on treatment effects (Makary2025).
+These considerations underscore the need for a structured review that consolidates existing approaches, clarifies their underlying assumptions, and evaluates estimator performance to support informed methodological decisions, particularly in finite-sample settings.
+Such a survey could also help to identify key methodological gaps that warrant further research.
 Related efforts have already been made to clarify methodological options in survival analysis.
-For instance, Denz et al., (2023) focus on confounders-adjusted survival curves under proportional hazards and explore model misspecification in that context.
-However, existing reviews often remain within the Cox model framework, and may not fully address the diversity of estimators and causal estimands available—particularly those applicable when the proportional hazards assumption does not hold or when censoring is not independent.
-Despite its widespread use in biomedical research, we decided not to cover the estimation of the Hazard Ratio (HR) which presents several well-documented limitations.
-First, when using Cox model, it relies on the proportional hazards assumption, which is not necessarely satisfied in practice.
-Second, hazard ratios are generally not causally interpretable—even in randomized controlled trials—due to built-in selection bias that can distort the comparison between groups unless the treatment has no effect at all (Martinussen and Vansteelandt,, 2013; Martinussen et al.,, 2020).
-Third, the HR is a non-collapsible measure, meaning that the effect estimated in the entire population (the marginal effect) can differ from the effect estimated within subgroups (the conditional effect), even when there is no confounding (Huitfeldt et al.,, 2019; Greenland et al.,, 1999).
-Finally, the HR could be difficult to interpret clinically, particularly because it reflects an instantaneous risk rather than a cumulative or absolute survival benefit (Averbuch et al.,, 2025).
-In contrast, survival curves are more informative than HRs and can be easily generated (Hernán,, 2010).
-For all these reasons, we focus in this paper on the estimation of the Restricted Mean Survival Time (RMST), a popular causal measure in survival analysis that offers an intuitive interpretation of the average survival time over a specified time horizon.
-Numerous methods have been developed for causal survival analysis, including approaches based on pseudo-observations (Andersen et al.,, 2017; Wang,, 2018; van Hage et al.,, 2024) and targeted maximum likelihood estimation (TMLE) (Stitelman et al.,, 2012; Talbot et al.,, 2025).
-In this work, we focus on estimators based on outcome regression, inverse probability weighting, and hybrid strategies.
-1.2 Definition of the estimand: the RMST
+For instance, denz2023 focus on confounders-adjusted survival curves under proportional hazards and assess the impact of model misspecification.
+Yet, most existing reviews remain confined to the Cox model framework and thus overlook the broader range of estimators and causal estimands now available.
+Despite its widespread use, we chose not to cover the estimation of the hazard ratio (HR), typically derived from a Cox model, because of several well-documented limitations.
+It relies on the proportional hazards assumption, which may not hold in practice.
+Moreover, the HR does not represent a causal contrast (Martinussen2020), even in randomized trials, due to inherent selection bias that can distort group comparisons unless the treatment has no effect (Martinussen2013; Hernan2010_HR).
+The HR is also non-collapsible, meaning marginal effects cannot be inferred from conditional ones (huitfeldt2019collapsibility; greenland1999).
+Finally, its clinical interpretation could be difficult, as it reflects an instantaneous risk rather than a cumulative or absolute survival benefit (averbuch2025).
+In this paper, after introducing the notations and the definition of the restricted mean survival time (Section 1.1), we present the statistical framework required to define and estimate RMST in both randomized controlled trials (Section 2) and observational settings (Section 3).
+We characterize their key statistical properties, such as consistency and asymptotic normality, and provide a unified, self-contained set of proofs.
+Where existing results are incomplete or unavailable, we supply original derivations; where they exist, we rederive them within a common formalism that enhances clarity and facilitates direct comparison between estimators.
+In Section 5, we complement this theoretical analysis with a simulation study assessing estimator performance under various conditions, including independent and conditionally independent censoring and both correctly specified and misspecified models.
+Finally, in Section 6, we offer practical recommendations for estimator selection, grounded in considerations such as finite-sample behavior and computational complexity.
+1.1 Notations and definition of the Restricted Mean Survival Time
 We set the analysis in the potential outcome framework, where a patient, described by a vector of covariates $X\in\mathbb{R}^{p}$, either receives a treatment ($A=1$) or is in the control group ($A=0$).
-The patient will then survive up to a certain time $T^{(0)}\in\mathbb{R}^{+}$ in the control group, or up to a time $T^{(1)}\in\mathbb{R}^{+}$ in the treatment group.
-In practice, we cannot simultaneously have access to $T^{(0)}$ and $T^{(1)}$, as one patient is either treated or control, but only to $T$ defined as follows:
+The patient’s survival time would then be $T^{(0)}\in\mathbb{R}^{+}$ under control and $T^{(1)}\in\mathbb{R}^{+}$ under treatment.
+In practice, both potential outcomes cannot be observed simultaneously for the same individual, since each patient receives only one treatment condition.
+Instead, we observe the realized survival time $T$, defined as follows:
 Assumption 1.1.
-(Stable Unit Treatment Value Assumption — SUTVA)
-$T=AT^{(1)}+(1-A)T^{(0)}$
-Due to potential censoring, the outcome $T$ is not completely observed.
-The most common form of censoring is right-censoring (also known as type II censoring), which occurs when the event of interest has not taken place by the end of the observation period, indicating that it may have occurred later if the observation had continued (Turkson et al.,, 2021).
-We focus in this study on this type of censoring only and we assume that we observe $\tilde{T}=T\wedge C=\min(T,C)$ for some censoring time $C\in\mathbb{R}^{+}$.
-When an observation is censored, the observed time is equal to the censoring time.
-We also assume that we know whether an outcome is censored or not.
-In other words, we observe the event status variable $\Delta=\mathbb{I}\{T\leqslant C\}$, where $\mathbb{I}\{\cdot\}$ is the indicator function.
-$\Delta$ is $1$ if the true outcome is observed, and 0 if it is censored.
-We assume observing a $n$-sample of variables $(X,A,\widetilde{T},\Delta)$ stemming from an $n$-sample of the partially unobservable variables $(X,A,T^{(0)},T^{(1)},C)$.
-A toy example of such data is given in Table 1.
+(Stable Unit Treatment Value Assumption — SUTVA) $T=AT^{(1)}+(1-A)T^{(0)}$.
+Because of potential censoring, the outcome $T$ is not always fully observed.
+The most common form is right-censoring, which occurs when the event has not occurred by the end of follow-up (Turkson2021).
+In what follows, we focus on this case: for each subject, we observe an $n$-sample $(X,A,\tilde{T},\Delta)$, where $\tilde{T}=\min(T,C)$, $C\in\mathbb{R}_{+}$, and $\Delta=\mathbb{I}\{T\leq C\}$ (equal to $1$ if the event time is observed and 0 otherwise).
+These observed variables are derived from the unobservable tuple $(X,A,T^{(0)},T^{(1)},C)$ illustrated in Table 1.
 Table 1: Example of a survival dataset.
-Each observation corresponds to a single patient.
+Each observation corresponds to a patient.
 In practice, only $X$, AA, $\Delta$, and $\widetilde{T}$ are observed.
@@
 9 0.5 3 1 ? ? ? 200 0 200
-Our aim is to estimate the Average Treatment Effect (ATE) defined as the difference between the Restricted Mean Survival Time (Royston and Parmar 2013).
+Our aim is to estimate the Average Treatment Effect (ATE) defined as the difference between the Restricted Mean Survival Time (royston2013).
 Definition 1.1.
-(Causal effect: Difference between Restricted Mean Survival Time)
-$\theta_{\mathrm{RMST}}=\mathbb{E}\left[T^{(1)}\wedge\tau-T^{(0)}\wedge\tau\right],$
-where $a\wedge b:=\min(a,b)$ for $a,b\in\mathbb{R}$, and $\tau>0$ is a pre-specified time horizon over which survival is evaluated.
+(Causal effect: Difference between Restricted Mean Survival Time) We define $\theta_{\mathrm{RMST}}=\mathbb{E}\!\left[T^{(1)}\wedge\tau-T^{(0)}\wedge\tau\right],$ where $a\wedge b:=\min(a,b)$ for $a,b\in\mathbb{R}$, and $\tau>0$ is a pre-specified time horizon over which survival is evaluated.
 We define the survival functions $S^{(a)}(t):=\mathbb{P}(T^{(a)}>t)$ for $a\in\{0,1\}$, i.e., the probability that an individual will survive beyond a given time $t$ if treated or if not treated.
@@
 $\mathbb{E}(T^{(a)}\wedge\tau)=\int_{0}^{\infty}\mathbb{P}(T^{(a)}\wedge\tau>t)dt=\int_{0}^{\tau}S^{(a)}(t)dt.$} (1)
-Consequently, $\theta_{\mathrm{RMST}}$ can be interpreted as the mean difference between the survival function of treated and control until a fixed time horizon $\tau$.
+Consequently, $\theta_{\mathrm{RMST}}$ can be interpreted as the mean difference between the survival function of treated and control until a fixed time horizon $\tau$ (See Figure 1).
 A RMST difference $\theta_{\mathrm{RMST}}=10$ days with $\tau=200$ means that, on average, the treatment increases the survival time by 10 days at 200 days.
-We give a visual interpretation of RMST in Figure 1.
-Figure 1: Plot of the estimated survival curves on synthetic toy-data.
-The $\theta_{\mathrm{RMST}}$ at $\tau=50$ corresponds to the yellow shaded area between the two survival curves.
-The curves have been estimated using the Kaplan-Meier estimator, see Section 2.1.
-Although the present work focuses on the estimation of the difference in RMST, we would like to stress that the causal effect can be assessed through other measures, such as for instance the difference between the survival functions
-$\theta_{\mathrm{SP}}:=S^{(1)}(\tau)-S^{(0)}(\tau)$
-for some time $\tau$, see for instance (Ozenne et al.,, 2020). As mentioned in Section 1.1, another widely used measure (though not necessarily causal) is the hazards ratio, defined as
-$\theta_{\mathrm{HR}}:=\frac{\lambda^{(1)}(\tau)}{\lambda^{(0)}(\tau)},$
-where the hazard function $\lambda^{(a)}$ is defined as
-$\lambda^{(a)}(t):=\lim_{h\to 0^{+}}\frac{\mathbb{P}(T^{(a)}\in[t,t+h]|T^{(a)}\geqslant t)}{h}.$
-in a continuous setting, or as $\lambda^{(a)}(t):=\mathbb{P}(T^{(a)}=t|T^{(a)}\geqslant t)$ when the survival times are discrete.
-The hazard functions and the survival functions are linked through the identities
-$S^{(a)}(t)=\exp\left(-\Lambda^{(a)}(t)\right)\quad\text{where}\quad\Lambda^{(a)}(t):=\int_{0}^{t}\lambda^{(a)}(s)\mathop{}\!\mathrm{d}s,$} (2)
-in the continuous case. The functions $\Lambda^{(a)}$ are called the cumulative hazard functions. In the discrete case, we have in turn
-$S^{(a)}(t)=\prod_{t_{k}\leqslant t}\left(1-\lambda^{(a)}(t_{k})\right),$} (3)
-where $\{t_{1},\dots,t_{K}\}$ are the discrete event times of $T^{(a)}$. These hazard functions are classically used to model the survival times and the censoring times, see Section 2.2.1.
-1.3 Organization of the paper
-In this paper, we detail the minimal theoretical framework that allows the analysis of established RMST estimators in the context of both Randomized Controlled Trials (Section 2) and observational data (Section 3).
-We give their statistical properties (consistency, asymptotic normality) along with proofs when possible.
-We then conduct in Section 6 a numerical study of these estimators through simulations under various experimental conditions, including independent and conditionally independent censoring and correct and incorrect model specifications.
-We conclude in Section 7 with practical recommendations on estimator selection, based on criteria such as asymptotic behavior, computational complexity, and efficiency.
-1.4 Notations
-We provide in Table LABEL:tbl-reminder_notations a summary of the notation used throughout the paper.
+Although the present work focuses on the estimation of the difference of RMST, we would like to stress that the causal effect can be assessed through other measures, such as for instance the difference between the survival functions $\theta_{\mathrm{SP}}:=S^{(1)}(\tau)-S^{(0)}(\tau)$ for some time $\tau$ (Ozenne20_AIPTW_AIPCW).
+Figure 1: Estimated survival curves on synthetic data.
+The $\theta_{\mathrm{RMST}}$ at $\tau=50$ corresponds to the yellow area between them.
+Curves are estimated with the Kaplan-Meier estimator, see Section 2.1.
+We provide in Table LABEL:tbl-reminder_notations a summary of the notations used throughout the paper.
 Table 2: Summary of the notations.
@@
 $\lambda^{(a)},a\in\{0,1\}$ Hazard function $\lambda^{(a)}(t)=\lim_{h\to 0^{+}}\mathbb{P}(T^{(a)}\in[t,t+h)|T^{(a)}\geqslant t)/h$ of the potential survival times
-$\Lambda^{(a)},a\in\{0,1\}$ Cumulative hazard function of the potential survival times
+$\Lambda^{(a)},a\in\{0,1\}$ Cumulative hazard function $\Lambda^{(a)}(t):=\int_{0}^{t}\lambda^{(a)}(s)\mathop{}\!\mathrm{d}s$ of the potential survival times
 $C$ Censoring time
@@
 $G$ Left-limit of the survival function $G(t)=\mathbb{P}(C\geqslant t)$ of the censoring time
-$\widetilde{T}$ Observed time ($T\wedge C$)
+$\widetilde{T}$ Observed time ($T\wedge C=\min(T,C)$)
 $\Delta$ Event status $\mathbb{I}\{T\leqslant C\}$
@@
 Randomized Controlled Trials (RCTs) are the gold standard for establishing the effect of a treatment on an outcome, because treatment allocation is controlled through randomization, which ensures (asymptotically) the balance of covariates between treated and controls, and thus avoids problems of confounding between treatment groups.
-The core assumption in a classical RCT is the random assignment of the treatment (D.
-B.
-Rubin 1974).
+In a classical RCT, treatment is randomly assigned (rubin_1974), and everyone could, in principle, be assigned to either arm.
 Assumption 2.1.
-(Random Treatment Assignment)
-$A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X$
-We also assume that there is a positive probability of receiving the treatment, which we rephrase under the following assumption.
+(Random Treatment Assignment) $A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X$.
 Assumption 2.2.
-(Trial Positivity)
-$0<\mathbb{P}(A=1)<1$
-Under Assumptions 1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) and 2.2 (.Causal survival analysis in Randomized Controlled Trials), classical causal identifiability equations can be written to express $\theta_{\mathrm{RMST}}$ without potential outcomes.
+(Trial Positivity) $0<\mathbb{P}(A=1)<1$.
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials) and 2.2 (.Causal survival analysis in Randomized Controlled Trials), classical causal identifiability equations can be written to express $\theta_{\mathrm{RMST}}$ without potential outcomes.
 $\theta_{\mathrm{RMST}}$ $=\mathbb{E}[T^{(1)}\wedge\tau-T^{(0)}\wedge\tau]$ (2)
@@
 $=\mathbb{E}[T\wedge\tau|A=1]-\mathbb{E}[T\wedge\tau|A=0].$ (SUTVA)
-However, Equation 4 still depends on $T$, which remains only partially observed due to censoring.
-To ensure that censoring does not compromise the identifiability of treatment effects, we must impose certain assumptions on the censoring process, standards in survival analysis, namely, independent censoring and conditionally independent censoring.
-These assumptions lead to different estimation approaches.
-We focus on two strategies: those that aim to estimate $\mathbb{E}[T\wedge\tau|A=a]$ directly (e.g., through censoring-unbiased transformations, see Section 2.2), and those that first estimate the survival curves to derive RMST via Equation 1 (such as the Kaplan-Meier estimator and all its variants, see the next Section).
+However, Equation 2 still depends on $T$, which is only partially observed due to censoring.
+To ensure identifiability of treatment effects, standard assumptions on the censoring mechanism are required, namely independent and conditionally independent censoring.
+When censoring is independent, standard survival estimators, such as the Kaplan–Meier estimator, directly accommodate censoring, allowing RMST to be obtained from the estimated survival curve via Equation 1 (Section 2.1).
+Under the more general condition of conditionally independent censoring, censoring-unbiased transformations of the event time can be used to recover $\mathbb{E}[T\wedge\tau\mid A=a]$ (Section 2.2).
+These transformations enable two estimation strategies: (i) direct estimation, by computing the transformed outcome and averaging it and (ii) adjusted survival-curve estimation, in which the transformation is incorporated within a survival estimator, and RMST is then derived using Equation 1.
+These two families of methods form the basis of the estimators reviewed in the next sections.
 2.1 Independent censoring: the Kaplan-Meier estimator
-In a first approach, one might assume that the censoring times are independent on the rest of the variables.
 Assumption 2.3.
-(Independent Censoring)
-$C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X,A.$
+(Independent Censoring) $C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)},X,A.$
 Under Assumption 2.3, subjects censored at time $t$ are representative of all subjects who remain at risk at time $t$.
-Figure 2 represents the causal graph when the study is randomized and outcomes are observed under independent censoring.
-The treatment assignment $A$ is randomized and hence independent of baseline covariates $X$, which may influence the potential survival outcomes $T^{(a)}$.
-The observed survival time $T$ depends on both $A$ and $X$, while the censoring time $C$ is assumed to be conditionally independent of the potential outcomes given $X$, reflecting an independent censoring mechanism.
-This structure encodes key assumptions used throughout this section, including consistency, random treatment assignment, and independent censoring.
+Figure 2 shows the causal graph where the treatment assignment is random, the time $T$ depends on $(A,X)$, and the censoring time $C$ is independent of all variables.
 Figure 2: Causal graph in RCT survival data with independent censoring.
-We also assume that there is no almost-sure upper bound on the censoring time before $\tau$, which we rephrase under the following assumption.
+We assume that there is no almost-sure upper bound on the censoring time before $\tau$.
 Assumption 2.4.
-(Positivity of the Censoring Process)
-There exists $\varepsilon>0$ such that
-$G(t)\geqslant\varepsilon\quad\text{for all}\quad t\in[0,\tau).$
+(Positivity of the Censoring Process) There exists $\varepsilon>0$ such that $G(\tau)\geqslant\varepsilon$.
 If indeed it was the case that $\mathbb{P}(C<t)=1$ for some $t<\tau$, then we would not be able to infer anything on the survival function on the interval $t,\tau]$ as all observation times $\widetilde{T}_{i}$ would be in $[0,t]$ almost surely.
-In practice, adjusting the threshold time $\tau$ can help satisfy the positivity assumption.
-For instance, in a clinical study, if a subgroup of patients has zero probability of remaining uncensored at a given time, $\tau$ can be modified to ensure that participants have a feasible chance of remaining uncensored up to the revised threshold.
-The two Assumptions 2.3 (.Independent censoring: the Kaplan-Meier estimator) and 2.4 (.Independent censoring: the Kaplan-Meier estimator) together allow the distributions of $T^{(a)}$ to be identifiable, in the sense that there exists an identity that expresses $S^{(a)}$ as a function of the joint distribution of $(\widetilde{T},\Delta,A=a)$, see for instance (Ebrahimi et al.,, 2003) for such a formula in a non-causal framework.
-This enables several estimation strategies, the most well-known being the Kaplan-Meier product-limit estimator.
-To motivate the definition of the latter and make explicit the identifiability identity, we set the analysis in the discrete case.
-We let $\{t_{k}\}_{k\geqslant 1}$ be a set of positive and increasing times and assume that $T\in\{t_{k}\}_{k\geqslant 1}$ almost surely.
-Then for any $t\in[0,\tau]$, it holds, letting $t_{0}=0$ by convention, thanks to Equation 3,
-$S(t|A=a)$ $=\mathbb{P}(T>t|A=a)=\prod_{t_{k}\leqslant t}\left(1-\mathbb{P}(T=t_{k}|T>t_{k-1},A=a)\right)$
-$=\prod_{t_{k}\leqslant t}\left(1-\frac{\mathbb{P}(T=t_{k},A=a)}{\mathbb{P}(T\geqslant t_{k},A=a)}\right).$
-Using Assumptions 2.3 (.Independent censoring: the Kaplan-Meier estimator) and 2.4 (.Independent censoring: the Kaplan-Meier estimator), we find that
-$\frac{\mathbb{P}(T=t_{k},A=a)}{\mathbb{P}(T\geqslant t_{k},A=a)}=\frac{\mathbb{P}(T=t_{k},C\geqslant t_{k},A=a)}{\mathbb{P}(T\geqslant t_{k},C\geqslant t_{k},A=a)}=\frac{\mathbb{P}(\widetilde{T}=t_{k},\Delta=1,A=a)}{\mathbb{P}(\widetilde{T}\geqslant t_{k},A=a)},$} (5)
-yielding the final identity
-$S(t|A=a)=\prod_{t_{k}\leqslant t}\left(1-\frac{\mathbb{P}(\widetilde{T}=t_{k},\Delta=1,A=a)}{\mathbb{P}(\widetilde{T}\geqslant t_{k},A=a)}\right).$} (6)
-Notice that the right hand side only depends on the distribution of the observed tuple $(A,\widetilde{T},\Delta)$. This last equation suggests in turn to introduce the quantities
-$D_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A=a)\quad\text{and}\quad N_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A=a),$} (7)
-which correspond respectively to the number of deaths $D_{k}(a)$ and of individuals at risk $N_{k}(a)$ at time $t_{k}$ in the treated group (a=1) or in the control group (a=0).
+In practice, adjusting the threshold time $\tau$ can help satisfy this positivity assumption.
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials), 2.2 (.Causal survival analysis in Randomized Controlled Trials), 2.3 (.Independent censoring: the Kaplan-Meier estimator), and 2.4 (.Independent censoring: the Kaplan-Meier estimator), the potential-outcome survival $S^{(a)}(t)$ is identified as a functional of the joint law of ($\tilde{T},\Delta,A$).
+This enables several estimation strategies, the most well-known being the Kaplan-Meier, the Maximum Likelihood Estimator (MLE) of the survival functions, see for instance Kaplan1958_1.
 Definition 2.1.
-(Kaplan-Meier estimator, (Kaplan and Meier,, 1958)) With $D_{k}(a)$ and $N_{k}(a)$ defined in Equation 7, we let
-$\widehat{S}_{\mathrm{KM}}(t|A=a):=\prod_{t_{k}\leqslant t}\left(1-\frac{D_{k}(a)}{N_{k}(a)}\right).$} (8)
-The associated RMST estimator is then simply defined as
-$\widehat{\theta}_{\mathrm{KM}}=\int_{0}^{\tau}\widehat{S}_{KM}(t|A=1)-\widehat{S}_{KM}(t|A=0)dt.$} (9)
-The Kaplan-Meier estimator is the Maximum Likelihood Estimator (MLE) of the survival functions, see for instance Kaplan and Meier, (1958).
-Furthermore, because $D_{k}(a)$ and $N_{k}(a)$ are sums of i.i.d.
-random variables, the Kaplan-Meier estimator inherits some convenient statistical properties.
-In Annex 9 (Section 9.1) we show that it exhibits exponentially fast bias decay (Proposition 9.1), asymptotic normality with explicitly characterized variances (Proposition 9.2).
-We finally mention that the KM estimator as defined in Definition 2.1 still makes sense in a non-discrete setting, and one only needs to replace the fixed grid $\{t_{k}\}$ by the values at which we observed an event $(\widetilde{T}_{i}=t_{k},\Delta_{i}=1)$.
-We refer to Breslow, (1974) for a study of this estimator in the continuous case and to Aalen et al., (2008), Sec 3.2 for a general study of the KM estimator through the prism of point processes.
+(Kaplan-Meier estimator, (Kaplan1958_1))
+$\widehat{S}_{\mathrm{KM}}(t|A=a):=\prod_{t_{k}\leqslant t}\left(1-\frac{D_{k}(a)}{N_{k}(a)}\right).$
+with $D_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A=a)\quad\text{and}\quad N_{k}(a):=\sum_{i=1}^{n}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A=a),$ which correspond respectively to the number of deaths $D_{k}(a)$ and of individuals at risk $N_{k}(a)$ at time $t_{k}$ in the treated group ($a=1$) or in the control group ($a=0$).
+Furthermore, since $D_{k}(a)$ and $N_{k}(a)$ are sums of i.i.d.
+random variables, the Kaplan–Meier estimator inherits convenient statistical properties.
+In Appendix A.1.1.
+Proofs of Section 2.1, we show that it exhibits exponentially fast bias decay (Proposition 6.1) and asymptotic normality with explicitly characterized variances (Proposition 6.2).
+Finally, the estimator in Definition 2.1 extends naturally to the continuous setting by replacing the fixed grid $\{t_{k}\}$ with observed event times $(\widetilde{T}_{i}=t_{k},\Delta_{i}=1)$.
+We refer to Breslow1974 for a study of this version and to (Aalen2008, Sec.
+3.2) for a broader point-process perspective.
+The associated RMST estimator is then defined as
+$\widehat{\theta}_{\mathrm{KM}}=\int_{0}^{\tau}\widehat{S}_{\mathrm{KM}}(t|A=1)-\widehat{S}_{\mathrm{KM}}(t|A=0)dt.$} (3)
+This estimator inherits KM properties.
+It is consistent for the target RMST and its bias decays exponentially fast.
+In RCTs, the independence of arms delivers asymptotic normality with variance equal to the sum of the arm-specific RMST variances.
 2.2 Conditionally independent censoring
-An alternative hypothesis in survival analysis that relaxes the assumption of independent censoring is conditionally independent censoring, also referred sometimes as informative censoring.
-It allows to model more realistic censoring processes, in particular in situations where there are reasons to believe that $C$ may be dependent on $A$ and $X$ (for instance, if a patient is more likely to leave the study when treated because of side-effects of the treatment).
+Conditionally independent censoring relaxes the assumption of independent censoring by allowing censoring to depend on treatment or covariates — for instance, when treated patients withdraw from a study due to side effects.
 Assumption 2.5.
-(Conditionally Independent Censoring)
-$C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X,A$
+(Conditionally Independent Censoring) $C\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X,A$
 Under Assumption 2.5, subjects within a same stratum defined by $X=x$ and $A=a$ have equal probability of censoring at time $t$, for all $t$.
@@
 Assumption 2.6.
-(Positivity / Overlap for Censoring)
-There exists $\varepsilon>0$ such that for all $t\in[0,\tau]$, it almost surely holds
-$G(t|A,X)\geqslant\varepsilon.$
-Figure 3 represents the causal graph when the study is randomized with conditionally independent censoring.
+(Conditional Positivity / Overlap for Censoring) There exists $\varepsilon>0$ such that for all $t\in[0,\tau]$, it almost surely holds $G(t|A,X)\geqslant\varepsilon.$
+Figure 3 represents the causal graph when treatment assignment is randomized and when there is a conditionally independent censoring mechanism.
 Figure 3: Causal graph in RCT survival data with dependent censoring.
-Under dependent censoring, the Kaplan-Meier estimator as defined in Definition 2.1 can fail to estimate survival, in particular because Equation 5 no longer holds.
+Under dependent censoring, the Kaplan-Meier estimator as defined in Definition 2.1 is not suited to estimate survival.
 Alternatives include plug-in G-formula estimators (Section 2.2.1) and unbiased transformations (Section 2.2.2).
-2.2.1 The G-formula and the Cox Model
-Because the censoring is now independent on the potential outcome conditionally to the covariates, it would seem natural to model the response of the survival time conditionally to these covariates too:
-$\mu(x,a):=\mathbb{E}[T\wedge\tau|X=x,A=a].$
-Building on Equation 4, one can express the RMST as a function of $\mu$:
+2.2.1 The G-formula
+Building on Equation 2, one can express the RMST conditioning on the covariates:
 $\theta_{\mathrm{RMST}}=\mathbb{E}\left[\mathbb{E}[T\wedge\tau|X,A=1]\right]-\mathbb{E}\left[T\wedge\tau|X,A=0]\right]=\mathbb{E}[\mu(X,1)-\mu(X,0)].$
-An estimator $\widehat{\mu}$ of $\mu$ would then straightforwardly yields an estimator of the RMST difference through the so-called G-formula plug-in estimator:
+with $\mu(x,a):=\mathbb{E}[T\wedge\tau|X=x,A=a].$
+An estimator of the RMST difference via the G-formula arises by plugging in any estimator $\widehat{\mu}$ of $\mu$:
 $\widehat{\theta}_{\mathrm{G-formula}}=\frac{1}{n}\sum_{i=1}^{n}\widehat{\mu}\left(X_{i},1\right)-\widehat{\mu}\left(X_{i},0\right).$} (4)
-We would like to stress that a G-formula approach works also in a observational context as the one introduced in Section 3.2.
-However, because the estimation strategies presented in the next sections rely on estimating nuisance parameters, and that this latter task is often done in the same way as we estimate the conditional response $\mu$, we decided to not delay the introduction of the G-formula any further, and we present below a few estimation methods for $\mu$.
-These methods are subdivided in two categories: $T$-learners (i.e.
-two learners), where $\mu(\cdot,1)$ is estimated separately from $\mu(\cdot,0)$, and $S$-learner (i.e.
-single learner), where $\widehat{\mu}$ is obtained by fitting a single model based on covariates $(X,A)$.
-Cox Model.
-There are many ways to model $\mu$ in a survival context, the most notorious of which being the Cox proportional hazards model (Cox,, 1972).
-It relies on a semi-parametric model of the conditional hazard functions and specifies the function under treatment $a$, $\lambda^{(a)}(t|X)$, as
+The G-formula accommodates right censoring under conditionally independent censoring, without specifying a separate censoring model: the target (e.g., $S^{(a)}(t)$ or $\mathrm{RMST}^{(a)}$) is expressed as a functional of the conditional outcome regression $\mu(\cdot)$.
+The same G-formula can be applied to observational settings (in Section 3.2) to address non-random treatment assignment.
+Different strategies can be used to estimate $\mu$, ranging from semi-parametric models such as the Cox model to fully non-parametric approaches like survival forests.
+For any chosen model, implementation follows one of two paradigms: T-learner approaches (i.e., two separate learners), where $\mu(\cdot,1)$ and $\mu(\cdot,0)$ are estimated separately; and S-learner approaches (i.e., a single learner), where $\widehat{\mu}$ is obtained by fitting one model using covariates $(X,A)$.
+Cox Model. The Cox proportional hazards model (Cox1972) relies on a semi-parametric model of the conditional hazard functions and specifies the function under treatment $a$, $\lambda^{(a)}(t|X)$, as
 $\lambda^{(a)}(t|X)=\lambda_{0}^{(a)}(t)\exp(X^{\top}\beta^{(a)}),$
-where $\lambda^{(a)}_{0}$ is a baseline hazard function under treatment $a$ and $\beta^{(a)}$ has the same dimension as the vector of covariate $X$ (We use the superscript $(a)$ here to emphasize that a separate Cox model is fit for each treatment level, since the treatment assignment $A$ is not included as a covariate in $X$).
-The conditional survival function then takes the simple form (in the continuous case)
-$S^{(a)}(t|X)=S^{(a)}_{0}(t)^{\exp(X^{\top}\beta^{(a)})},$
-where $S^{(a)}_{0}(t)$ is the survival function associated with $\lambda_{0}^{(a)}$.
-The vector $\beta^{(a)}$ is classically estimated by maximizing the so-called partial likelihood function as introduced in the original paper of Cox, (1972).
-The cumulative baseline hazard function can be estimated through the Breslow’s estimator (Breslow,, 1974):
-$\widehat{\Lambda}_{0}^{(a)}(t)=\sum_{\Delta_{i}=1,\widetilde{T}_{i}\leqslant t}\frac{1}{\displaystyle\sum_{\widetilde{T}_{j}\geqslant\widetilde{T}_{i}}\exp(X_{j}^{\top}\widehat{\beta}^{(a)})}$
-where $\widehat{\beta}^{(a)}$ is a partial likelihood maximizer.
-One can show that $(\widehat{\beta}^{(a)},\widehat{\Lambda}_{0}^{(a)})$ is the MLE of the true likelihood, when $\widehat{\Lambda}_{0}^{(a)}$ is optimized over all step functions of the form
-$\Lambda_{0}(t):=\sum_{\Delta_{i}=1}h_{i},\quad h_{i}\in\mathbb{R}^{+}.$
-This fact was already hinted in the original paper by Cox and made rigorous in many subsequent papers, see for instance Fan et al., (2010).
-Furthermore, if the true distribution follows a Cox model, then both $\widehat{\beta}^{(a)}$ and $\widehat{\Lambda}_{0}^{(a)}$ are strongly consistent and asymptotically normal estimator of the true parameters $\beta^{(a)}$ and $\Lambda^{(a)}$, see Kalbfleisch and Prentice, (2002), Sec 5.7.
-When using a $T$-learner approach, one simply finds $(\widehat{\beta}^{(a)},\widehat{\Lambda}_{0}^{(a)})$ for $a\in\{0,1\}$ by considering the control group and the treated group separately.
-When using a $S$-learner approach, the treatment status $A$ becomes a covariate a the model becomes
-$\lambda(t|X,A)=\lambda_{0}(t)\exp(X^{\top}\beta+\alpha A).$} (11)
-for some $\alpha\in\mathbb{R}$.
-One main advantage of Cox model is that it makes it very easy to interpret the effect of a covariate on the survival time.
-If indeed $\alpha>0$, then the treatment has a negative effect on the survival times.
-Likewise, if $\beta_{i}>0$, then the $i$-th coordinate of $X$ has a negative effect as well.
-We finally mention that the hazard ratio takes a particularly simple form under the later model since
-$\theta_{\mathrm{HR}}=e^{\alpha}.$
-In particular, it does not depends on the time horizon $\tau$, and is thus sometimes referred to as proportional hazard.
-Figure 4 illustrates the estimation of the difference in Restricted Mean Survival Time using G-formula with Cox models.
-Finally, while we have focused here on Cox-based specifications due to its widespread use in biomedical research.
-However, alternative approaches exist.
-For example, Accelerated Failure Time (AFT) models (Saikia and Barman,, 2017) provide a semi-parametric alternative that directly models the survival time and does not rely on the proportional hazards (PH) assumption.
-Figure 4: Illustration of the G-formula for estimating $\theta_{\mathrm{RMST}}$ in an RCT when only one covariate $X_{1}$ influences the outcome.
-Weibull Model. A popular parametric model for survival is the Weibull Model, which amounts to assume that
-$\lambda^{(a)}(t|X)=\lambda_{0}^{(a)}(t)\exp(X^{\top}\beta)$
-where $\lambda_{0}^{(a)}(t)$ is the instantaneous hazard function of a Weibull distribution, that is to say a function proportional to $t^{\gamma}$ for some shape parameter $\gamma>0$.
-We refer to Zhang, (2016) for a study of this model.
-Survival Forests. On the non-parametric front, we mention the existence of survival forests (Ishwaran et al.,, 2008).
-2.2.2 Censoring-unbiased transformations
+where $\lambda^{(a)}_{0}$ is a baseline hazard function under treatment $a$ and $\beta^{(a)}$ has the same dimension as the vector of covariate $X$.
+The model is semi-parametric in the sense that the baseline hazard is left unspecified, while the effect of covariates enters through a multiplicative proportional hazards structure.
+Alternative semi-parametric models include Accelerated Failure Time (AFT) (saikia2017) which directly models the survival time and does not rely on the proportional hazards (PH) assumption.
+Weibull Model.
+A popular parametric model for survival is the Weibull Model, which amounts to assume that $\lambda^{(a)}(t|X)=t^{\gamma(a)}\exp(X^{\top}\beta)$ where $\gamma(a)>0$.
+We refer to zhang2016 for a study of this model.
+Survival Forests.
+On the nonparametric side, survival forests (ishwaran2008) aggregate survival trees that use log-rank splitting and terminal-node Nelson–Aalen estimates to learn $S(t\mid X)$ under right censoring, capturing nonlinearities and high-order interactions without parametric form.
+2.2.2 Censoring-unbiased transformations (CUT)
 Censoring-unbiased transformations involve applying a transformation to $T$. Specifically, we compute a new time $T^{*}$ of the form
 $T^{*}:=T^{*}(\widetilde{T},X,A,\Delta)=\begin{cases}\phi_{0}(\widetilde{T}\wedge\tau,X,A)\quad&\text{if}\quad\Delta^{\tau}=0,\\ \phi_{1}(\widetilde{T}\wedge\tau,X,A)\quad&\text{if}\quad\Delta^{\tau}=1.\end{cases}$} (5)
-for two wisely chosen transformations $\phi_{0}$ and $\phi_{1}$, and where
-$\Delta^{\tau}:=\mathbb{I}\{T\wedge\tau\leqslant C\}=\Delta+(1-\Delta)\mathbb{I}(\widetilde{T}\geqslant\tau)$} (13)
-is the indicator of the event where the individual is either uncensored or censored after time $\tau$.
-The idea behind the introduction of $\Delta^{\tau}$ is that because we are only interested in computing the expectation of the survival time thresholded by $\tau$, any censored observation coming after time $\tau$ can in fact be considered as uncensored ($\Delta^{\tau}=1$).
-A censoring-unbiased transformation $T^{*}$ shall satisfy: for $a\in\{0,1\}$, it holds
+for two wisely chosen transformations $\phi_{0}$ and $\phi_{1}$, and where $\Delta^{\tau}:=\mathbb{I}\{T\wedge\tau\leqslant C\}=\Delta+(1-\Delta)\mathbb{I}(\widetilde{T}\geqslant\tau)$ indicates whether an individual is either uncensored or censored after time $\tau$.
+Since we only need the expectation of survival time thresholded at $\tau$, any censoring beyond $\tau$ can be treated as uncensored ($\Delta^{\tau}=1$).
+A censoring-unbiased transformation $T^{*}$ is defined such that, for any $a\in\{0,1\}$,
 $\mathbb{E}[T^{*}|A=a,X]=\mathbb{E}[T^{(a)}\wedge\tau|X]\quad\text{almost surely.}$} (6)
-A notable advantage of this approach is that it enables the use of the full transformed dataset $(X_{i},A_{i},T^{*}_{i})$ as if no censoring occurred. Because it holds
+A notable advantage of this approach is that it enables the use of the full transformed dataset $(X_{i},A_{i},T^{*}_{i})$ as if no censoring occurred. Moreover, by the reweighting identity
 $\mathbb{E}[\mathbb{E}[T^{*}|A=a,X]]=\mathbb{E}\left[\frac{\mathbb{I}\{A=a\}}{\mathbb{P}(A=a)}T^{*}\right],$} (7)
-One can derive an estimator of the RMST difference from any censoring-unbiased transformation $T^{*}$ as:
-$\widehat{\theta}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{i}$} (16)
+this equality links the arm-specific mean of $T^{*}$ to its full-sample inverse-probability–weighted representation. The RMST difference can be estimated by
+$\widehat{\theta}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{i},$} (8)
 where $\pi=\mathbb{P}(A=1)\in(0,1)$ by Assumption 2.2 and where $T^{*}_{i}=T^{*}(\widetilde{T}_{i},X_{i},A_{i},\Delta_{i})$.
-The following proposition establishes the theoretical guarantees for the class of estimators constructed from bounded censoring-unbiased transformations:
-Proposition 2.1.
+The following proposition establishes the theoretical guarantees for the class of estimators constructed from bounded censoring-unbiased transformations.
+Proposition 2.1 (Consistency of CUT as RMST estimator with known $\pi$).
 Under Assumptions 2.1 (.Causal survival analysis in Randomized Controlled Trials) and 2.2 (.Causal survival analysis in Randomized Controlled Trials), the estimator $\widehat{\theta}$ derived as in Equation 8 from square integrable censoring-unbiased transformations satisfying Equation 6 is an unbiased, strongly consistent, and asymptotically normal estimator of the RMST difference.
@@
 It is natural in a RCT setting to assume that probability of being treated $\pi$ is known.
-If not, it is usual to replace $\pi$ by its empirical counterpart $\widehat{\pi}=n_{1}/n$ where $n_{a}=\sum_{i}\mathds{1}\{A=a\}$.
+If not, it is usual to replace $\pi$ by its empirical counterpart $\widehat{\pi}=n_{1}/n$ where $n_{a}=\sum_{i}\mathbbm{1}\{A=a\}$.
 The resulting estimator takes the form
 $\widehat{\theta}=\frac{1}{n_{1}}\sum_{A_{i}=1}T^{*}_{i}-\frac{1}{n_{0}}\sum_{A_{i}=0}T^{*}_{i}.$} (9)
-While this estimator introduces a slight bias due to the estimation of $\pi$ (see for instance Colnet et al., (2022), Lemma 2), it remains strongly consistent and asymptotically normal under the same assumptions:
-Proposition 2.2.
+While this estimator introduces a slight bias due to the estimation of $\pi$ (see for instance colnet2022, Lemma 2), it remains strongly consistent and asymptotically normal under the same assumptions:
+Proposition 2.2 (Consistency of CUT as RMST estimator with estimated $\hat{\pi}$).
 Under Assumptions 2.1 (.Causal survival analysis in Randomized Controlled Trials) and 2.2 (.Causal survival analysis in Randomized Controlled Trials), the estimator $\widehat{\theta}$ derived as in Equation 9 from a square integrable censoring-unbiased transformations satisfying Equation 6 is a strongly consistent, and asymptotically normal estimator of the RMST difference.
-The two most popular transformations are Inverse-Probability-of-Censoring Weighting (Koul et al.,, 1981) and Buckley-James (Buckley and James,, 1979), both illustrated in Figure 5 and detailed below.
+For the later, we consider only estimator with estimated probability of being treated in Equation 9.
+The two most popular transformations are Inverse-Probability-of-Censoring Weighting (koul1981) and Buckley-James (buckley1979), both illustrated in Figure 4 and detailed below.
+Figure 4: Illustration of Inverse-Probability-of-Censoring and Buckley-James transformations.
 In the former, only non-censored observations are considered and they are weighted while in the latter, censored observations are imputed with an estimated survival time.
-Figure 5: Illustration of Inverse-Probability-of-Censoring and Buckley-James transformations.
 The Inverse-Probability-of-Censoring Weighted transformation
-The Inverse-Probability-of-Censoring Weighted (IPCW) transformation, introduced by Koul et al., (1981) in the context of censored linear regression, involves discarding censored observations and applying weights to uncensored data.
+The Inverse-Probability-of-Censoring Weighted (IPCW) transformation, introduced by koul1981 in the context of censored linear regression, involves discarding censored observations and applying weights to uncensored data (see Figure 4).
 More precisely, we let
@@
 where we recall that $G(t|X,A):=\mathbb{P}(C\geqslant t|X,A)$ is the left limit of the conditional survival function of the censoring.
-This estimator assigns higher weights to uncensored subjects within a covariate group that is highly prone to censoring, thereby correcting for conditionally independent censoring and reducing selection bias (Howe et al.,, 2016).
-We establish in Appendix 9, Section 9.2, (Proposition 9.3) that the IPCW transformation is indeed a censoring-unbiased transformation in the sense of Equation 14.
+This estimator assigns higher weights to uncensored subjects within a covariate group that is highly prone to censoring, thereby correcting for conditionally independent censoring and removing selection bias (howe2016).
+Using proposition 2.1 and 2.2, we show in Appendix that the IPCW transformation is a censoring-unbiased transformation in the sense of Equation 6 (Proposition 6.3).
 The IPCW depends on the unknown conditional survival function of the censoring $G(\cdot|X,A)$, which thus needs to be estimated.
-Estimating conditional censoring or the conditional survival function can be approached similarly, as both involve estimating a time—whether for survival or censoring.
-Consequently, we can use semi-parametric methods, such as the Cox model, or non-parametric approaches like survival forests, and we refer to Section 2.2.1 for a development on these approaches.
-Once an estimator $\widehat{G}(\cdot|A,X)$ of the later is provided, one can construct an estimator of the difference in RMST based on Equation 16 or Equation 17
-$\widehat{\theta}_{\mathrm{IPCW}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{\mathrm{IPCW},i},$} (19)
-or
+Estimating conditional censoring or the conditional survival function can be approached similarly, as both involve estimating a time, whether for survival or censoring.
+Consequently, we can use parametric methods, or semi-parametric methods, such as the Cox model, or non-parametric approaches like survival forests (see Section 2.2.1).
+Once $\widehat{G}(\cdot|A,X)$ is provided, one can construct an estimator of the difference of RMST based on Equation 9,
 $\widehat{\theta}_{\mathrm{IPCW}}=\frac{1}{n_{1}}\sum_{A_{i}=1}T^{*}_{\mathrm{IPCW},i}-\frac{1}{n_{0}}\sum_{A_{i}=0}T^{*}_{\mathrm{IPCW},i}.$} (11)
-where we recall that $n_{a}:=\#\{i\in[n]~|~A_{i}=a\}$.
-By Proposition 2.1, Proposition 2.2 and Proposition 9.3, we easily deduce that $\widehat{\theta}_{\mathrm{IPCW}}$ is asymptotically consistent as soon as $\widehat{G}$ is.
+Alternatively, instead of rescaling by $n_{a}$, the estimator can be stabilized by dividing by the sum of the weights, yielding a Hájek-type normalization that reduces variance without affecting consistency (Hajek1971; Robins2000; Hernan2020).
+By Proposition 2.1 ( (Consistency of CUT as RMST estimator with known $\pi$).), Proposition 2.2 ( (Consistency of CUT as RMST estimator with estimated $\hat{\pi}$).) and Proposition 6.3 ( (IPCW is a Censoring Unbiased Transformation).) (in Appendix A.1.2.
+Proofs of Section 2.2), we easily deduce that $\widehat{\theta}_{\mathrm{IPCW}}$ is asymptotically consistent as soon as $\widehat{G}$ is.
 Corollary 2.1.
-Under Assumptions1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.2 (.Causal survival analysis in Randomized Controlled Trials), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{G}$ is uniformly weakly (resp.
-strongly) consistent then so is $\widehat{\theta}_{\mathrm{IPCW}}$, either as in defined in Equation 19 or in Equation 20.
 This result simply comes from the fact that $\widehat{\theta}_{\mathrm{IPCW}}$ depends continuously on $\widehat{G}$ and that $G$ is lower-bounded (Assumption 2.6 .Conditionally independent censoring).
-Surprisingly, we found limited use of this estimator in the literature, beside its first introduction in Koul et al., (1981).
-This could potentially be explained by the fact that, empirically, we observed that this estimator is highly variable.
-Consequently, we do not explore its properties further and will not include it in the numerical experiments.
-A related and more popular estimator is the IPCW-Kaplan-Meier, defined as follows.
+Surprisingly, we found limited use of this estimator in the literature, beside its first introduction in koul1981.
+A more common variant is the IPCW–Kaplan–Meier (IPCWrobins; IPCW_consistency; IPCW), defined as follows.
+In this approach, censoring weights are dynamically evaluated at each distinct event time, allowing the estimator to update risk sets and weighted failures throughout the follow-up period.
 Definition 2.2.
 (IPCW-Kaplan-Meier) We let again $\widehat{G}(\cdot|X,A)$ be an estimator of the (left limit of) the conditional censoring survival function and we introduce
-$D_{k}^{\mathrm{IPCW}}(a)$ $:=\sum_{i=1}^{n}\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
-$\quad\text{and}\quad N^{\mathrm{IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
+$D_{k}^{\mathrm{IPCW}}(a)$ $:=\sum_{i=1}^{n}\frac{\Delta_{i}^{\tau}}{\widehat{G}(t_{k}|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
+$\quad\text{and}\quad N^{\mathrm{IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\frac{1}{\widehat{G}(t_{k}|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
 be the weight-corrected numbers of deaths and of individuals at risk at time $t_{k}$. The IPCW version of the KM estimator is defined as:
@@
 Like before for the classical KM estimator, this new reweighted KM estimator enjoys good statistical properties.
-We show in Proposition 9.4 (Appendix 9) that the oracle version of this estimator is strongly consistent and asymptotically normal.
+We show in Proposition 6.4 (Appendix A.1.
+Appendix A: Proofs) that the oracle version of this estimator is strongly consistent and asymptotically normal.
 These properties also hold for the estimator using a consistently estimated $\widehat{G}$.
 Corollary 2.2.
-Under Assumptions1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{G}$ is uniformly weakly (resp.
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{G}$ is uniformly weakly (resp.
 strongly) consistent then so is $\widehat{S}_{\mathrm{IPCW}}(t|A=a)$.
-Because the evaluation of $N_{k}^{\textrm{IPCW}}(a)$ now depends on information gathered after time $t_{k}$ (through the computation of the weights), the previous properties on the absence of bias (Proposition 9.1) and on the derivation of the asymptotic variance (Proposition 9.2) do not hold in this setting.
-Whether its bias is exponentially small and whether the asymptotic variance can be derived in a closed form are questions left open.
-We are also not aware of any estimation schemes for the asymptotic variance in this context.
-In the case where we do not have access to the oracle survival function $G$, we can again still achieve consistency if the estimator $\widehat{G}(\cdot|A,X)$ that we provide is consistent.
+Because the evaluation of $N_{k}^{\textrm{IPCW}}(a)$ now depends on information gathered after time $t_{k}$, the previous derivations of bias (Appendix A.1.1.
+Proofs of Section 2.1, proposition 6.1) and asymptotic variance (Appendix A.1.1.
+Proofs of Section 2.1, proposition 6.2) no longer hold.
+Overgaard2025 derived asymptotic variance expressions for related Kaplan–Meier–based inverse probability-of-censoring weighted estimators using influence-function representations.
+While these results theoretically extend to our IPCW–Kaplan–Meier formulation, the resulting expressions remain difficult to interpret or implement in practice.
+Based on our own results, we can nevertheless establish that when $G$ is unknown, consistency still holds provided that the estimator $\widehat{G}(\cdot|A,X)$ is itself consistent.
 The Buckley-James transformation
-One weakness of the IPCW transformation is that it discards all censored data.
-The Buckley-James (BJ) transformation, introduced by Buckley and James, (1979), takes a different path by leaving all uncensored values untouched, while replacing the censored ones by an extrapolated value.
+The Buckley-James (BJ) transformation, introduced by buckley1979, takes a different path by leaving all uncensored values untouched, while replacing the censored ones by an extrapolated value (see Figure 4).
 Formally, it is defined as follows:
@@
 $Q_{S}(t|X,A):=\mathbb{E}[T\wedge\tau|X,A,T\wedge\tau>t]=\frac{1}{S(t|X,A)}\int_{t}^{\tau}S(u|X,A)\mathop{}\!\mathrm{d}u$
-where $S(t|X,A=a):=\mathbb{P}(T^{(a)}>t|X)$ is the conditional survival function. We refer again to Figure 5 for a diagram of this transformation.
-We show in Appendix 9 (Proposition 9.5) that this transformation satisfies the censoring-unbiasedness condition defined in Equation 14.
-Again, the BJ transformation depends on a nuisance parameter (here $Q_{S}(\cdot|X,A)$) that needs to be estimated.
-We again refer to Section 2.2.1 for a brief overview of possible estimation strategies for $Q_{S}$.
-Once provided with an estimator $\widehat{Q}_{S}(\cdot|A,X)$, a very natural estimator of the RMST based on the BJ transformation and on Equation 16 or Equation 17 would be
-$\widehat{\theta}_{\mathrm{BJ}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\pi}-\frac{1-A_{i}}{1-\pi}\right)T^{*}_{\mathrm{BJ},i},$} (23)
-or
+and $S(t|X,A=a):=\mathbb{P}(T^{(a)}>t|X)$ is the conditional survival function.
+This transformation satisfies the censoring-unbiasedness condition defined in Equation 6 (Proposition 6.5 in Appendix A.1.
+Appendix A: Proofs).
+The BJ transformation depends on a nuisance parameter ($Q_{S}(\cdot|X,A)$) that needs to be estimated.
+Estimation strategies described in Section 2.2.1 can be used.
+Once provided with an estimator $\widehat{Q}_{S}(\cdot|A,X)$, a natural estimator of the RMST based on the BJ transformation and Equation 9 is
 $\widehat{\theta}_{\mathrm{BJ}}=\frac{1}{n_{1}}\sum_{A_{i}=1}T^{*}_{\mathrm{BJ},i}-\frac{1}{n_{0}}\sum_{A_{i}=0}T^{*}_{\mathrm{BJ},i}.$} (14)
@@
 Corollary 2.3.
-Under Assumptions1.1 (.Definition of the estimand: the RMST), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{Q}_{S}$ is uniformly weakly (resp.
-strongly) consistent then so is $\widehat{\theta}_{\mathrm{BJ}}$ defined as in Equation 23 or Equation 24.
-The proof is again a mere application of Propositions 2.1, 2.2 and 9.5, and relies on the continuity of $S\mapsto Q_{S}$.
-Additionally, the Buckley-James transformation enjoys a notable optimality property: it minimizes the mean squared error among all square-integrable censoring-unbiased transformations.
-We restate formally as Theorem 9.1 in Appendix 9.
-This result, originally stated without proof in Fan and Gijbels, (1994), is proved there for completeness.
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), if $\widehat{Q}_{S}$ is uniformly weakly (resp.
+strongly) consistent then so is $\widehat{\theta}_{\mathrm{BJ}}$ defined as in Equation 14.
+The proof is an application of Propositions 2.1 ( (Consistency of CUT as RMST estimator with known $\pi$).), 2.2 ( (Consistency of CUT as RMST estimator with estimated $\hat{\pi}$).) and 6.5 ( (BJ is a censoring unbiased transformation).
+in Appendix, Section A.1.2.
+Proofs of Section 2.2), and relies on the continuity of $S\mapsto Q_{S}$.
+Additionally, the Buckley-James transformation enjoys a notable optimality property.
+Theorem 2.1 (BJ minimizes the mean squared error among CUT).
+For any transformation $T^{*}$ of the form 5, it holds
+$\mathbb{E}[(T^{*}_{\mathrm{BJ}}-T\wedge\tau)^{2}]\leqslant\mathbb{E}[(T^{*}-T\wedge\tau)^{2}].$
+This result, originally stated without proof in fan1994, is proved there for completeness in Appendix, Section A.1.2. Proofs of Section 2.2.
 2.2.3 Augmented corrections
 The main disadvantage of the two previous transformations is that they heavily rely on the specification of good estimator $\widehat{G}$ (for IPCW) or $\widehat{S}$ (for BJ).
-In order to circumvent this limitation, Rubin and van der Laan, (2007) proposed the following transformations, whose expression is based on theory of semi-parametric estimation developed in Van der Laan and Robins, (2003),
+In order to circumvent this limitation, rubin2007 proposed the following transformations, whose expression is based on theory of semi-parametric estimation developed in vanderLaan2003,
 $T^{*}_{\mathrm{DR}}=\frac{\Delta^{\tau}\widetilde{T}\wedge\tau}{G(\widetilde{T}\wedge\tau|X,A)}+\frac{(1-\Delta^{\tau})Q_{S}(\widetilde{T}\wedge\tau|X,A)}{G(\widetilde{T}\wedge\tau|X,A)}-\int_{0}^{\widetilde{T}\wedge\tau}\frac{Q_{S}(t|X,A)}{G(t|X,A)^{2}}\mathop{}\!\mathrm{d}\mathbb{P}_{C}(t|X,A),$} (15)
 where $\mathop{}\!\mathrm{d}\mathbb{P}_{C}(t|X,A)$ is the distribution of $C$ conditional on the covariates $X$ and $A$.
-We stress that this distribution is entirely determined by the $G(\cdot|X,A)$, so that this transformation only depends on the knowledge of both conditional survival functions $G$ and $S$, will be thus sometimes denoted $T^{*}_{\mathrm{DR}}(G,S)$ to underline this dependency.
+As this transformation depends on both conditional survival functions $G$ and $S$, we will denote it $T^{*}_{\mathrm{DR}}(G,S)$ to make this dependence explicit.
 Importantly, this transformation not only satisfies the censoring-unbiasedness condition of Equation 6, but also enjoys a double robustness property: it remains valid as long as either $G$ or $S$ is correctly specified.
-The formal statement of this result is given in Proposition 9.6 in Appendix 9.
-The statement and proof of this results is found in Rubin and van der Laan, (2007) in the case where $C$ and $T$ are continuous.
+The formal statement of this result is given in Proposition 6.6 in Appendix A.1.
+Appendix A: Proofs.
+The statement and proof of this results is found in rubin2007 in the case where $C$ and $T$ are continuous.
 3 Causal survival analysis in observational studies
-Unlike RCT, observational data — such as from registries, electronic health records, or national healthcare systems — are collected without controlled randomized treatment allocation.
+Unlike RCT, observational data (such as from registries, electronic health records, or national healthcare systems) are collected without controlled randomized treatment allocation.
 In such cases, treated and control groups are likely unbalanced due to the non-randomized design, which results in the treatment effect being confounded by variables influencing both the survival outcome $T$ and the treatment allocation $A$.
@@
 Assumption 3.1.
-(Conditional Exchangeability / Unconfoundedness)
-It holds
-$A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X$
+(Conditional Exchangeability / Unconfoundedness) It holds $A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X$
 Under Assumption 3.1, the treatment assignment is randomly assigned conditionally on the covariates $X$.
-This assumption ensures that there are no unmeasured confounders as the latter would make it impossible to distinguish correlation from causality.
+This assumption states that there are no unmeasured confounders as the latter would make it impossible to distinguish correlation from causality.
 Assumption 3.2.
-(Positivity / Overlap for Treatment)
-Letting $e(X):=\mathbb{P}(A=1\mid X)$ be the propensity score, there holds
-$0<e(X)<1\quad\text{almost surely.}$
-The Assumption 3.2 requires adequate overlap in covariate distributions between treatment groups, meaning every observation must have a non-zero probability of being treated.
+(Positivity / Overlap for Treatment) Letting $e(X):=\mathbb{P}(A=1\mid X)$ be the propensity score, there holds $0<e(X)<1\quad\text{almost surely.}$
+The Assumption 3.2 requires adequate overlap in covariate distributions between treatment groups, meaning every observation must have a non-zero probability of being treated and not being treated.
 Because Assumption 2.1 (.Causal survival analysis in Randomized Controlled Trials) no longer holds anymore, neither does the previous identifiability Equation 2.
@@
 $=\mathbb{E}\left[\mathbb{E}[T^{(1)}\wedge\tau|X]-\mathbb{E}[T^{(0)}\wedge\tau|X]\right]$
-$=\mathbb{E}\left[\mathbb{E}[T^{(1)}\wedge\tau|X,A=1]-\mathbb{E}[T^{(0)}\wedge\tau|X,A=0]\right]$ (unconfoundness)
+$=\mathbb{E}\left[\mathbb{E}[T^{(1)}\wedge\tau|X,A=1]-\mathbb{E}[T^{(0)}\wedge\tau|X,A=0]\right]$ (Unconfoundedness)
 $=\mathbb{E}\left[\mathbb{E}[T\wedge\tau|X,A=1]-\mathbb{E}[T\wedge\tau|X,A=0]\right].$ (SUTVA)
-In another direction, one could wish to identify the treatment effect through the survival curve as in Equation 1:
-$S^{(a)}(t)=\mathbb{P}(T^{(a)}>t)=\mathbb{E}\left[\mathbb{P}(T>t|X,A=a)\right].$} (27)
-Again, both identities still depend on the unknown quantity $T$ and suggest two different estimation strategies. These strategies differ according to the censoring assumptions and are detailed below.
+A direct application of the above equation could be a direct plug-in G-formula estimator developed in Section 2.2.1, equation 4.
+In particular, chen2001 prove consistency and asymptotic normality results for Cox estimators in an observational study, and they give an explicit formulation of the asymptotic variance as a function of the parameters of the Cox model.
+In the non-parametric literature, foster2011 and kunzel2019 empirically study this estimator using survival forests, with the former employing it as a T-learner (referred to as Virtual Twins) and the latter as an S-learner.
+In another direction, the next sections describe alternative estimation strategies that also rely on equation 16 but adapt the estimation of the conditional expectations to the specific censoring assumptions—either independent or conditionally independent censoring.
 3.1 Independent censoring
-Figure 6 illustrates a causal graph in observational survival data with independent censoring (Assumptions 3.1 and 2.3).
+Under Assumptions 3.1 (.Causal survival analysis in observational studies) and 2.3 (.Independent censoring: the Kaplan-Meier estimator), the data exhibit confounding effect induced by the confounders.
+The corresponding causal graph is shown in Figure 5.
 Figure 5: Causal graph in observational survival data with independent censoring.
@@
 Building on Equation 16, we can write
-$S^{(1)}(t)=\mathbb{E}\left[\frac{\mathbb{E}[\mathbb{I}\{A=1,T>t\}|X]}{\mathbb{E}[\mathbb{I}\{A=1\}|X]}\right]=\mathbb{E}\left[\frac{A\mathbb{I}\{T>t\}}{e(X)}\right],$
-which suggests to adapt the classical KM estimator by reweighting it by the propensity score.
-The use of propensity score in causal inference has been initially introduced by Rosenbaum and Rubin, (1983) and further developed in Hirano et al., (2003).
-It was extended to survival analysis by Xie and Liu, (2005) through the adjusted Kaplan-Meier estimator as defined below.
+$\theta_{\mathrm{RMST}}$ $=\mathbb{E}\left[\mathbb{E}[T\wedge\tau|X,A=1]-\mathbb{E}[T\wedge\tau|X,A=0]\right]$ (17)
+$=\mathbb{E}\left[\frac{\mathbb{E}[\mathbb{I}\{A=1\}T\wedge\tau|X]}{\mathbb{E}[\mathbb{I}\{A=1\}|X]}\right]-\mathbb{E}\left[\frac{\mathbb{E}[\mathbb{I}\{A=0\}T\wedge\tau|X]}{\mathbb{E}[\mathbb{I}\{A=0\}|X]}\right]$ (Positivity for treatment)
+$=\mathbb{E}\left[\frac{AT\wedge\tau}{e(X)}\right]-\mathbb{E}\left[\frac{(1-A)T\wedge\tau}{1-e(X)}\right],$
+which suggests to adapt the classical KM estimator by reweighting it by the propensity score (see equation 1).
+The use of propensity score in causal inference has been initially introduced by rosenbaum1983 and further developed in hirano2003.
+It was extended to survival analysis by xie2005 through the adjusted Kaplan-Meier estimator as defined below.
 Definition 3.1.
 […]
-$D_{k}^{\mathrm{IPTW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{a}{\widehat{e}(X_{i})}+\frac{1-a}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A_{i}=a)$
-$\quad\text{and}\quad N^{\mathrm{IPTW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{a}{\widehat{e}(X_{i})}+\frac{1-a}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
+$D_{k}^{\mathrm{IPTW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A_{i}=a)$
+$\quad\text{and}\quad N^{\mathrm{IPTW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
 be the numbers of deaths and of individuals at risk at time $t_{k}$, reweighted by the propensity score.
@@
 Although the reweighting slightly changes the analysis, the good properties of the classical KM carry on to this setting.
-In Appendix 9, Proposition 9.7 establishes that $S^{*}_{\mathrm{IPTW}}(t\mid A=a)$ is a strongly consistent and asymptotically normal estimator of the survival function $S^{(a)}(t)$.
+In Appendix A.1.
+Appendix A: Proofs, Proposition 6.7 establishes that $S^{*}_{\mathrm{IPTW}}(t\mid A=a)$ is a strongly consistent and asymptotically normal estimator of the survival function $S^{(a)}(t)$.
 Corollary 3.1 further shows that the same properties hold for the practical estimator provided the estimated propensity score converges uniformly.
@@
 which will be consistent under the same Assumptions as the previous Corollary.
-Note that we are not aware of any formal results concerning the bias and the asymptotic variance of the oracle estimator $S^{*}_{\mathrm{IPTW}}(t|A=a)$, and we refer to Xie and Liu, (2005) for heuristics concerning these questions.
+Note that we are not aware of any formal results concerning the bias and the asymptotic variance of the oracle estimator $S^{*}_{\mathrm{IPTW}}(t|A=a)$, and we refer to Xie and Liu (2005) for heuristics concerning these questions.
 3.2 Conditional independent censoring
-Under Assumptions 3.1 (.Causal survival analysis in observational studies) and 2.5 (.Conditionally independent censoring), the causal effect is affected both by confounding variables and by conditional censoring.
-The associated causal graph is depicted in Figure 7.
-In this setting, one can still use the $G$-formula exactly as in Section 2.2.1.
-A natural alternative approach is to weight the IPCW and BJ transformations from Section 2.2.2 by the propensity score to disentangle both confounding effects and censoring at the same time.
+Under Assumptions 3.1 (.Causal survival analysis in observational studies) and 2.5 (.Conditionally independent censoring), the data exhibit both confounding and conditionally independent censoring.
+The corresponding causal graph is shown in Figure 6.
+To address this, weight the IPCW and Buckley–James (BJ) transformations from Section 2.2.2 by the propensity score.
 Figure 6: Causal graph in observational survival data with dependent censoring.
-We mention that the G-formula approach developed in Section 2.2.1 does work in that context.
-In particular, Chen and Tsiatis, (2001) prove consistency and asymptotic normality results for Cox estimators in an observational study, and they give an explicit formulation of the asymptotic variance as a function of the parameters of the Cox model.
-In the non-parametric world, Foster et al., (2011) and Künzel et al., (2019) empirically study this estimator using survival forests, with the former employing it as a T-learner (referred to as Virtual Twins) and the latter as an S-learner.
 3.2.1 IPTW-IPCW transformations
-One can check that the IPCW transformation as introduced in Equation 18 is also a censoring-unbiased transformation in that context (formally stated in Proposition 9.8).
-Deriving an estimator of the difference in RMST is however slightly different in that context.
+One can check that the IPCW transformation as introduced in Equation 10 is also a censoring-unbiased transformation in observational study (see Appendix in Proposition 6.8).
+Deriving an estimator of the difference of RMST is however slightly different in that context.
 In particular, Equation 7 rewrites
 $\mathbb{E}[\mathbb{E}[T^{*}|X,A=1]]=\mathbb{E}\left[\frac{A}{e(X)}T^{*}\right],$
-Which in turn suggests to define
-$\widehat{\theta}_{\mathrm{IPTW-IPCW}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)T^{*}_{\mathrm{IPCW},i}.$} (30)
+which suggests defining an estimator of the RMST difference that can, similarly to IPCW estimators, be stabilized by normalizing with the sum of the weights, and is identified by combining Equation 17 with the IPCW representation in Equation 10.
+$\widehat{\theta}_{\mathrm{IPTW-IPCW}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{e(X_{i})}-\frac{1-A_{i}}{1-e(X_{i})}\right)T^{*}_{\mathrm{IPCW},i}.$} (20)
 This transformation now depends on two nuisance parameters, namely the conditional survival function of the censoring (through $T^{*}_{\mathrm{IPCW}}$) and the propensity score.
 Proposition 3.1 ensures consistency of this estimator when both nuisance quantities are consistently estimated.
-Proposition 3.1.
+Proposition 3.1 (Consistency of $\widehat{\theta}_{\mathrm{IPTW-IPCW}}$ with estimated $\hat{e}$ and $\hat{G}$).
 […]
-We can also use the same strategy as for the IPCW transformation and incorporate the new weights into a Kaplan-Meier estimator.
+We can also use the same strategy as for the IPCW transformation and incorporate the time-dependent weights into a Kaplan-Meier estimator.
 Definition 3.2.
 […]
-$D_{k}^{\mathrm{IPTW-IPCW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
-$\quad\text{and}\quad N^{\mathrm{IPTW-IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
+$D_{k}^{\mathrm{IPTW-IPCW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{\Delta_{i}^{\tau}}{\widehat{G}(t_{k}|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
+$\quad\text{and}\quad N^{\mathrm{IPTW-IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{1}{\widehat{G}(t_{k}|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
 be the weight-corrected numbers of deaths and of individual at risk at time $t_{k}$. The IPTW-IPCW version of the KM estimator is defined as:
@@
 $\widehat{\theta}_{\mathrm{IPTW-IPCW-KM}}=\int_{0}^{\tau}\widehat{S}_{\mathrm{IPTW-IPCW}}(t|A=1)-\widehat{S}_{\mathrm{IPTW-IPCW}}(t|A=0)dt.$} (21)
-In Appendix 9, Proposition 9.9 shows that the oracle version of this estimator is consistent and asymptotically normal.
-The result extends to the practical estimator: as shown in the corollary below (Corollary 3.2), uniform consistency of the nuisance estimators ensures the consistency of the estimated survival curves $\widehat{\theta}_{\mathrm{IPTW-IPCW-KM}}$.
+In Appendix A.1.
+Appendix A: Proofs, Proposition 6.9 shows that the oracle version of this estimator is consistent and asymptotically normal.
+The result extends to the empirical estimator: as shown in Corollary 3.2, uniform consistency of the nuisance estimators ensures the consistency of the estimated survival curves $\widehat{\theta}_{\mathrm{IPTW-IPCW-KM}}$.
 Corollary 3.2.
-Under Assumptions 1.1 (.Definition of the estimand: the RMST), 3.1 (.Causal survival analysis in observational studies), 3.2 (.Causal survival analysis in observational studies), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), and for all $t\in[0,\tau]$, if the nuisance estimators $\widehat{G}(\cdot|A,X)$ and $\widehat{e}$ are weakly (resp.
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 3.1 (.Causal survival analysis in observational studies), 3.2 (.Causal survival analysis in observational studies), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), and for all $t\in[0,\tau]$, if the nuisance estimators $\widehat{G}(\cdot|A,X)$ and $\widehat{e}$ are weakly (resp.
 strongly) uniformly consistent, then $\widehat{S}_{\mathrm{IPTW-IPCW}}(t|A=a)$ is a weakly (resp.
@@
 We are not aware of general formulas for the asymptotic variances in this context.
-We mention nonetheless that Schaubel and Wei, (2011) have been able to derive asymptotic laws in this framework in the particular case of Cox-models.
+We mention nonetheless that schaubel2011 have been able to derive asymptotic laws in this framework in the particular case of Cox-models.
 3.2.2 IPTW-BJ transformations
-Like IPCW transformation, BJ transformation is again a censoring-unbiased transformation in an observational context. This result is formalized in Proposition 9.10 in Appendix 9.
+As with IPCW, the BJ transformation in Equation 13 is censoring-unbiased in observational contexts (Proposition 6.10, Appendix A.1.
+Appendix A: Proofs).
 The corresponding estimator of the difference in RMST is
-$\hat{\theta}_{\mathrm{IPTW-BJ}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)T^{*}_{\mathrm{BJ},i}.$} (32)
+$\hat{\theta}_{\mathrm{IPTW-BJ}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{e(X_{i})}-\frac{1-A_{i}}{1-e(X_{i})}\right)T^{*}_{\mathrm{BJ},i}.$} (22)
 This transformation depends on the conditional survival function $S$ (through $T^{*}_{\mathrm{BJ}}$) and the propensity score.
 Proposition 3.2 stated that consistency of the nuisance parameter estimators implies again consistency of the RMST estimator.
-Proposition 3.2.
+Proposition 3.2 (Consistency of $\widehat{\theta}_{\mathrm{IPTW-BJ}}$ with estimated $\hat{e}$ and $\hat{S}$).
 […]
 3.2.3 Double augmented corrections
-Building on the classical doubly-robust AIPTW estimator from causal inference (Robins et al.,, 1994), we could incorporate the doubly-robust transformations for censoring of Section 2.2.3 to obtain a doubly robust transformation for censoring and treatment assignment
+Building on Equations 16 and 17, it follows that
+$\mathbb{E}\!\left[\frac{\mathbb{I}\{A=a\}}{\mathbb{E}[\mathbb{I}\{A=a\}\mid X]}\,\Big((T\wedge\tau)-\mathbb{E}[T\wedge\tau\mid X,A=a]\Big)\,\middle|\,X\right]=0.$
+Thus,
+$\theta_{\mathrm{RMST}}$ $=\mathbb{E}\left[\mathbb{E}[T\wedge\tau\mid X,A=1]-\mathbb{E}[T\wedge\tau\mid X,A=0]\right]$ (23)
+$=\mathbb{E}\left[\mathbb{E}[T\wedge\tau\mid X,A=1]-\mathbb{E}[T\wedge\tau\mid X,A=0]\right$.
+$\qquad\left.+\underbrace{\frac{\mathbb{I}\{A=1\}}{e(X)}\Big(T\wedge\tau-\mathbb{E}[T\wedge\tau\mid X,A=1]\Big)-\frac{\mathbb{I}\{A=0\}}{1-e(X)}\Big(T\wedge\tau-\mathbb{E}[T\wedge\tau\mid X,A=0]\Big)}_{\text{\small add mean-zero terms}}\right].$
+This representation coincides with the classical doubly robust augmented inverse probability weighting (AIPTW) functional from causal inference (robins1994).
+Incorporating the doubly robust censoring transformations from Section 2.2.3 yields a transformation that is doubly robust with respect to both censoring and treatment assignment
 $\Delta^{*}_{\mathrm{QR}}=\Delta^{*}_{\mathrm{QR}}(G,S,\mu,e):=\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)(T^{*}_{\mathrm{DR}}(G,S)-\mu(X,A))+\mu(X,1)-\mu(X,0),$
@@
 That is, among the four nuisance functions involved, only one from each pair $(S,G)$ and $(e,\mu)$ needs to be correctly specified.
-This property (similar to Ozenne et al., (2020), Thm 1) is formally stated in Proposition 9.11 in Appendix 9.
+This property (similar to Ozenne20_AIPTW_AIPCW, Thm 1) is formally stated in Proposition 6.11 in Appendix A.1.
+Appendix A: Proofs.
 […]
@@
 $=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}-\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)(T^{*}_{\mathrm{DR}}(\widehat{G},\widehat{S})_{i}-\widehat{\mu}(X_{i},A_{i}))+\widehat{\mu}(X_{i},1)-\widehat{\mu}(X_{i},0).$
-This estimator enjoys good asymptotic properties under parametric models, as detailed in Ozenne et al., (2020)
-4 Variance estimation
-Some survival estimators presented in this work have known asymptotic variance expressions, such as the Kaplan-Meier estimator (via Greenwood’s formula) or weighted Kaplan-Meier estimators.
-Under assumptions of asymptotic normality and independence between the two survival curves, the variance of the RMST difference can then be straightforwardly obtained through the delta method.
-However, in practical situations, the covariance between estimated survival curves is typically nonzero, complicating variance estimation.
-While explicit variance formulas have been derived in certain settings—such as for the G-formula combined with semiparametric estimators (see Chen and Tsiatis, (2001))—these expressions often remain complex and difficult to implement.
-Therefore, we recommend using the nonparametric bootstrap (Efron and Tibshirani,, 1994) as a robust and practical solution for variance estimation and confidence interval construction in general settings.
-The effectiveness of bootstrap methods in the context of censored data has been investigated in several studies, including Efron, (1981) and Chen and Lo, (1996), which provide both theoretical justification and empirical validation for its use with Kaplan–Meier estimator.
-5 Implementation
-In this section, we first summarize the various estimators and their properties.
-We then provide custom implementations for all estimators, even those available in existing packages.
-These manual implementations serve two purposes: first, to make the methods accessible to the community when no existing implementation is available; and second, to facilitate a deeper understanding of the methods by detailing each step, even when a package solution exists.
-Finally, we present the packages available for directly computing $\theta_{\mathrm{RMST}}$.
-5.1 Summary of the estimators
-Table LABEL:tbl-nuisance provides an overview of the estimators introduced in this paper, along with the corresponding nuisance parameters needed for their estimation and an overview of their statistical properties in particular regarding their sensitivity to misspecification of the nuisance parameters.
+Asymptotic properties of this estimator, such as asymptotic normality and semiparametric efficiency, are detailed in Ozenne20_AIPTW_AIPCW.
+4 Summary of the estimators and variance estimation
+For completeness, Table LABEL:tbl-nuisance offers a compact synopsis of the estimators, context of application, associated nuisance models, and large-sample behavior, with a focus on misspecification sensitivity.
 Table 3: Estimators of the RMST difference and nuisance parameters needed to compute each estimator.
 Empty boxes indicate that the nuisance parameter is not needed in the estimator thus misspecification has no effect.
-Estimators in italic are those that are already implemented in available packages.
 Estimator RCT Obs Ind Cens Dep Cens Outcome model Censoring model Treatment model Robustness
@@
 […]
-5.2 Implementation of the estimators
-Across different implementations, we use the following shared functions provided in the utilitary.R file.
-  • estimate_propensity_score: function to estimate propensity scores $e(X)$ using either parametric (i.e.
-logistic regression with the argument type_of_model = "glm") or non-parametric methods (i.e.
-probability forest with the argument type_of_model ="probability forest" based on the probability_forest function from the grf (Tibshirani et al.,, 2017) package).
-This latter can include cross-fitting (n.folds 1).
-  • estimate_survival_function: function to estimate the conditional survival model, which supports either Cox models (argument type_of_model = "cox") or survival forests (argument type_of_model = "survival forest") which uses the function survival_forest from the grf (Tibshirani et al.,, 2017) package.
-This latter can include cross-fitting (n.folds 1).
-The estimation can be done with a single learner (argument learner = "S-learner") or two learners (argument learner = "T-learner").
-  • estimate_hazard_function: function to estimate the instantaneous hazard function by deriving the cumulative hazard function at each time point.
-This cumulative hazard function is estimated from the negative logarithm of the survival function.
-  • Q_t_hat: function to estimate the remaining survival function at all time points and for all individuals which uses the former estimate_survival_function.
-  • Q_Y: function to select the value of the remaining survival function from Q_t_hat at the specific time-to-event.
-  • integral_rectangles: function to estimate the integral of a decreasing step function using the rectangle method.
-  • expected_survival: function to estimate the integral with $x,y$ coordinate (estimated survival function) using the trapezoidal method.
-  • integrate: function to estimate the integral at specific time points Y.grid of a given integrand function which takes initially its values on times using the trapezoidal method.
+Some survival estimators presented in this work have known asymptotic variance expressions, such as the Kaplan-Meier estimator (via Greenwood’s formula, derived in Proposition 6.2) or AIPTW-AIPCW estimator (derived in Ozenne20_AIPTW_AIPCW).
+Under standard regularity conditions ensuring asymptotic normality and assuming independence between the two survival curves, the variance of the RMST difference can be readily obtained via the delta method.
+However, in practice the two estimated survival curves are typically correlated, which complicates variance estimation.
+While explicit variance formulas have been derived in certain settings (such as for the G-formula combined with Cox estimators (see chen2001)) these expressions often remain complex and difficult to implement.
+We therefore recommend the nonparametric bootstrap (efron1994) as a robust and practical solution for variance estimation and confidence interval construction in general settings.
+The effectiveness of bootstrap methods in the context of censored data has been investigated in several studies, including efron1981 and chen1996, which provide both theoretical justification and empirical validation for its use with Kaplan–Meier estimator.
+5 Simulations
+5.1 Implementation
+Maintained implementations for estimating RMST under right censoring remain limited.
+Available tools include survRM2 (Hajime2015), grf (Tibshirani2017), and RISCA (Foucher2019).
+The Appendix details these packages and their functions, mapping each to the estimators studied in this paper.
+We implement the remaining estimators and, for a unified and pedagogical framework, provide custom implementations for all estimators, including those with existing software.
+The complete codebase is available on GitHub at github.com/Sanofi-Public/causal_survival_analysis , with function-level usage documented in the Appendix A.2.
+Appendix B: Implementations.
+Building on these implementations, we compare the behaviors and performances of the estimators through simulations conducted across various experimental contexts.
+These contexts include scenarios based on RCTs and observational data, with both independent and dependent censoring (see Appendix A.3.
+Appendix C: Data Generating Process and Figures for details about the data generating process).
+For each setting, we generated 100 independent datasets, computed the estimator on each, and summarized the resulting distribution using boxplots.
+5.2 RCT
+We sample $n$ i.i.d.
+observations $(X_{i},A_{i},C_{i},T_{i}^{(0)},T_{i}^{(1)})_{i\in[n]}$ where covariates $X\sim\mathcal{N}(\mu,\Sigma)$, treatment is randomized with $e(X)=\Pr(A=1\mid X)=0.5$, and the potential event times follow Cox models.
+Specifically, the control potential time has hazard $\lambda^{(0)}(t\mid X)=\lambda_{0}\,\exp\{\beta_{0}^{\top}X\},$ and the treated potential time is obtained via a constant additive shift, $T^{(1)}=T^{(0)}+10.$ Censoring is either independent, with a constant hazard $\lambda_{C}(t)=\lambda_{C,0}$ (Scenario 1), or conditionally independent given $X$, with a Cox-type hazard $\lambda_{C}(t\mid X)=\lambda_{C,0}\,\exp\{\beta_{C}^{\top}X\}$ (Scenario 2).
+The truncation horizon is fixed at $\tau=25$.
+For RCT scenarios 1 and 2, the ground truth value of the RMST difference at time $\tau=25$ was computed analytically and found to be approximately 7.1.
+For each scenario, we estimate the RMST difference using the methods summarized in Table LABEL:tbl-nuisance.
+The methods used to estimate the nuisance components are indicated in brackets: either logistic regression or random forests for propensity scores and either Cox models or survival random forests for survival and censoring models.
+A naive estimator where censored observations are simply removed and the survival time is averaged for treated and controls is also provided for a naive baseline.
+Figure 7 shows the distribution of the RMST difference for 100 simulations in Scenario 1 and different sample sizes: 500, 1000, 2000, 4000.
+The true value of $\theta_{\mathrm{RMST}}$ is indicated by red dashed line.
+For sake of clarity, only a subset of estimators with particularly informative or interesting results is displayed.
+Figure 7: ATEs estimation in RCT setting with independent censoring.
+In this setting, and in accordance with the theory, the simplest estimator (unadjusted KM) performs just as well as the others, and presents, as expected, an extremely small bias.
+The naive estimator is biased, as expected, and the bias in the G-formula S-learner arises because the treatment effect is additive $T^{(1)}=T^{(0)}+10$ and violates the assumption that $T$ would follow a Cox model in the variables $(X,A)$.
+However, $T|A=a$ is a Cox-model for $a\in\{0,1\}$, which explains the remarkable performance of G-formula (Cox/T-learners) achieves the lowest finite-sample variance.
+Some forest-based versions also yield nearly unbiased results, although their performance can vary: for instance, the T-learner remains well-behaved, while the S-learner appears more sensitive and may require larger sample sizes to achieve convergence.
+For completeness, we have also evaluated the other estimators (e.g.
+IPTW KM, IPCW, IPCW KM, BJ, IPTW-IPCW, IPTW-IPCW KM, IPTW-BJ, AIPTW-AIPCW, causal survival forest) in Figure 11 (Appendix A.3.
+Appendix C: Data Generating Process and Figures).
+As expected, they yield no convergence gains, exhibit larger finite-sample variance, and rely on unnecessary model specification.
+Figure 8: ATEs estimation in RCT setting with conditionally independent censoring.
+Figure 8 shows the results for the RCT simulation with conditionally independent censoring (Scenario 2), focusing on a subset of estimators.
+The naive estimator is biased, as expected.
+The IPCW (Cox) is slightly biased up to 2,000 observations.
+In contrast, both IPCW KM (Cox) and Buckley–James (Cox) are unbiased even with only 500 observations, with a smaller variance for BJ.
+G-formula (Cox/T-learners) performs well, even with small sample size.
+As in Scenario 1, G-formula (Cox/ S-learner) remains biased.
+The forest-based estimators are biased.
+Notably, all estimators exhibit higher variability compared to Scenario 1 likely due to the higher censoring rate in this scenario and the estimation of censoring model.
+For completeness, we have also evaluated the other estimators (e.g.
+IPTW KM, IPTW-IPCW, IPTW-IPCW KM, IPTW-BJ, AIPTW-AIPCW, causal survival forest) in Figure 12 (Appendix A.3.
+Appendix C: Data Generating Process and Figures).
+Introducing a superfluous treatment model via the propensity score inflates finite-sample variance, with a marked effect at 500 observations.
+5.3 Observational data
+We simulate an observational study under both independent and conditionally independent censoring (as in Scenarios 1–2).
+The data consist of an i.i.d.
+$n$-sample $(X_{i},A_{i},C_{i},T_{i}^{(0)},T_{i}^{(1)})_{i\in[n]}$ generated as in Section 5.2, except that treatment allocation now follows a logistic propensity model, $\operatorname{logit}(e(X))=\beta_{A}^{\top}X,$ so that $e(X)$ is no longer constant.
+The survival data-generating mechanism is unchanged; hence the target parameter $\theta_{\mathrm{RMST}}$ coincides with that of the RCT simulations, namely $7.1$ (see Section 5.2).
+Figure 9 below shows the distribution of the estimators of $\theta_{\mathrm{RMST}}$ for the observational study with independent censoring.
+Figure 9: ATEs estimation in observational setting with independent censoring.
+As expected, the naive estimator is biased while IPTW KM (Log.Reg.) is unbiased.
+The top-performing estimator in this scenario is the G-formula (Cox/T-learners), which remains unbiased even with as few as 500 observations and achieves the lowest variance.
+All estimators that use forests to estimate nuisance parameters are biased across sample sizes from 500 to 8000.
+This setting thus highlights that one should either have an a priori knowledge of the specification of the models or large sample size.
+Figure 10 below shows the distribution of the $\theta_{\mathrm{RMST}}$ estimates for the observational study with conditionally independent censoring.
+Figure 10: ATEs estimation in observational setting with conditionally independent censoring.
+The top-performing estimators in this scenario are G-formula (Cox/T-learners) and AIPCW-AIPTW (Cox & Cox & Log.Reg.), which are unbiased even with 500 observations.
+The G-formula (Cox/T-learners) has the lowest variance.
+Surprisingly, the G-formula (Cox/S-learner) performs quite competitively, showing only a slight bias despite the violation of the proportional hazards assumption.
+IPTW-IPCW KM (Cox & Log.Reg.) and IPTW-IPCW (Cox & Log.Reg.) exhibit different finite-sample behaviors.
+The Kaplan–Meier version is already unbiased with 500 observations, whereas the moment-based estimator requires around 8,000 observations to achieve unbiasedness.
+This pattern mirrors the results observed for IPCW and IPCW KM in the randomized trial scenario with conditionally independent censoring.
+All estimators that use forests to estimate nuisance parameters are biased across sample sizes from 500 to 8000.
+Although Causal Survival Forest (Cui2023) and AIPTW-AIPCW (Forest) are expected to converge, they remain extremely demanding in terms of sample size.
+For completeness, we have also evaluated the other estimators (e.g. KM, IPCW KM, IPCW, BJ, IPTW KM and package versions) in Figure 14 (Appendix A.3. Appendix C: Data Generating Process and Figures).
+5.4 Misspecification of nuisance components
+We generate an observational study with covariate interactions and conditionally independent censoring.
+The objective is to assess the impact of misspecifying nuisance components; specifically, we will use models that omit interactions to estimate these components.
+This approach enables us to evaluate the robustness properties of various estimators.
+In addition, in this setting forest-based methods are expected to behave better.
+We generate an i.i.d.
+sample $(X_{i},A_{i},C_{i},T_{i}^{(0)},T_{i}^{(1)})_{i=1}^{n}$ where $X\sim\mathcal{N}(\mu,\Sigma)$, the potential event times and censoring follow Cox models with interactions, and treatment follows a logistic propensity with interactions.
+Specifically, the control potential time has hazard $\lambda^{(0)}(t\mid X)=\exp\{\beta_{0}^{\top}Z(X)\},$ the treated potential time is obtained via a constant shift, $T^{(1)}=T^{(0)}+0.3,$ censoring follows $\lambda_{C}(t\mid X)=\exp\{\beta_{C}^{\top}Z(X)\},$ and the treatment allocation satisfies $\operatorname{logit}(e(X))=\beta_{A}^{\top}Z(X).$ Under correct specification, the nuisance models use $(X,Z(X))$ as inputs; under misspecification, only $X$ and a strict subset of Z(X) (the main-effect terms) are supplied.
+The true RMST difference at time $1$ is estimated to be $0.21$.
+We first confirm consistency by estimating $\theta_{\mathrm{RMST}}$ at 10,000 observations with correctly specified nuisance models and assess robustness under (i) single-model misspecification (Appendix, Figure 15) and (ii) paired or more misspecifications—(outcome, censoring), (outcome, treatment), (treatment, censoring) and (outcome, treatment, censoring) (Appendix, Figure 16).
+Given that, in the parametric setting, IPTW-IPCW KM and IPTW-IPCW behaved similarly asymptotically in the previous simulations, we report results for IPTW-IPCW KM only.
+As expected, when exactly one nuisance component is misspecified (censoring, treatment, or outcome; see Appendix, Figure 15), AIPTW–AIPCW (Cox & Cox & Log.Reg) remains unbiased.
+In contrast, the other estimators are biased when any of their nuisance models is misspecified; for example, if the treatment model is misspecified, IPTW-IPCW KM is biased.
+This finding was previously reported in Ozenne20_AIPTW_AIPCW.
+When two nuisance components are misspecified (see Appendix, figure 16), AIPTW-AIPCW (Cox & Cox & Log.Reg.) seems to converge in case where the treatment and censoring models are misspecificed.
+Furthemore, when all nuisance models are misspecified, all estimators exhibit bias.
+6 Conclusion
+Summary of objectives and findings Restricted mean survival time (RMST)-based estimators provide a valuable alternative to hazard ratios, offering interpretable and robust effect estimates that are particularly well-suited for complex, real-world data.
+In this work, we provided a comprehensive overview of causal survival estimators targeting RMST, emphasizing both their theoretical foundations and practical performance.
+As well as implementation (available on a public GitHub repository), we reviewed a wide range of methods—including regression-based models, weighting approaches, or hybrid estimators-and assessed their behavior in both their parametric and non-parametric versions through extensive simulations.
+These scenarios covered diverse settings, including randomized trials and observational data, with varying censoring mechanisms and model specifications.
+Particular attention was given to robustness, sensitivity to misspecification, and estimator variability in finite samples.
+Practical recommendations for RCTs Whether subject to independent or dependent censoring, we advise against overcomplicating estimation strategies.
+Classical estimators like the Kaplan-Meier estimator (with weights for dependent censoring if necessary) remain highly effective in these settings without requiring adjustment on covariates.
+We also recommend using the parametric G-formula.
+Although randomization can make certain covariates seem unnecessary, theoretical results (Karrison2018) and our simulations show that adjusting for prognostic covariates reduces the asymptotic variance of the RMST-difference estimator.
+We also advocate for the use of the Buckley-James (BJ) transformation in this context, which demonstrated excellent performance in our simulations, particularly in reducing bias under independent or dependent censoring without introducing excessive variability.
+Moreover, BJ offers an attractive compromise between ease of implementation and statistical efficiency.
+Practical recommendations for Observational studies Non-robust weighted estimators such as IPTW-IPCW and IPTW-BJ were consistently marked by high variability, particularly when applied to dependent censoring.
+Parametric G-formula estimators yielded the best performance under correct model specification (as in denz2023), showing both low finite sample bias and variance.
+However, their sensitivity to model misspecification remains a limitation.
+Under dependent censoring, we therefore recommend more robust estimators such as AIPTW-AIPCW, which maintain good performance despite their complexity.
+Parametric estimators perform best when correctly specified; yet with interactions or nonlinearities, nonparametric methods can dominate, as they both avoid functional assumptions and automatically learn complex interactions that parametric models would require explicit specification for.
+That said, parametric versions may still be preferable in small to moderate sample sizes, as they exhibited lower bias and variance in our simulations, especially for samples under 4,000 observations.
+In contrast, when sample sizes are sufficiently large, non-parametric estimators become viable alternatives.
+In such cases, we specifically recommend the AIPTW-AIPCW estimator, since the non-parametric G-formula lacks formal asymptotic guarantees and is prone to plug-in bias (chernozhukov2018).
+Limitations It is important to highlight that our simulation design may disadvantage nonparametric methods, as survival and censoring times were generated from exponential distributions and treatment assignments via logistic regression, naturally favoring parametric models.
+Perspectives Future work should address variable selection in causal survival analysis.
+In causal inference, selection serves for first identification to adjust for confounders while avoiding instrumental variables (IVs) that inflate variance (lunceford2004) and second to efficiency—include precision variables predictive of outcome (lunceford2004; rotnitzky2019).
+With survival data, censoring introduces a third class: covariates tied only to censoring; whether to include them (e.g., in IPCW) is unclear, as they may reduce bias in RMST differences yet raise variance.
+Systematic evaluation of confounders, precision variables, IVs, and censoring-only covariates on RMST bias–variance across modeling strategies is warranted.
+Other estimators could be benchmarked such as pseudo-observation approaches (andersen2017; wang2018; vanhage2024) and targeted maximum likelihood estimation (TMLE) (Stitelman2012; Talbot2025).
+Acknowledgement
+The research was supported by Sanofi and INRIA (Institut National de Recherche en Informatique et en Automatique) through the CIFRE program (Convention Industrielle de Formation par la Recherche).
+Charlotte Voinot and Bernard Sebastien are Sanofi employees and may hold shares and/or stock options in the company.
+The work of CB was supported by the Deutsche Foschungsgemeinschaft (German Research Foundation) on the French-German PRCI ANR ASCAI CA 1488/4-1 “Aktive und Batch-Segmentierung, Clustering und Seriation: Grundlagen der KI”.
+Imke Mayer and Julie Josse have nothing to disclose.
+Conflict of Interest
+The authors declare no conflicts of interest.
+Appendix
+A.1. Appendix A: Proofs
+Convergence notions and consistency.
+Let $(\hat{\theta}_{n})_{n\in\mathbb{N}}$ be a sequence of estimators for a parameter $\theta$. We recall three standard notions of convergence, commonly used to define statistical consistency:
+  1. 1.
+    Weak consistency means that $\hat{\theta}_{n}$ converges in probability to $\theta$, i.e.,
+    $\hat{\theta}_{n}\overset{P}{\to}\theta\quad\text{if and only if}\quad\forall\varepsilon>0,\;\mathbb{P}(|\hat{\theta}_{n}-\theta|>\varepsilon)\to 0\quad\text{as }n\to\infty.$
+    This implies that the estimator gets arbitrarily close to the true value with high probability as the sample size increases.
+  2. 2.
+    Strong consistency means that $\hat{\theta}_{n}$ converges almost surely to $\theta$, i.e.,
+    $\hat{\theta}_{n}\overset{\text{a.s.}}{\to}\theta\quad\text{if and only if}\quad\mathbb{P}\left(\lim_{n\to\infty}\hat{\theta}_{n}=\theta\right)=1.$
+    This is a stronger form of convergence, requiring the sequence to converge pointwise on almost every sample path.
+  3. 3.
+    Convergence in distribution (or weak convergence in the sense of probability theory) refers to the convergence of the distribution of $\hat{\theta}_{n}$ to that of a limiting random variable $Z$, not necessarily equal to $\theta$.
+It is denoted:
+    $\hat{\theta}_{n}\overset{d}{\to}Z,$
+    and defined by the condition that for every bounded continuous function $f$,
+    $\mathbb{E}[f(\hat{\theta}_{n})]\to\mathbb{E}[f(Z)].$
+    This notion is typically used in the context of asymptotic distributional results, such as central limit theorems.
+A.1.1. Proofs of Section 2.1
+Note that not all proofs are novel work.
+Consistency results on many survival estimators have been reported multiple times as explained in the main paper.
+For clarity we still recall them.
+We indicate when the proofs are not novel or when similar proofs exist elsewhere.
+When we indicate nothing, this means that we have not found those results in other published work.
+Proposition 6.1 (Consistency of the Kaplan-Meier estimator).
+(Adapted from gill1983, zhou1988)
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials), 2.2 (.Causal survival analysis in Randomized Controlled Trials), 2.3 (.Independent censoring: the Kaplan-Meier estimator) and 2.4 (.Independent censoring: the Kaplan-Meier estimator), and for all $t\in[0,\tau]$, the estimator $\widehat{S}_{\mathrm{KM}}(t|A=a)$ of $S^{(a)}(t)$ is strongly consistent and admits the following bounds for its bias:
+$0\leqslant S^{(a)}(t)-\mathbb{E}[\widehat{S}_{\mathrm{KM}}(t|A=a)]\leqslant O(\mathbb{P}(N_{k}(a)=0)),$
+where $k$ is the greatest time $t_{k}$ such that $t\geqslant t_{k}$.
+The bound we give, although slightly looser, still exhibits the same asymptotic regime.
+In particular, as soon as $S^{(a)}(t)>0$ (and Assumption 2.4 holds), then the bias decays exponentially fast towards 0.
+Proof.
+(Proposition 6.1).
+Consistency is a trivial consequence of the law of large number.
+To show that $\widehat{S}_{\mathrm{KM}}$ is unbiased, let us introduce $\mathcal{F}_{k}$ be the filtration generated by the set of variables
+$\{A_{i},\mathbb{I}\{\widetilde{T}_{i}=t_{j}\},\mathbb{I}\{\widetilde{T}_{i}=t_{j},\Delta_{i}=1\}~|~j\in[k],i\in[n]\}.$
+which corresponds to the known information up to time $t_{k}$, so that $D_{k}(a)$ is $\mathcal{F}_{k}$-measurable but $N_{k}(a)$ is $\mathcal{F}_{k-1}$-measurable.
+One can write that, for $k\geqslant 2$
+$\mathbb{E}[\mathbb{I}\{\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A_{i}=a\}~|~\mathcal{F}_{k-1}]$ $=\mathbb{E}[\mathbb{I}\{\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A_{i}=a\}~|~\mathbb{I}\{\widetilde{T}_{i}\geqslant t_{k}\},A_{i}]$
+$=\mathbb{I}\{A_{i}=a\}\mathbb{E}[\mathbb{I}\{T_{i}=t_{k},C_{i}\geqslant t_{k}\}~|~\mathbb{I}\{T_{i}\geqslant t_{k},C_{i}\geqslant t_{k}\},A_{i}]$
+$=\mathbb{I}\{A_{i}=a\}\mathbb{I}\{C_{i}\geqslant t_{k}\}\mathbb{E}[\mathbb{I}\{T_{i}=t_{k}\}~|~\mathbb{I}\{T_{i}\geqslant t_{k}\},A_{i}]$
+$=\mathbb{I}\{\widetilde{T}_{i}\geqslant t_{k},A_{i}=a\}\left(1-\frac{S^{(a)}(t_{k})}{S^{(a)}(t_{k-1})}\right),$
+where we used that $T_{i}(a)$ is independent from $A_{i}$ by Assumption 2.1 (.Causal survival analysis in Randomized Controlled Trials). We then easily derive from this that
+$\mathbb{E}\left[\left(1-\frac{D_{k}(a)}{N_{k}(a)}\right)\mathbb{I}\{N_{k}(a)>0\}\middle|\mathcal{F}_{k-1}\right]=\frac{S^{(a)}(t_{k})}{S^{(a)}(t_{k-1})}\mathbb{I}\{N_{k}(a)>0\},$
+and then that
+$\mathbb{E}\left[\widehat{S}_{\mathrm{KM}}(t_{k}|A=a)\middle|\mathcal{F}_{k-1}\right]=\frac{S^{(a)}(t_{k})}{S^{(a)}(t_{k-1})}\widehat{S}_{\mathrm{KM}}(t_{k-1}|A=a)+O(\mathbb{I}\{N_{k}(a)=0\}),$
+By induction, we easily find that
+$\mathbb{E}[\widehat{S}_{\mathrm{KM}}(t|A=a)]=\prod_{t_{j}\leqslant t}\frac{S^{(a)}(t_{j})}{S^{(a)}(t_{j-1})}+O\left(\sum_{t_{j}\leqslant t}\mathbb{P}(N_{j}(a)=0)\right)=S^{(a)}(t)+O(\mathbb{P}(N_{k}(a)=0))$
+where $t_{k}$ is the greatest time such that $t_{k}\leqslant t$. ∎
+Proposition 6.2 (Variance of the Kaplan-Meier estimator).
+(Adapted from Kaplan1958_1)
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials), 2.2 (.Causal survival analysis in Randomized Controlled Trials) , 2.3 (.Independent censoring: the Kaplan-Meier estimator) and 2.4 (.Independent censoring: the Kaplan-Meier estimator), and for all $t\in[0,\tau]$, $\widehat{S}_{\mathrm{KM}}(t|A=a)$ is asymptotically normal and $\sqrt{n}\left(\widehat{S}_{\mathrm{KM}}(t|A=a)-S^{(a)}(t)\right)$ converges in distribution towards a centered Gaussian of variance
+$V_{\mathrm{KM}}(t|A=a):=S^{(a)}(t)^{2}\sum_{t_{k}\leqslant t}\frac{1-s_{k}(a)}{s_{k}(a)r_{k}(a)},$
+where $s_{k}(a)=S^{(a)}(t_{k})/S^{(a)}(t_{k-1})$ and $r_{k}(a)=\mathbb{P}(\widetilde{T}\geqslant t_{k},A=a)$.
+The proof of Proposition 6.2 is derived below.
+Because $D_{k}(a)/N_{k}(a)$ is a natural estimator of $1-s_{k}(a)$ and, $\frac{1}{n}N_{k}(a)$ a natural estimator for $r_{k}(a)$, the asymptotic variance of the Kaplan-Meier estimator can be estimated with the so-called Greenwood formula, as already derived heuristically in Kaplan1958_1:
+$\widehat{\mathrm{Var}}\left(\widehat{S}_{\mathrm{KM}}(t|A=a)\right):=\widehat{S}_{\mathrm{KM}}(t|A=a)^{2}\sum_{t_{k}\leqslant t}\frac{D_{k}(a)}{N_{k}(a)(N_{k}(a)-D_{k}(a))}.$} (25)
+Proof.
+(Proposition 6.2).
+The asymptotic normality is a mere consequence of the joint asymptotic normality of $(N_{k}(a),D_{k}(a))_{t_{k}\leqslant t}$ with an application of the $\delta$-method.
+To access the asymptotic variance, notice that, using a similar reasonning as in the previous proof:
+$\mathbb{E}[(1-D_{k}(a)/N_{k}(a))^{2}|\mathcal{F}_{k-1}]$ $=\mathbb{E}[1-D_{k}(a)/N_{k}(a)|\mathcal{F}_{k-1}(a)]^{2}+\frac{1}{N_{k}(a)^{2}}\mathrm{Var}(D_{k}(a)|\mathcal{F}_{k-1})$
+$=s_{k}^{2}(a)+\frac{s_{k}(a)(1-s_{k}(a))}{N_{k}(a)}\mathbb{I}\{N_{k}(a)>0\}+O(\mathbb{I}\{N_{k}(a)=0\}).$
+Now we know that $N_{k}(a)=nr_{k}(a)+\sqrt{n}O_{\mathbb{P}}(1)$, with the $O_{\mathbb{P}}(1)$ having uniformly bounded moments. So that we deduce that
+$\mathbb{E}[(1-D_{k}(a)/N_{k}(a))^{2}|\mathcal{F}_{k-1}]$ $=s_{k}^{2}(a)+\frac{s_{k}(a)(1-s_{k}(a))}{nr_{k}(a)}+\frac{1}{n^{3/2}}O_{\mathbb{P}}(1),$
+where $O_{\mathbb{P}}(1)$ has again bounded moments. Using this identity, we find that
+$n\mathrm{Var}\widehat{S}_{\mathrm{KM}}(t|A=a)$ $=n\left(\mathbb{E}S_{\mathrm{KM}}(t|A=a)^{2}-S^{(a)}(t)^{2}\right)$
+$=nS^{(a)}(t)^{2}\left(\mathbb{E}\left[\prod_{t_{k}\leqslant t}\left(1+\frac{1}{n}\frac{1-s_{k}(a)}{s_{k}(a)r_{k}(a)}+\frac{1}{n^{3/2}}O_{\mathbb{P}}(1)\right)\right]-1\right).$
+Expending the product and using that the $O_{\mathbb{P}}(1)$’s have bounded moments, we finally deduce that
+$\mathbb{E}\left[\prod_{t_{k}\leqslant t}\left(1+\frac{1}{n}\frac{1-s_{k}(a)}{s_{k}(a)r_{k}(a)}+\frac{1}{n^{3/2}}O_{\mathbb{P}}(1)\right)\right]-1=\frac{1}{n}\sum_{t_{k}\leqslant t}\frac{1-s_{k}(a)}{s_{k}(a)r_{k}(a)}+\frac{1}{n^{3/2}}O(1),$
+$n\mathrm{Var}\widehat{S}_{\mathrm{KM}}(t|A=a)=V_{\mathrm{KM}}(t|A=a)+O(n^{-1/2}),$
+which is what we wanted to show. ∎
+A.1.2. Proofs of Section 2.2
+Proposition 6.3 (IPCW is a Censoring Unbiased Transformation).
+Proof.
+(Proposition 6.3). Assumption 2.6 allows the transformation to be well-defined. Furthermore, it holds
+$E[T^{*}_{\mathrm{IPCW}}|A=a,X]$ $=E\left[\frac{\Delta^{\tau}\times\widetilde{T}\wedge\tau}{G(\widetilde{T}\wedge\tau|A,X)}\middle|A=a,X\right]$
+$=E\left[\frac{\Delta^{\tau}\times T^{(a)}\wedge\tau}{G(T^{(a)}\wedge\tau|A,X)}\middle|A=a,X\right]$
+$=E\left[E\left[\frac{\mathbb{I}\{T^{(a)}\wedge\tau\leqslant C\}\times T^{(a)}\wedge\tau}{G(T^{(a)}\wedge\tau|A,X)}\middle|A,X,T^{(a)}\right]\middle|A=a,X\right]$
+$=E\left[T^{(a)}\wedge\tau|A=a,X\right]$
+$=E\left[T^{(a)}\wedge\tau|X\right].$
+We used in the second equality that on the event $\{\Delta^{\tau}=1,A=a\}$, it holds $\widetilde{T}\wedge\tau=T\wedge\tau=T^{(a)}\wedge\tau$.
+We used in the fourth equality that $G(T^{(a)}\wedge\tau|A,X)=E[\mathbb{I}\{T^{(a)}\wedge\tau\leqslant C\}|X,T^{(a)},A]$ thanks to Assumption 2.5, and in the last one that $A$ is independent of $X$ and $T^{(a)}$ thanks to Assumption 2.1.
+∎
+Proposition 6.4 (Consistency of $S^{*}_{\mathrm{IPCW}}(t|A=a)$ with known $G$).
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials), 2.2 (.Causal survival analysis in Randomized Controlled Trials), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), and for all $t\in[0,\tau]$, the oracle estimator $S^{*}_{\mathrm{IPCW}}(t|A=a)$ defined as in Definition 2.2 with $\widehat{G}=G$ is a stronlgy consistent and asymptotically normal estimator of $S^{(a)}(t)$ .
+Proof.
+(Proposition 6.4). Similarly to the computations done in the proof of Proposition 6.3, it is easy to show that, for $t_{k}\leqslant\tau$,
+$\mathbb{E}\left[\frac{\Delta^{\tau}}{G(t_{k}|X,A)}\mathbb{I}(\widetilde{T}=t_{k},A=a)\right]$ $=\mathbb{E}\left[\frac{\mathbb{I}(t_{k}\leqslant C)}{G(t_{k}|X,A)}\mathbb{I}(T^{(a)}=t_{k},A=a)\right],$
+$=\mathbb{P}(A=a)\mathbb{P}(T^{(a)}=t_{k}),$
+and likewise that
+$\mathbb{E}\left[\frac{1}{G(t_{k}|X,A)}\mathbb{I}(\widetilde{T}\geqslant t_{k},A=a)\right]$ $=\mathbb{E}\left[\frac{1}{G(t_{k}|X,A)}\mathbb{I}(T^{(a)}\wedge C\geqslant t_{k},A=a)\right]$
+$=\mathbb{E}\left[\frac{\mathbb{I}(T^{(a)}\geqslant t_{k})\mathbb{I}(A=a)}{G(t_{k}|X,A)}\mathbb{E}\left[\mathbb{I}(C\geqslant t_{k})|X,A\right]\right]$
+$=\mathbb{P}(A=a)\mathbb{P}(T^{(a)}\geqslant t_{k}),$
+so that $\widehat{S}_{\mathrm{IPCW}}(t)$ converges almost surely towards the product limit
+$\prod_{t_{k}\leqslant t}\left(1-\frac{\mathbb{P}(T^{(a)}=t_{k})}{\mathbb{P}(T^{(a)}\geqslant t_{k})}\right)=S^{(a)}(t),$
+yielding strong consistency. Asymptotic normality is straightforward. ∎
+Proposition 6.5 (BJ is a censoring unbiased transformation).
+Proof.
+(Proposition 6.5). There holds
+$\mathbb{E}[T_{\mathrm{BJ}}^{*}|X,A=a]$ $=\mathbb{E}\left[\Delta^{\tau}T^{(a)}\wedge\tau+(1-\Delta^{\tau})\frac{\mathbb{E}[T\wedge\tau\times\mathbb{I}\{T\wedge\tau>C\}|C,A,X]}{\mathbb{P}(T>C|C,A,X)}\middle|X,A=a\right]$
+$=\mathbb{E}[\Delta^{\tau}T^{(a)}\wedge\tau|X]+\underbrace{\mathbb{E}\left[\mathbb{I}\{T\wedge\tau>C\}\frac{\mathbb{E}[T\wedge\tau\times\mathbb{I}\{T\wedge\tau>C\}|C,A,X]}{\mathbb{E}[\mathbb{I}\{T\wedge\tau\geqslant C\}|C,A,X]}\middle|X,A=a\right]}_{(\star)}.$
+Now we easily see that conditionning wrt $X$ in the second term yields
+$(\star)$ $=\mathbb{E}\left[\mathbb{E}[T\wedge\tau\times\mathbb{I}\{T\wedge\tau>C\}|C,A,X]\middle|X,A=a\right]$
+$=\mathbb{E}[(1-\Delta^{\tau})T\wedge\tau|X,A=a]$
+$=\mathbb{E}[(1-\Delta^{\tau})T^{(a)}\wedge\tau|X],$
+ending the proof. ∎
+Proof.
+(Theorem 2.1, (BJ minimizes the mean squared error among CUT).). We let $T^{*}=\Delta^{\tau}\phi_{1}+(1-\Delta^{\tau})\phi_{0}$ be a transformation of the form Equation 5. There holds
+$\mathbb{E}[(T^{*}-T\wedge\tau)^{2}]=\mathbb{E}[\Delta^{\tau}(\phi_{1}-T\wedge\tau)^{2}]+\mathbb{E}[(1-\Delta^{\tau})(\phi_{0}-T\wedge\tau)^{2}].$
+The first term is non negative and is zero for the BJ transformation.
+Since $\phi_{0}$ is a function of $(\widetilde{T},X,A)$ and that $\widetilde{T}=C$ on $\{\Delta^{\tau}=0\}$, the second term can be rewritten in the following way.
+We let $R$ be a generic quantity that does not depend on $\phi_{0}$.
+$\mathbb{E}$ $[(1-\Delta^{\tau})(\phi_{0}-T)^{2}]=\mathbb{E}\left[\mathbb{I}\{T\wedge\tau>C\}\phi_{0}^{2}-2\mathbb{I}\{T\wedge\tau>C\}\phi_{0}T\wedge\tau\right]+R$
+$=\mathbb{E}\left[\mathbb{P}(T\wedge\tau>C|C,A,X)\phi_{0}^{2}-2\mathbb{E}[T\wedge\tau\mathbb{I}\{T\wedge\tau>C\}|C,A,X]\phi_{0}\right]+R$
+$=\mathbb{E}\left[\mathbb{P}(T\wedge\tau>C|C,A,X)\left(\phi_{0}-\frac{\mathbb{E}[T\wedge\tau\mathbb{I}\{T\wedge\tau>C\}|C,A,X]}{\mathbb{P}(T\wedge\tau>C|C,A,X)}\right)^{2}\right]+R.$
+Now the first term in the right hand side is always non-negative, and is zero for the BJ transformation. ∎
+Proposition 6.6 (AIPCW is a censoring-unbiased transformation).
+We let $F,R$ be two conditional survival functions.
+Under Assumptions1.1 (.Notations and definition of the Restricted Mean Survival Time), 2.1 (.Causal survival analysis in Randomized Controlled Trials) , 2.2 (.Causal survival analysis in Randomized Controlled Trials), 2.5 (.Conditionally independent censoring) and if $F$ also satisfies Assumption 2.6 (.Conditionally independent censoring), and if $F(\cdot|X,A)$ is absolutely continuous wrt $G(\cdot|X,A)$, then the transformation $T^{*}_{\mathrm{DR}}=T^{*}_{\mathrm{DR}}(F,R)$ is a censoring-unbiased transformation in the sense of Equation 6 whenever $F=G$ or $R=S$.
+A careful examination of the proofs from D. Rubin and Laan (2007) show that the proof translates straight away to our discrete setting.
+A.1.3. Proofs of Section 3.1
+Proposition 6.7 (Consistency of $S^{*}_{\mathrm{IPTW}}(t|A=a)$ with known $e$).
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 3.1 (.Causal survival analysis in observational studies), 3.2 (.Causal survival analysis in observational studies), 2.3 (.Independent censoring: the Kaplan-Meier estimator) and 2.4 (.Independent censoring: the Kaplan-Meier estimator) The oracle IPTW Kaplan-Meier estimator $S^{*}_{\mathrm{IPTW}}(t|A=a)$ with $\hat{e}=e$ is a strongly consistent and asymptotically normal estimator of $S^{(a)}(t)$.
+The proof of this result simply relies again on the law of large number and the $\delta$-method.
+Proof.
+(Proposition 6.7).
+The fact that it is strongly consistent and asymptotically normal is again a simple application of the law of large number and of the $\delta$-method.
+We indeed find that, for $t_{k}\leqslant\tau$
+$\mathbb{E}\left[\frac{1}{e(X_{i})}\mathbbm{1}\{\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A_{i}=1\}\right]$ $=\mathbb{E}\left[\frac{A_{i}}{e(X_{i})}\mathbbm{1}\{T_{i}=t_{k},C_{i}\geqslant t_{k}\}\right]$
+$=\mathbb{E}\left[\mathbb{E}\left[\frac{A_{i}}{e(X_{i})}\mathbbm{1}\{T_{i}=t_{k},C_{i}\geqslant t_{k}\}\middle|X_{i}\right]\right]$
+$=\mathbb{E}\left[\mathbb{E}\left[\frac{A_{i}}{e(X_{i})}\middle|X_{i}\right]\mathbb{P}(T_{i}=t_{k}|X_{i})\mathbb{P}(C_{i}\geqslant t_{k})\right]$
+$=\mathbb{P}(T_{i}=t_{k})\mathbb{P}(C_{i}\geqslant t_{k}),$
+where we used that $A$ is independent of $T$ conditionnaly on $X$, and that $C$ is independent of everything. Likewise, one would get that
+$\mathbb{E}\left[\frac{1}{e(X_{i})}\mathbbm{1}\{\widetilde{T}_{i}\geqslant t_{k},A_{i}=1\}\right]=$ $=\mathbb{P}(T_{i}\geqslant t_{k})\mathbb{P}(C_{i}\geqslant t_{k}).$
+Similar computations hold for $A=0$, ending the proof. ∎
+A.1.4. Proofs of Section 3.2
+Proposition 6.8 (IPTW-IPCW is a censoring unbiased transformation).
+Proof.
+(Proposition 6.8). On the event $\{\Delta^{\tau}=1,A=1\}$, there holds $\widetilde{T}\wedge\tau=T\wedge\tau=T^{(1)}\wedge\tau$, whence we find that,
+$\mathbb{E}[T^{*}_{\mathrm{IPCW}}|X,A=1]$ $=\mathbb{E}\left[\frac{A}{e(X)}\frac{\mathbb{I}\{T^{(1)}\wedge\tau\leqslant C\}}{G(T^{(1)}\wedge\tau|X,A)}T^{(1)}\wedge\tau\middle|X\right]$
+$=\mathbb{E}\left[\frac{A}{e(X)}\mathbb{E}\left[\frac{\mathbb{I}\{T^{(1)}\wedge\tau\leqslant C\}}{G(T^{(1)}\wedge\tau|X,A)}\middle|X,A,T^{(1)}\right]T^{(1)}\wedge\tau\middle|X\right]$
+$=\mathbb{E}\left[\frac{A}{e(X)}T^{(1)}\wedge\tau\middle|X\right]$
+$=\mathbb{E}\left[T^{(1)}\wedge\tau\middle|X\right],$
+and the same holds on the event $A=0$. ∎
+Proof.
+(Proposition 3.1, (Consistency of $\widehat{\theta}_{\mathrm{IPTW-IPCW}}$ with estimated $\hat{e}$ and $\hat{G}$).).
+By consistency of $\widehat{G}(\cdot|X,A)$ and $\widehat{e}$ and by continuity, it suffices to look at the asymptotic behavior of the oracle estimator
+$\theta^{*}_{\mathrm{IPTW-IPCW}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{e(X_{i})}-\frac{1-A_{i}}{1-e(X_{i})}\right)\frac{\Delta_{i}^{\tau}}{G(\widetilde{T}_{i}\wedge\tau|A_{i},X_{i})}\widetilde{T}_{i}\wedge\tau.$
+The later is converging almost towards its mean, which, following similar computations as in the previous proof, write
+$\mathbb{E}\left[\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)\frac{\Delta^{\tau}}{G(\widetilde{T}\wedge\tau|A,X)}\widetilde{T}\wedge\tau\right]$ $=\mathbb{E}\left[\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)T\wedge\tau\right]$
+$=\mathbb{E}\left[T^{(1)}\wedge\tau\right]-\mathbb{E}\left[T^{(0)}\wedge\tau\right].$
+Proposition 6.9 (Consistency of $S^{*}_{\mathrm{IPTW-IPCW}}(t|A=a)$ with known $e$ and $G$).
+Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 3.1 (.Causal survival analysis in observational studies), 3.2 (.Causal survival analysis in observational studies), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), and for all $t\in[0,\tau]$, if the oracle estimator $S^{*}_{\mathrm{IPTW-IPCW}}(t|A=a)$ defined as in Definition 3.2 with $\widehat{G}(\cdot|A,X)=G(\cdot|A,X)$ and $\widehat{e}=e$ is a strongly consistent and asymptotically normal estimator of $S^{(a)}(t)$.
+Proof.
+(Proposition 6.9).
+Asymptotic normality comes from a mere application of the $\delta$-method, while strong consistency follows from the law of large number and the follozing computations.
+Like for the proof of Proposition 6.4, one find, by first conditionning wrt $X,A,T^{(a)}$, that, for $t_{k}\leqslant\tau$,
+$\mathbb{E}\left[\left(\frac{A}{e(X)}+\frac{1-A}{1-e(X)}\right)\frac{\Delta^{\tau}}{G(t_{k}|A,X)}\mathbb{I}\{\widetilde{T}=t_{k},A=a\}\right]=\mathbb{P}(T^{(a)}=t_{k})$
+and likewise that
+$\mathbb{E}\left[\left(\frac{A}{e(X)}+\frac{1-A}{1-e(X)}\right)\frac{1}{G(t_{k}|A,X)}\mathbb{I}\{\widetilde{T}\geqslant t_{k},A=a\}\right]=\mathbb{P}(T^{(a)}\geqslant t_{k})$
+so that indeed $S^{*}_{\mathrm{IPTW-IPCW}}(t|A=a)$ converges almost surely towards $S^{(a)}(t)$. ∎
+Proposition 6.10 (IPTW-BJ is a censoring-unbiased transformation).
+Proof.
+(Proposition 6.10). We write
+$\mathbb{E}[T^{*}_{\mathrm{IPTW-BJ}}|X,A=1]$ $=\mathbb{E}\left[\frac{A}{e(X)}\Delta^{\tau}\times\widetilde{T}\wedge\tau\middle|X\right]+\mathbb{E}\left[\frac{A}{e(X)}(1-\Delta^{\tau})Q_{S}(\widetilde{T}\wedge\tau|A,X)\middle|X\right].$
+On the event $\{\Delta^{\tau}=1,A=1\}$, there holds $\widetilde{T}\wedge\tau=T\wedge\tau=T^{(1)}\wedge\tau$, whence we find that the first term on the the right hand side is equal to
+$\mathbb{E}\left[\frac{A}{e(X)}\Delta^{\tau}\times\widetilde{T}\wedge\tau\middle|X\right]$ $=\mathbb{E}\left[\frac{A}{e(X)}\Delta^{\tau}\times T^{(1)}\wedge\tau\middle|X\right]$
+$=\mathbb{E}\left[\Delta^{\tau}\times T^{(1)}\wedge\tau\middle|X\right].$
+For the second term in the right hand side, notice that on the event $\{\Delta^{\tau}=0,A=1\}$, there holds $\widetilde{T}=C<T^{(1)}\wedge\tau$, so that
+$\mathbb{E}$ $\left[\frac{A}{e(X)}\mathbb{I}\{C<T^{(1)}\wedge\tau\}\frac{\mathbb{E}[T^{(1)}\wedge\tau\times\mathbb{I}\{C<T^{(1)}\wedge\tau\}|X,A,C]}{\mathbb{P}(C<T^{(1)}\wedge\tau|C,X,A)}\middle|X\right]$
+$=\mathbb{E}\left[\frac{A}{e(X)}\mathbb{E}[T^{(1)}\wedge\tau\times\mathbb{I}\{C<T^{(1)}\wedge\tau\}|X,A,C]\middle|X\right]$
+$=\mathbb{E}\left[T^{(1)}\wedge\tau\times\mathbb{I}\{C<T^{(1)}\wedge\tau\}\middle|X\right]$
+$=\mathbb{E}\left[(1-\Delta^{\tau})T^{(1)}\wedge\tau\middle|X\right],$
+and the sane holds on the event $\{A=0\}$, which ends the proof. ∎
+Proof.
+(Proposition 3.2, (Consistency of $\widehat{\theta}_{\mathrm{IPTW-BJ}}$ with estimated $\hat{e}$ and $\hat{S}$).).
+By consistency of $\widehat{G}(\cdot|X,A)$ and $\widehat{e}$ and by continuity, it suffices to look at the asymptotic behavior of the oracle estimator
+$\theta^{*}_{\mathrm{IPTW-BJ}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{e(X_{i})}-\frac{1-A_{i}}{1-e(X_{i})}\right)\left(\Delta_{i}^{\tau}\times\widetilde{T}_{i}\wedge\tau+(1-\Delta_{i}^{\tau})Q_{S}(\widetilde{T}_{i}\wedge\tau|A_{i},X_{i})\right).$
+The later is converging almost towards its mean, which, following similar computations as in the previous proof, is simply equal to the RMST difference. ∎
+Proposition 6.11 (Double/Triple robustness of $\Delta^{*}_{\mathrm{QR}}$).
+(adapted from Ozenne20_AIPTW_AIPCW)
+Let $F,R$ be two conditional survival function, $p$ be a propensity score, and $\nu$ be a conditional response.
+Then, under the same assumption on $F,R$ as in Proposition 6.6, and under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 3.1 (.Causal survival analysis in observational studies), 3.2 (.Causal survival analysis in observational studies), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), the transformations $\Delta^{*}_{\mathrm{QR}}=\Delta^{*}_{\mathrm{QR}}(F,R,p,\nu)$ satisfies, fo $a\in{0,1}$,
+$\mathbb{E}[\Delta^{*}_{\mathrm{QR}}|X]=\mathbb{E}[T^{(1)}\wedge\tau-T^{(0)}\wedge\tau|X]\quad\text{if}\quad\begin{cases}F=G\quad&\text{or}\quad R=S\quad\text{and}\\ p=e\quad&\text{or}\quad\nu=\mu.\end{cases}$
+Proof.
+(Proposition 6.11). We can write that
+$\Delta^{*}_{\mathrm{QR}}=\underbrace{\frac{A}{p(X)}(T^{*}_{\mathrm{DR}}(F,R)-\nu(X,1))+\nu(X,1)}_{(\mathrm{A})}-\left(\underbrace{\frac{1-A}{1-p(X)}(T^{*}_{\mathrm{DR}}(F,R)-\nu(X,0))+\nu(X,0)}_{(\mathrm{B})}\right).$
+Focusing on term $(\mathrm{A})$, we easily find that
+$\mathbb{E}[(\mathrm{A})|X]$ $=\mathbb{E}\left[\frac{A}{p(X)}(T^{*}_{\mathrm{DR}}(F,R)-\nu(X,1))+\nu(X,1)\middle|X\right]$
+$=\frac{e(X)}{p(X)}(\mu(X,1)-\nu(X,1))+\nu(X,1).$
+Where we used that $T^{*}_{\mathrm{DR}}(F,R)$ is a censoring-unbiased transform when $F=G$ or $R=S$. Now we see that if $p(X)=e(X)$, then
+$\mathbb{E}[(\mathrm{A})|X]=\mu(X,1)-\nu(X,1)+\nu(X,1)=\mu(X,1),$
+and if $\nu(X,1)=\mu(X,1)$, then
+$\mathbb{E}[(\mathrm{A})|X]=\frac{e(X)}{p(X)}\times 0+\mu(X,1)=\mu(X,1).$
+Likewise, we would show that $\mathbb{E}[(\mathrm{B})|X]=\mu(X,0)$ under either alternative, ending the proof. ∎
+A.2. Appendix B: Implementations
+A.2.1. Available packages
+SurvRM2
+The RMST difference with Unadjusted Kaplan-Meier $\hat{\theta}_{KM}$ (Equation 3) can be obtained using the function rmst2 which requires as input the observed time-to-event, the status, the arm which corresponds to the treatment and $\tau$.
+RISCA
+The RISCA package (Foucher2019) provides several methods for estimating $\theta_{\mathrm{RMST}}$.
+The RMST difference with Unadjusted Kaplan-Meier $\hat{\theta}_{KM}$ (Equation 3) can be derived using the survfit function from the survival package (Therneau2001) which estimates Kaplan-Meier survival curves for treated and control groups, and then the rmst function calculates the RMST by integrating these curves, applying the rectangle method (type=“s”), which is well-suited for step functions.
+The IPTW Kaplan-Meier (Equation 18) can be applied using the ipw.survival and rmst functions.
+The ipw.survival function requires user-specified weights (i.e.
+propensity scores).
+To streamline this process, we define the RISCA_iptw function, which combines these steps and utilizes the estimate_propensity_score from the utilitary.R file.
+A single-learner version of the G-formula, as introduced in Section 2.2.1, can be implemented using the gc.survival function.
+This function requires as input the conditional survival function which should be estimated beforehand with a Cox model via the coxph function from the survival package (Therneau2001).
+Specifically, the single-learner approach applies a single Cox model incorporating both covariates and treatment, rather than separate models for each treatment arm.
+We provide a function RISCA_gf that consolidates these steps.
+grf
+The grf package (Tibshirani2017) enables estimation of the difference between RMST using the Causal Survival Forest approach (Cui2023), which extends the non-parametric causal forest framework to survival data.
+The RMST can be estimated with the causal_survival_forest function, requiring covariates $X$, observed event times, event status, treatment assignment, and the time horizon $\tau$ as inputs.
+The average_treatment_effect function then evaluates the treatment effect based on predictions from the fitted forest.
+A.2.2. Implementation of the estimators
+The complete codebase is available on github.com/Sanofi-Public/causal_survival_analysis
 Unadjusted Kaplan-Meier
-Although Kaplan-Meier is implemented in the survival package (Therneau,, 2001), we provide a custom implementation, Kaplan_meier_handmade, for completeness.
+Although Kaplan-Meier is implemented in the survival package (Therneau2001), we provide a custom implementation, Kaplan_meier_handmade, for completeness.
 The difference in Restricted Mean Survival Time, estimated using Kaplan-Meier as in Equation 3 can then be calculated with the RMST_1 function.
-As an alternative, one can also use the survfit function in the survival package (Therneau,, 2001) for Kaplan-Meier and specify the rmean argument equal to $\tau$ in the corresponding summary function:
+As an alternative, one can also use the survfit function in the survival package (Therneau2001) for Kaplan-Meier and specify the rmean argument equal to $\tau$ in the corresponding summary function:
 IPCW Kaplan-Meier
@@
 The survival censoring function $G(t|X)$ is computed with the estimate_survival_function utility function from the utilitary.R file.
-One could also use the survfit function in the survival package (Therneau,, 2001) by adding IPCW weights for treated and control group and specify the rmean argument equal to $\tau$ in the corresponding summary function:
+One could also use the survfit function in the survival package (Therneau2001) by adding IPCW weights for treated and control group and specify the rmean argument equal to $\tau$ in the corresponding summary function:
 This alternative approach for IPCW Kaplan-Meier would also be valid for IPTW and IPTW-IPCW Kaplan-Meier.
@@
 In g_formula_T_learner, separate models estimate survival curves for treated and control groups, whereas g_formula_S_learner uses a single model incorporating both covariates and treatment status to estimate survival time.
-The latter approach is also available in the RISCA package (Foucher et al.,, 2019) but is limited to Cox models.
+The latter approach is also available in the RISCA package (Foucher2019) but is limited to Cox models.
 IPTW-IPCW Kaplan-Meier
@@
 The AIPTW_AIPCW function implements the AIPTW_AIPCW estimator in Equation 24 using the utilitary function from the utilitary.R file estimate_propensity_score, Q_t_hat, Q_Y, and estimate_survival_function to estimate the nuisance parameters.
-5.3 Available packages
-Currently, there are limited number of maintained implementations available for estimating RMST in the presence of right censoring.
-Among the available tools, noteworthy packages include survRM2 (Hajime et al.,, 2015), grf (Tibshirani et al.,, 2017) and RISCA (Foucher et al.,, 2019).
-These implementations are used and wrapped in our provided utilitary.R code files.
-SurvRM2
-The RMST difference with Unadjusted Kaplan-Meier $\hat{\theta}_{KM}$ (Equation 9) can be obtained using the function rmst2 which requires as input the observed time-to-event, the status, the arm which corresponds to the treatment and $\tau$.
-RISCA
-The RISCA package (Foucher et al.,, 2019) provides several methods for estimating $\theta_{\mathrm{RMST}}$.
-The RMST difference with Unadjusted Kaplan-Meier $\hat{\theta}_{KM}$ (Equation 9) can be derived using the survfit function from the survival package (Therneau,, 2001) which estimates Kaplan-Meier survival curves for treated and control groups, and then the rmst function calculates the RMST by integrating these curves, applying the rectangle method (type=“s”), which is well-suited for step functions.
-The IPTW Kaplan-Meier (Equation 28) can be applied using the ipw.survival and rmst functions.
-The ipw.survival function requires user-specified weights (i.e.
-propensity scores).
-To streamline this process, we define the RISCA_iptw function, which combines these steps and utilizes the estimate_propensity_score from the utilitary.R file.
-A single-learner version of the G-formula, as introduced in Section 2.2.1, can be implemented using the gc.survival function.
-This function requires as input the conditional survival function which should be estimated beforehand with a Cox model via the coxph function from the survival package (Therneau,, 2001).
-Specifically, the single-learner approach applies a single Cox model incorporating both covariates and treatment, rather than separate models for each treatment arm.
-We provide a function RISCA_gf that consolidates these steps.
-grf
-The grf package (Tibshirani et al.,, 2017) enables estimation of the difference between RMST using the Causal Survival Forest approach (Cui et al.,, 2023), which extends the non-parametric causal forest framework to survival data.
-The RMST can be estimated with the causal_survival_forest function, requiring covariates $X$, observed event times, event status, treatment assignment, and the time horizon $\tau$ as inputs.
-The average_treatment_effect function then evaluates the treatment effect based on predictions from the fitted forest.
-6 Simulations
-We compare the behaviors and performances of the estimators through simulations conducted across various experimental contexts.
-These contexts include scenarios based on RCTs and observational data, with both independent and dependent censoring.
-We also use semi-parametric and non-parametric models to generate censoring and survival times, as well as logistic or nonlinear models to simulate treatment assignment.
-For each setting, we generated 100 independent datasets, computed the estimator on each, and summarized the resulting distribution using boxplots.
-6.1 RCT
-Data Generating Process
-We generate RCTs with independent censoring (Scenario 1) and conditionally independent censoring (Scenario 2).
-We sample $n$ iid datapoints $(X_{i},A_{i},C,T_{i}(0),T_{i}(1))_{i\in[n]}$ where $T_{i}(0),T_{i}(1)$ and $C$ follows Cox models.
-More specifically, we set
-  • $X\sim\mathcal{N}\left(\mu,\Sigma\right)$ where $\mu=(1,1,-1,1)$ and $\Sigma=\mathrm{Id}_{4}$.
-  • The hazard function of $T^{(0)}$ is
-    $\lambda^{(0)}(t|X)=0.01\exp\left\{\beta_{0}^{\top}X\right\}\quad\text{where}\quad\beta_{0}=(0.5,0.5,-0.5,0.5).$
-  • The survival times in the treatment group are given by $T^{(1)}=T^{(0)}+10$.
-  • The hazard function of the censoring time $C$ is simply taken as $\lambda_{C}(t|X)=0.03$ in Scenario 1, and in Scenario 2 as
-    $\lambda_{C}(t|X)=0.03\cdot\exp\left\{\beta_{C}^{\top}X\right\}\quad\text{where}\quad\beta_{C}=(0.7,0.7,-0.25,-0.1).$
-  • The treatment allocation is independent of $X$: $e(X)=0.5$.
-  • The threshold time $\tau$ is set to 25.
-The descriptive statistics of the two datasets are displayed in Annex (Section 10.1).
-The graph of the RMST difference as a function of $\tau$ for the two scenarios are displayed below; $\theta_{\mathrm{RMST}}$ is the same in both settings.
-For RCT scenarios 1 and 2, the ground truth value of the RMST difference at time $\tau=25$ was computed analytically and found to be approximately 7.1.
-Estimation of the RMST
-For each Scenario, we estimate the RMST difference using the methods summarized in Section 5.1.
-The methods used to estimate the nuisance components are indicated in brackets: either logistic regression or random forests for propensity scores and either Cox models or survival random forests for survival and censoring models.
-A naive estimator where censored observations are simply removed and the survival time is averaged for treated and controls is also provided for a naive baseline.
-Figure 8 shows the distribution of the RMST difference for 100 simulations in Scenario 1 and different sample sizes: 500, 1000, 2000, 4000.
-The true value of $\theta_{\mathrm{RMST}}$ is indicated by red dashed line.
-For clarity, only a subset of estimators with particularly informative or interesting results is displayed; the full set of estimators is provided in Figure 15 (Appendix 11).
-Figure 8: Estimation results of the ATE in the simulation of an RCT with independent censoring, using a restricted subset of estimators.
-The complete results are provided in the appendix (Section 11, Figure 15).
-In this setting, and in accordance with the theory, the simplest estimator (unadjusted KM) performs just as well as the others, and presents an extremely small bias (as derived in Section 2.1).
-The naive estimator is biased, as expected, and the bias in the manual G-formula S-learner arises because the treatment effect is additive $T^{(1)}=T^{(0)}+10$ and violates the assumption that $T$ would follow a Cox model in the variables $(X,A)$.
-However, $T|A=a$ is a Cox-model for $a\in\{0,1\}$, which explains the remarkable performance of G-formula (Cox/T-learners) and some of the other models based on a Cox estimation of $S$.
-Some forest-based versions also yield nearly unbiased results, although their performance can vary: for instance, the T-learner remains well-behaved, while the S-learner appears more sensitive and may require larger sample sizes to achieve convergence.
-G-formula (Cox/T-learners) achieves the lowest finite-sample variance at 8,000 observations.
-For completeness, we have also evaluated the other estimators (e.g. IPTW KM, IPCW KM, BJ, IPTW-IPCW KM, IPTW-BJ, AIPTW-AIPCW, causal survival forest and package versions) in Figure 15 (Appendix 11).
-Figure 9: Estimation results of the ATE for the simulation of a RCT with dependent censoring, using a restricted subset of estimators.
-The complete results are provided in the appendix (Section 11, Figure 16).
-Figure 9 shows the results for the RCT simulation with conditionally independent censoring (Scenario 2), focusing on a subset of estimators.
-The naive estimator is biased, as expected.
-The IPCW KM (Cox) is slightly biased up to 4,000 observations and quite variable due to extreme values of censoring probabilities.
-In contrast, the Buckley-James estimator BJ (Cox) is unbiased even with as few as 500 observations.
-The BJ estimator also demonstrates smaller variance than IPCW methods.
-G-formula (Cox/T-learners) seem to perform well, even in small samples.
-As in Scenario 1, G-formula (Cox/ S-learner) remains biased.
-The forest-based estimators shown in Figure 9 appear biased.
-Notably, all estimators exhibit higher variability compared to Scenario 1 likely due to the higher censoring rate in this scenario and the estimation on censoring model.
-For completeness, we have also evaluated the other estimators (e.g. IPTW KM, IPTW-IPCW KM, IPTW-BJ, AIPTW-AIPCW, causal survival forest and package versions) in Figure 16 (Appendix 11).
-6.2 Observational data
-Data Generating Process
-We carry out two simulations of an observational study with both independent and conditional independent censoring as in scenario 1 and 2.
-The only difference lies in the simulation of the propensity score, which is no longer constant.
-For the simulation, an iid $n$-sample $(X_{i},A_{i},C,T_{i}(0),T_{i}(1))_{i\in[n]}$ is generated as in Section 6.1, except for the treatment allocation process that is given by:
-$\operatorname{logit}(e(X))=\beta_{A}^{\top}X\quad\text{where}\quad\beta_{A}=(-1,-1,-2.5,-1),$
-where we recall that $\operatorname{logit}(p)=\log(p/(1-p))$.
-The descriptive statistics for the two observational data with independent (Obs1) and conditionally independent censoring (Obs2) are displayed in Appendix (Section 10.2).
-Note that the survival distribution was not modified, so the target RMST difference remains identical to that of the RCT simulation, namely 7.1 (see Section 6.1).
-Estimation of the RMST
-Figure 10 below shows the distribution of the estimators of $\theta_{\mathrm{RMST}}$ for the observational study with independent censoring.
+A.3. Appendix C: Data Generating Process and Figures
+Simulation coefficients (Section 5.2).
+In the RCT setting, we set $\mu=(1,1,-1,1)$, $\Sigma=\mathrm{Id}_{4}$, and use a Cox baseline rate $\lambda_{0}=0.01$ with $\beta_{0}=(0.5,0.5,-0.5,0.5)$ for $\lambda^{(0)}(t\mid X)=\lambda_{0}\exp\{\beta_{0}^{\top}X\}$.
+Censoring uses $\lambda_{C,0}=0.03$: in Scenario 1, $\lambda_{C}(t)=\lambda_{C,0}$; in Scenario 2, $\lambda_{C}(t\mid X)=\lambda_{C,0}\exp\{\beta_{C}^{\top}X\}$ with $\beta_{C}=(0.7,0.3,-0.25,-0.1)$.
+Treatment is randomized with $e(X)=0.5$.
+Figure 11: Results of the ATE for the simulation of a RCT with independent censoring.
+Figure 12: Estimation results of the ATE for the simulation of a RCT with dependent censoring.
+Simulation coefficients (Section 5.3).
+In the observational setting, we use a logistic propensity model $\operatorname{logit}(e(X))=\beta_{A}^{\top}X$ with $\beta_{A}=(-1,\,-1,\,-2.5,\,-1)$.
+Recall that $\operatorname{logit}(p)=\log\!\big(p/(1-p)\big)$.
+All other components of the data-generating process are identical to those in Section 5.2.
 Figure 13: Estimation results of the ATE for the simulation of an observational study with independent censoring.
-In the simulation of an observational study with independent censoring, confounding bias is introduced, setting it apart from RCT simulations.
-As expected, the naive estimator is biased while IPTW KM (Log.Reg.) is unbiased.
-The top-performing estimator in this scenario is the G-formula (Cox/T-learners), which remains unbiased even with as few as 500 observations and achieves the lowest variance.
-All estimators that use forests to estimate nuisance parameters are biased across sample sizes from 500 to 8000.
-This setting thus highlights that one should either have an a priori knowledge of the specification of the models or large sample size.
-For completeness, we have also evaluated the other estimators (e.g. KM, IPCW KM, BJ, IPTW-IPCW KM, IPTW-BJ, AIPTW-AIPCW, causal survival forest and package versions) in Figure 17 (Appendix 11).
-Figure 11 below shows the distribution of the $\theta_{\mathrm{RMST}}$ estimates for the observational study with conditionally independent censoring.
-The red dashed line represents the true $\theta_{\mathrm{RMST}}$ for $\tau=25$.
 Figure 14: Estimation results of the ATE for the simulation of an observational study with dependent censoring.
-In the simulation of an observational study with conditionally independent censoring, the top-performing estimators in this scenario are G-formula (Cox/T-learners) and AIPCW-AIPTW (Cox & Cox & Log.Reg.), which are unbiased even with 500 observations.
-The G-formula (Cox/T-learners) has the lowest variance.
-Surprisingly, the G-formula (Cox/S-learner) performs quite competitively, showing only a slight bias despite the violation of the proportional hazards assumption.
-All estimators that use forests to estimate nuisance parameters are biased across sample sizes from 500 to 8000.
-Although Causal Survival Forest and AIPTW-AIPCW (Forest) are expected to eventually converge, they remain extremely demanding in terms of sample size.
-For completeness, we have also evaluated the other estimators (e.g. KM, IPCW KM, BJ, IPTW KM and package versions) in Figure 18 (Appendix 11).
-6.3 Misspecification of nuisance components
-Data Generating Process
-We generate an observational study with covariate interactions and conditionally independent censoring.
-The objective is to assess the impact of misspecifying nuisance components; specifically, we will use models that omit interactions to estimate these components.
-This approach enables us to evaluate the robustness properties of various estimators.
-In addition, in this setting forest-based methods are expected to behave better.
-We generate $n$ samples $(X_{i},A_{i},C,T_{i}(0),T_{i}(1))$ as follows:
-  • $X\sim\mathcal{N}\left(\mu,\Sigma\right)$ and $\mu=(0.5,0.5,0.7,0.5)$, $\Sigma=\mathrm{Id}_{4}$.
-  • The hazard function of $T^{(0)}$ is given by
+Simulation coefficients (Section 5.4).
+We set $\mu=(0.1,0.5,0.7,0.4)$ and $\Sigma=\mathrm{Id}_{4}$.
+Let $Z(X)=(X_{1}^{2},X_{2}^{2},X_{3}^{2},X_{4}^{2},\;X_{1}X_{2},X_{1}X_{3},X_{1}X_{4},X_{2}X_{3},X_{2}X_{4},X_{3}X_{4}).$ The control hazard uses
 […]
-    $\text{and}\quad Z$ $=(X_{1}^{2},X_{2}^{2},X_{3}^{2},X_{4}^{2},X_{1}X_{2},X_{1}X_{3},X_{1}X_{4},X_{2}X_{3},X_{2}X_{4},X_{3}X_{4}).$
-  • The distribution of $T^{(1)}$ is the one of $T^{(0)}$ but shifted: $T^{(1)}=T^{(0)}+1$.
-  • The hazard function of $C$ is given by
-[…]
-  • The propensity score is
-[…]
-    When the model is well-specified, the full vector $(X,Z)$ is given as an input of the nuisance parameter models.
-When it is not, only $X$ and the first half of $Z$ corresponding to $(X_{1}^{2},X_{2}^{2},X_{3}^{2},X_{4}^{2})$ is given as an input.
-The true RMST difference at time 0.45 for this scenarios is estimated to be 0.26. The descriptive statistics are given in Appendix (Section 10.3).
-Estimation of the RMST
-First, we estimate $\theta_{\mathrm{RMST}}$ without any model misspecification to confirm the consistency of the estimators under correctly specified nuisance models.
-More specifically, it means that for parametric propensity score models, semi-parametric censoring and survival models, we use models including interactions and squared terms, assuming knowledge on the data generating process.
-Next, we introduce misspecification individually for the treatment model, censoring model, and outcome model (Figure 13), i.e., we use models without interactions to estimate parametric and semi-parametric nuisance components while the data are generated with interactions.
-We further examine combined misspecifications for pairs of models: treatment and censoring, treatment and outcome, and outcome and censoring.
-Finally, we assess the impact of misspecifying all nuisance models simultaneously (Figure 14).
-Figure 12: Estimation results of the ATE for the simulation of an observational study with dependent censoring with interaction.
-As before, for completeness, we have also evaluated the other estimators (e.g. KM, IPCW KM, BJ, IPTW KM and package versions) in Figure 19 (Appendix 11).
 Figure 15: Estimation results of the ATE for an observational study with dependent censoring in case of a single misspecification.
-When there is no misspecification in Figure 12, as expected, IPTW-BJ (Cox & Log.Reg), G-formula (Cox/T-learners) and AIPTW-AIPCW (Cox & Cox & Log.Reg) are unbiased.
-IPTW-IPCW KM (Cox & Log.Reg) exhibits a bias but seems to converge at larger sample size.
-Regarding forest-based methods, IPTW-BJ (Forest), AIPTW-AIPCW (Forest) and Causal Survival Forest estimate accurately the RMST difference.
-Surprisingly, G-formula (Forest/ T-learners), G-formula (Forest/ S-learner) and IPTW-IPCW KM (Forest) exhibit small bias.
-Figure 13 shows that AIPTW-AIPCW (Cox & Cox & Log.Reg) remains unbiased when there is one nuisance misspecification.
-In contrary, the other estimators are biased when one of its nuisance parameter is misspecified.
-The results for the other estimators (e.g., KM, IPCW KM, BJ, IPTW KM) are shown in Figure 20 (Appendix 11).
-Figure 14: Estimation results of the ATE for an observational study with dependent censoring in case of two or more misspecifications.
-Figure 14 shows that, as expected, when all nuisance models are misspecified, all estimators exhibit bias.
-AIPTW-AIPCW (Cox & Cox & Log.Reg.) seems to converge in case where either the outcome and censoring models, or the treatment and censoring models are misspecificed, which deviates from initial expectations (see Section 3.2.3.
-It was anticipated that AIPTW-AIPCW would converge solely when both the censoring and treatment models were misspecified.
-The results for the other estimators (e.g., KM, IPCW KM, BJ, IPTW KM) are shown in Figure 21 (Appendix 11).
-7 Conclusion
-Summary of objectives and findings In this work, we provided a comprehensive overview of causal survival estimators targeting the restricted mean survival time (RMST), emphasizing both their theoretical foundations and practical performance.
-We reviewed a wide range of methods—including regression-based models, weighting approaches, or hybrid estimators-and assessed their behavior in both their parametric and non-parametric versions through extensive simulations.
-These scenarios covered diverse settings, including randomized trials and observational data, with varying censoring mechanisms and model specifications.
-Particular attention was given to robustness, sensitivity to misspecification, and estimator variability in finite samples.
-Practical recommendations for RCTs In simple contexts such as randomized controlled trials (RCTs), whether subject to independent or dependent censoring, we advise against overcomplicating estimation strategies.
-Classical estimators like the Kaplan-Meier survival curve estimator (with correction for dependent correction if necessary) remain highly effective in these settings, providing reliable performance without requiring adjustment.
-Additionally, we recommend the use of the parametric G-formula, even in RCTs, despite its inclusion of covariates which may seem unnecessary when treatment assignment is randomized.
-Indeed, both theoretical results (Karrison and Kocherginsky,, 2018) and our simulations suggest that adjusting for prognostic covariates—even in randomized settings—can reduce the asymptotic variance of the RMST difference estimator.
-We also advocate for the use of the Buckley-James (BJ) transformation in this context, which demonstrated excellent performance in our simulations, particularly in reducing bias under independent or dependent censoring without introducing excessive variability.
-Practical recommendations for Observational studies Beyond simple settings, non-robust weighted estimators such as IPTW-IPCW and IPTW-BJ were consistently marked by high variability, particularly when applied to observational data with dependent censoring.
-Parametric G-formula estimators yielded the best performance under correct model specification (as in Denz et al., (2023)), showing both low finite sample bias and variance.
-However, their sensitivity to model misspecification remains a limitation.
-In observational settings with dependent censoring, we therefore recommend more robust estimators such as AIPTW-AIPCW, which maintain good performance despite their complexity.
-That said, parametric versions may still be preferable in small to moderate sample sizes, as they exhibited lower bias and variance in our simulations—especially for samples under 4,000 observations.
-In contrast, when sample sizes are sufficiently large and model structure is unknown or complex (e.g., involving interactions or nonlinearities), non-parametric estimators become viable alternatives.
-In such cases, we specifically recommend the AIPTW-AIPCW estimator, since the non-parametric G-formula lacks formal asymptotic guarantees and is prone to plug-in bias (Chernozhukov et al.,, 2018).
-Limitations and Strengths It is important to highlight that our simulation design may disadvantage nonparametric methods, as survival and censoring times were generated from exponential distributions and treatment assignments via logistic regression, naturally favoring parametric models.
-Finally, to promote further research and ensure reproducibility, we provide full R code implementing all the estimators discussed in this work and reproducing the simulation results.
-Perspectives A promising avenue for future research lies in exploring variable selection strategies within causal survival analysis.
-In standard causal inference, variable selection involves two main objectives.
-First, ensuring correct identification of the causal effect by appropriately adjusting for confounders—while avoiding the inclusion of instrumental variables, which may increase the variance without reducing bias (Lunceford and Davidian,, 2004).
-Second, improving the efficiency of the estimators by including precision variables—variables that are predictive of the outcome but not of the treatment assignment—known to reduce variance (Lunceford and Davidian,, 2004; Rotnitzky and Smucler,, 2019).
-In the context of causal survival analysis, a new variable appears: censoring.
-It is not clear a priori whether variables that are associated only with the censoring mechanism should be included in the models used for estimation (e.g., in IPCW-based methods).
-Including such variables might reduce bias in the estimation of the RMST difference but could also increase the variance of the final estimator.
-Thus, a promising line of research would be to investigate how different types of covariates—confounders, precision variables, instrumental variables, and censoring-related variables—affect both the bias and variance of RMST estimators under various modeling strategies.
-Finally, RMST-based estimators provide a valuable alternative to hazard ratios, offering interpretable and robust effect estimates that are particularly well-suited for complex, real-world data.
-The tools and simulations provided with this work aim to facilitate their broader adoption in applied causal survival analysis.
-8 Acknowledgments
-This study was funded by Sanofi.
-Charlotte Voinot and Bernard Sebastien are Sanofi employees and may hold shares and/or stock options in the company.
-Imke Mayer and Julie Josse have nothing to disclose.
-The work of CB was supported by the Deutsche Foschungsgemeinschaft (German Research Foundation) on the French-German PRCI ANR ASCAI CA 1488/4-1 “Aktive und Batch-Segmentierung, Clustering und Seriation: Grundlagen der KI”.
+Figure 16: ATEs estimation in observational study with conditionally independent censoring with two or more misspecifications.